A Complete Bibliography of Publications in Biometrika for the decade 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.03

Title word cross-reference

#15763 [BR81]. #15917 [HP80b]. #4143 [AR81].

2 × 2 [Gar85a, Zid84]. α [JE86]. b_{2} [AG83]. Ē^{2} [SK85]. c [JR80]. c^{+}c^{−} [JR80]. χ^{2} [Sie80]. D [AD89, GV83, Guy82]. E [GV83, GS89, Jac80]. F [BMW86, CF84, HW87, Jey82, Phi82, Reg80, Sie80, VS80]. G^{2} [Law84]. H_{o} [GP82]. K [Bar83, Jam87, KR87]. L [DG85]. M [BC88a, FH82, LM86a, Mar82b, Sen82, AR72, AR81, Azz83, FR87]. N [Joe87, Raf88, FR87]. P [FM82, SS82]. π [Sär80]. R [SS89a, MB88]. t [Ede86, Ede89, Hor82, LF80]. U [DR83, Fen83]. U_{N}^{2} [Fre81]. V [Fen83, Bur89]. W [Özt86]. X^{2} [LU80, Law84]. Y^{2} [Law84].

-designs [JE86]. -dimensional [Guy82]. -Estimates [Mar82b, LM86a]. -Estimators [SS89a, DG85, FH82]. -fold [Bur89]. -inverse [Sär80]. -Optimal [GV83, GS89, Jac80]. -Optimum [AD89]. -quantiles [BC88a].

-sample [KR87]. -statistics [DR83]. -Values [SS82]. -way [MB88].

1 [AR72].

49 [AR81].

52 [BR81, HP80b]. 59 [AR81].

62 [BR81, HP76, HP80b]. 65 [dBP81, Pre83]. 66 [CS82, JW81]. 67 [Fuj86, Hos82, Joh82, JM81, O’B82, WL84]. 68 [AO83, Hol82, WR82]. 69 [iA83, HR83, PS84, Pre84b, SL84, Ton83a, Wei83].

70 [FL84, HW84a, MP84, SS85, TJ89, TP84, Tyl84]. 71 [Azz87, Le86, LSW89, Sol86]. 72 [DA87, Fra88, Har87, Sch87b]. 73 [Dav90, Fin90, Hou88b, Kot87, Mar88, Roc90]. 74 [Bro88, Mat88b, Pos90, Ros88a, WFBH88]. 75 [Sch89]. 76 [CP91, Hal90, TKK91].

80a [Pre83]. 80h [dBP81]. 80i [JW81]. 81f [CS82]. 81h [Fuj86, Hos82, Joh82]. 82b [JM81]. 82c [O’B82]. 82e [WL84]. 82h [AO83]. 82m [Hol82]. 83b [WR82]. 83d [PS84, Wei83]. 83g [Ton83a]. 83j [Pre84b]. 84e [HR83]. 84g [iA83, SL84]. 84i [TP84]. 84m [MP84]. 85a [FL84]. 85b [HW84a]. 85c [Azz87]. 85e [Le86]. 85g [TJ89]. 85j [Sol86]. 86b [SS85]. 86e [LSW89]. 87a [Har87]. 87b [DA87]. 87c [Sch87b]. 87i [Dav90]. 87k [Fra88]. 88a [Mar88]. 88b [Roc90]. 88c [Hou88b, Ros88a]. 88e [Fin90]. 88g [Mat88b]. 88h [Bro88]. 88i [WFBH88]. 89c [Sch89].

90f [CP91, Pos90]. 90g [Hal90]. 91c [TKK91].

A. [SD86]. Accelerated [CJ88, Lou81]. Acceptance [GP82]. Adaptive [BK83, Res89]. Addendum [Le86, dBP81]. Additional [JR80]. Additive [O’N82]. Additivity [AO81, DW85, WR80b, AO83]. Adequacy [Mar81, New80a]. Adjusted [Bed83]. Adjusting [TRT85]. Adjustment [BNH88]. Adjustments [DR89, FR89, Har86]. Admissible [Kem84]. Affine [MK85b]. After [BNH88]. Against [Bar83, Des83, GG80, KG87, LM80, LST88, McA83, Nai86]. Aggregate [LM84]. Aggregated [AK83, KA83, Sun86]. Agreement [Pet82]. Agricultural [Mar81]. AIDS [LBD88, SF89]. Akaike [Tay87]. Albert [SD86]. algebraic [Hol86]. Algorithm [AK82, KA89, Lon87, Lun85, Mur83, Pea80, Pet86, RS82, Ton88]. Algorithms [Böh89]. Aligned [Sch85, Sch87b]. Alignment [Gat86]. Allocation [Bro87a, Rob83]. Allowing [GP82]. Alternative [Bow84, Dav87b, Gof87, HLH88, Per84, Tut86, WGPB89]. Alternatives [Bar83, Des83, GG80, Hir82, KY82, LM80, McA83, Nai86, RW82, SW87].

Amendments [iA83, Ano81a, AR81, AO83, Azz87, BR81, Bro88, Che84b, CP91, CS82, Dav90, DA87, Fin90, FL84, Fra88, Fuj86, HW84a, Hal90, HR83, Har87, Hol82, HP76, HP80b, Hos82, Hou88a, Hou88b, Joh82, JW81, JM81, Kee84, Ken82a, Kot87, LSW89, LSG83, Mar88, Mat88b, MP84, MK89, O’B82, Pos90, PS84, Pre83, Pre84b, Roc90, Ros88a, RO89, Sch87b, Sch89, Shi82, SL84, Sol86, TKK91, TJ89, TP84, Ton83a, Tyl84, Wei83, WFBH88, WL84, Yan80]. Among [Ano81a, BO78]. Amongst [Pet82]. Amplitude [HO81]. Amplitude-Dependent [HO81]. Analyse [Ste82]. Analyses [GPT85]. analysing [MB88]. Analysis [Abr80, Ait83, BW81, BP88b, Bar84, Bel86, Bes81, Bow80, BVJM84, BJV88, Bri80, Bro87b, Cas81, CB88, CV89, CS86b, Cri85, EK83, EGG89, Far82, Fri80, GAR85c, Gol86, Gre88, Hal81b, HCT83, Hug89, IO80, Jon83, KTK89, KD87, Kon81, KL83, Krz89, KY82, Kun85, LBD88, LZ86, Lou81, Lus80, MN81, MD89, Mar82a, Mar82c, Mat89, MT80, MW80, Nai86, OV81, Pat85b, Pat86, PWP81, QJ86, RA86, RRK87, Roc83, Ros87, Ros88b, SH80, SS83a, Sha85, SS84a, Sol84, Swe87, TT85, Tut86, Tyl87, Ver88, VS80, WR80a, ZLS85, Bro88, Ros88a, Sol86, WR82]. Ancillarity [iA82, Cox80, FM81, God80a, God84a, Lia83, Rya81, Sko86, iA83]. Ancillary [CS89a]. Anderson [SD86]. Angular [FL81, FL82, Tyl87, WJ80, Wel85]. Angular-Angular [FL82]. Angular-Linear [FL81]. Animal [BT83, Por86]. Animals [Ano81a, BO78]. Annealing [Lun85]. Antithetic [Hal89b]. Application [AD89, BLE88, BHS85, BT83, Blæ81, Dau80, Daw81, HS80, KS81, LBD88, MT81, TGG89]. Applications [Bas81b, CS89b, CFR88, Las82, Lun85, McC82, PS85, SCR86, Wor82]. Applied [SvW86]. Approach [BK87b, CB88, Dar80, Dea80, FR88, MN81, MOL86, O’N82, Per81, Pes84, Pet86, Raf88, Tho86, Wan87, Wel85, Yeh88]. Approximate [But89, CRC89, Dav86, Dav90, Dur80a, Fuj80b, Gar85a, Gar87, Hin80, Kai83, KTK89, Lia87a, Pet83, Shi84, Swe87, TGG89, TKK89, TKK91, Upt86]. Approximating [WT85]. Approximation [Ali84, Bau86, CS88a, DH85, FP80, Feu89, Gol89a, Hal82b, Joh80, Kim88, Lev84, Min83, MT80, RW83, Joh82, LU80]. Approximations [Bin81, Dan80, Dan83, DH88, Dur80b, SK85, SW87]. Arbitrary [PW82]. Arguing [Pal81]. Arising [Jew85]. Armed [HT81a, HT81b]. Arrays [GN87]. Arrival [PTF89]. Aspects [Mar82c]. Assays [DCS88]. Assessing [BE86, FS85, KTK89, Koz82, Min83, TC89, Yan84]. Assessment [Lau83]. Associated [CS89b, RW83]. Association [Cuz82, Dal84, Dau80, FL81, FL82, Goo81, Hir83, Oak86, Wah80]. Associations [CS86a]. Assumption [GS87, MOL86]. Assumptions [Gol80]. Asymmetric [DW82, LM80]. Asymmetrical [Lig89]. Asymmetry [Mor87a]. Asymptotic [iA82, BF88, CC88, Dav87c, Fuj80a, Fuj86, HM85, Har85a, HP85, Hil81, Hos80, Kai83, KR86, Kot86, Lar83, Lia84, LM85b, Moo86, MW80, Nur81, Pes84, Phi87, PR87a, RW87, Roy89, Sch81a, Sch81b, Sha85, Sic82, TGH83, TT88, Tsi81, Win80, Wu85, iA83, FH82, Har87, Hos82, Kot87, Wil84]. Asymptotically [AL89, SE86, Tam88a]. Autocorrelated [BB89b, GG80, Mat87, Mat88b]. Autocorrelation [Bel84, DA85, Mar80, DA87]. Autocorrelations [AL84, Hos80, Li85, MP83, Hos82, MP84]. Autocovariance [DA85, DA87]. Autocovariances [Ahn88, New80b]. Autoregression [Phi87, SS83b]. Autoregressions [Kün87b, Mar82b]. Autoregressive [Ahn88, AT84, AK83, Azz84, AG87, BHS85, Bha80, CW81b, EN88, Fra85, God80b, GU83, God84b, HO81, HR82, HK84a, Hos80, Kan87, Li85, Lju82, LST88, LT87, MP89, Mas87, McL84, Mon84, NP84, Pea80, PD86, Pie85, PT80, Pos87, Rei83, SD84, TG80, Ton82, Ton83b, Tsa86a, Yam81, Azz87, Fra88, HR83, Hos82, LM86a, Pos90, Ton83a]. Autoregressive- [McL84, Pea80]. Autoregressive-Moving [AK83, CW81b, Fra85, God80b, GU83, God84b, HR82, HK84a, Hos80, Lju82, Mas87, Mon84, PT80, Pos87, SD84, TG80, Yam81, Fra88, HR83, Hos82, LM86a, Pos90]. Autoregressive-Type [PD86]. Auxiliary [SJ81]. Average [AK83, AK85, CW81b, CL81, Des83, Fin86, Fra85, God80b, GD82, GU83, God84b, Hal89a, HLP88, HR82, HK84a, Hen86, Hos80, KP81, KA84, KA86, Lju82, Mas87, McL84, Mon84, New80b, Pea80, PT80, Pos87, SD84, TG80, Yam81, Fin90, Fra88, Hal90, HR83, Hos82, LM86a, Pos90]. Averages [Bro81, Dav85b]. Axis [Cla80].

Back [Ano80a, Ano80b, Ano80c, Ano81b, Ano81c, Ano81d, Ano82a, Ano82b, Ano82c, Ano83a, Ano83b, Ano83c, Ano84a, Ano84b, Ano84c, Ano85a, Ano85b, Ano85c, Ano86a, Ano86b, Ano86c, Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88b, Ano88c, Ano88d, Ano89a, Ano89b, Ano89c, Ano89d, OS88]. Bahadur [Pes84]. Balanced [Afs83, AP82, CW81a, Che86, GS85, GJ83, Kun87a, MK85b, RP82, SE86, Sre89, SC85]. Balancing [Rus80, Ste80]. Band [Wyn84]. Bandit [HT81a]. Bands [HW80, NP84]. Bandwidth [HM85, MSS87]. Bartlett [BNH88, Gla80, Har86, Nag84]. Based [BN86, CFR88, DR89, ESP82, EP83, HW83, Hal87, Hom88, Hor82, Kam89, KD87, Kim85, Kot88, Kou80, MV82, MI81, Mat88a, MBBW81, PW80a, QJ86, SH80, Ski88, Tam86, Tar85, Tsa88, Tut86, Wei88, HW84a]. Batches [FS84]. Bayes [BW81, CS85, Dea80, GV83, Gol80, Kuo88, Raf88, Smi80b, SV80]. Bayesian [Ban88, Bes89, Bin81, BK87b, Bro87a, Cas81, CB88, CS89d, Dar80, Daw81, Eav83, EGG89, KTK89, Naz87, Per81, Per84, RA86, RP82, SS84a, Swe87, Tho86, Yeh88]. be [Jey82]. Bean [Blæ81]. Behaviour [BF88, Jen83, Vas81]. Behrens [Fen83, FR82, Smi80a]. Bernoulli [Gün87, SS89b]. Best [Cur88, Tam86, Sär80]. Beta [Gla80]. Between [Don87, Fen83, Gil84, HM84, Kee82, Kee84, Krz83, MMM82, Per84, She82, Yan78, Yan80]. Bias [CTW86, Dav85a, GPT85, Mil84, Par83, PGT86, Ste80, TP83, TP84, Whi86]. Biased [Atk82, Bol88, GWP84]. Binary [AO81, CSL^{+}84, CL88, Fow87, GJ82, Hal81a, KL87, ZLS85, AO83]. Binomial [BP88b, Chi80, CS89d, Fuj80b, GAR85c, Hal82b, PGT86, Raf88, Ruk88, Wor83]. Bioassay [Dar80, Kuo88, MT83, MH80, RO88, RO89]. Biometrika [HP80b, dBP81, iA83, AR81, AO83, Azz87, BR81, Bro88, CP91, CS82, Dav90, DA87, Fin90, FL84, Fra88, Fuj86, HW84a, Hal90, HR83, Har87, Hol82, HP76, Hos82, Hou88b, Joh82, JW81, JM81, Kot87, Le86, LSW89, Mar88, Mat88b, MP84, O’B82, Pos90, PS84, Pre83, Pre84b, Roc90, Ros88a, Sch87b, SS85, Sch89, SL84, Sol86, TKK91, TJ89, TP84, Ton83a, Tyl84, WR82, Wei83, WFBH88, WL84]. Biplots [GH88]. Birth [AD80a, Bas81b]. Bivariate [Bur88, CMH87, Cam81, Cla80, CS89d, D’E81, Dal84, DCS87, Gen87, Goo81, Leu83a, LKP86, Oak86, PKL83, Taw88, Wah80, Wan87, WJ80]. Biweight [MI81]. Block [AP82, CW81a, Che86, GS85, GV83, GJ83, HJ81, Kun87a, Pat83a, Pat83b, Sch85, SE86, Sre89, WL81, Yeh86, GS89, Jac80, Sch87b]. Blocking [AD89]. Blocks [Bai87, ML88b]. Bonferroni [Fal89, Hoc88, Hom88, Hom89, Sim86, Wor82]. Bonferroni-Type [Fal89]. Bootstrap [Ban88, BM86, DHS86, DJRT85, Efr81, Efr85, Hal87, HM88, Hal89b, Hal89c, HS87, HS89, Par83, SY87, SvW86, Tay89, Tib88]. Bootstrapping [You88]. Boscovich [Sti84]. Both [Won89]. Bound [BB86, Mar82b, Pat83b, PS82, PS84]. Boundaries [Bau86, DW82, LD83]. Bounded [Smi80b, SCR86]. Bounds [Coo80, Ede86, GW80, Jar84, MV82, Sch84b, Van80]. Box [GJ82, Swe84]. Brain [TMKG87]. Branching [Bas81b]. Brownian [PS85]. Buckley [STV84]. Buffon [Hol82, Hol81]. Building [Lüt82].

Calculated [Dur80a, Tsi81]. Calculation [BG84, Mor80]. Calculations [GLR81, WGPB89]. Calibration [BS89]. Candidata [Ede89]. Candidate [Bes89]. Canonical [BFR89, DM85, Gil84, HA87, Krz89, MW80, Pat83b, Ren88, TT85]. Caps [ME82]. Capture [Ano81a, BO78, Hug89, SC84]. Capture- [SC84]. Carcinogenesis [Por86]. Carlo [BC89]. Carry [Rus80]. Carry-Over [Rus80]. Carryover [Mat87, Mat88b]. Case [BG87, BF88, BC88b, CF85, Ekb82, FP80, GLR81, Gar85b, Hal80, Lub81, Pre86, SW89, Shu86, VM88, WGPB89, Yan84, FJ89]. Case-Cohort [Pre86, WGPB89]. Case-Control [BG87, BC88b, FP80, GLR81, Gar85b, Hal80, Lub81, SW89, Shu86, VM88, Yan84]. Cases [BG87, Smi85, KR87]. Categorical [Bow80, Gre88, Krz83, LS85b, Pet84, Por87]. Categories [Goo81, Wah80]. Cauchy [Gab82, Kou82]. Causal [RR83]. Causes [Gof87]. Censored [BLS^{+}82, Cam81, Che84a, Cuz82, DL86, HW80, HF82, IR85, Jam86, JW79, KO88, Koz80, LR81, Leu83b, MMM82, Meh81, MH82, Nai81, PW80a, Pet83, Pet86, Pre78, Reg80, Sch85, Sch84a, SW86, SL82, STV84, Swe87, Tsa88, TRT85, Tur89, WL80, WL81, Zhe88, Jam87, JW81, Pre83, Sch87b, SL84, WL84]. Censoring [Bro82a, Jen83, Jen84, Leu83b, Leu87, TJW87, Var89]. Censorship [Bur88, CH81b, Kum87, PW82]. Central [RR83, Tyl87]. Certain [Ros87, TGH83, Ros88a]. Chain [Vas81]. Chains [LM84, Swa83, SS85]. Change [AR88, Fle87, JJS87, KS89, MF85, NRW84, Pet80b, Pet81, PS85, RA86, Wor83, Wor86]. Change-Over [Fle87]. Change-Point [JJS87, KS89, MF85, NRW84, Pet80b, Pet81, RA86, Wor86]. Changepoint [Lom87]. Changes [RZ87]. Changing [Pat86, Ver88]. Chao [Sen89]. Characteristic [EP83, HW83, Kou80, Kou82, MT81, Phi82, HW84a]. Characterization [Gla80, Tam86]. Checking [ABH89]. Chi [Bed83, FS85, Hir86, Kou86, LSW84, Sch80, TA83, LSW89]. Chi-Plots [FS85]. Chi-Squared [Bed83, Hir86, Kou86, LSW84, Sch80, TA83, LSW89]. Choice [Atk80, BLE88, Bro84, LG84, MSS87, Swe84, Tay89, Tut86]. Circle [HR85]. Circuits [Pat83a]. Circular [FL83a, FL83b, FL84]. Class [Des83, Efr85, EGG89, Fle87, Gil84, HF82, Hou86a, Hou88a, HHP82, Koz82, MBBW81, MM89, Smi85, SJ81, Vec85, WV81]. Classes [Gla86a, Leu83b]. Classical [BLE88, FS84]. Classification [Far82, Kee82, Kee84, Kok81, OV81, Shu82, WR82]. Classified [Bed83, BJV88]. Clinical [Atk82, DW82, DW80, GTP88, Gea88, GP82, HT81b, LD83, Pat85b, Pat86, Pet80a, TGG89, TSE88]. Closed [Ano81a, BO78, Cas81, Kim88, LOdBP88]. Cluster [Bri80, Mar82a]. Clustering [Bin81]. Clusters [WFBH87, WFBH88]. Coefficient [AD80a, FL83a, GL88, JM80, Jup87, LM89, Mal86, SK88, Wah80, You88, JM81]. Coefficients [Dur80a, FL86, Hil86, TC89]. Cohort [Pre86, WGPB89]. Coin [Atk82]. Collapsibility [AE83]. Collections [Azz84, Azz87]. Column [IJ85, JE86]. Columns [AP82, Che86, Sre89]. combination [Gar85a]. Combinatorial [Lun85]. Combining [WJ85]. Comments [Gab82, WR82]. Common [FL83b, Flu87, Nur81, Sch88a, Sch89, TGH83, FL84, Gar85a]. Commonly [TP83, TP84]. Commonly-Used [TP83, TP84]. Community [BH83]. Comparative [Pet80a, TSE88, Bur89]. Comparing [BCE83, DM88a, Ekb82, Gup89, Jen84, Pal87, PR87b, SO82, Sch81b, WGJJ89]. Comparison [AL89, CS86a, DDS89, Dav87a, DHR89, FD83, Fen83, Gab81, HLH88, Hom89, Lei88, MR87, Mat87, MB89, Meh81, NH80, Pet84, Rob83, RT88, Swa83, Mat88b, SS85]. Comparisons [Che82, Che84b, CW80, Law84, Pes84, SE86]. Competing [BDK89, HH89, SR83]. Competitors [KG87]. Complex [Por87]. Complexity [HH88]. Component [Ait83, DW85, Flu87, Gol87, Sch88a, Sch89]. Components [Cri85, DM85, FH83, MH88, Sol85, TGG89]. Composite [Sas80]. Compositional [Ait83, She82]. Concentration [DCS88]. Concept [FM81, Rya81]. Conceptual [ĆGKP82]. Concordance [Kra81]. Condition [Jey82]. Conditional [iA82, Bas81a, Bas81b, CL88, CRC89, FR88, Fri80, HS87, Lia84, Lia87a, Lin82, Lub81, SC87, iA83]. Conditionality [BN80]. Conditionally [Pal81]. Conditioning [Ski88, Ver86]. Conditions [BF88, BFR89, Gab84, Gof87, Yeh86, SD86]. Cones [DJRT85]. Confidence [AR72, Bas81a, Ber87, CT80, DD87, DM88a, Dav87c, DJRT85, Efr85, Fuj80b, HW80, Hal82b, Hal87, Ham86, HW84b, JT85, Krz89, Mat88a, MSS87, NP84, Owe88, RT88, Tab87, Upt86, Win80, Wor86, Wyn84, AR81, Sie80]. Confined [CS89a]. Confounding [Yan84]. Conjecture [Ste80]. Consequences [Fri80]. Conservative [DM88a, Gla86a]. Considerations [Daw81]. Consistency [Bow80, GL88, Mau83, Slu82]. Consistent [Bri89, JR81, Mas87, RW89]. Constancy [LM89]. Constraints [Las82, Lee85, Sha85]. Construct [Tab87]. Constructed [FS80]. Constructing [Che86, Hal82b, MPW88]. Construction [AP82, AD89, Sin87, Sre89, WW87, Jac80]. Constructions [Böh89]. Contemporaneous [CMH87]. Contingency [AE83, Bri80, CP80, EK83, EH85, EGG89, Fri80, Gar87, Gil84, Goo81, Hir83, Naz87, TA83, Wah80, Wan87, WL83, MB88]. Continuous [Krz83, LS85b, Pet84, Ste82, Ton83b]. Contrast [ABS84]. Contrasts [ZH89]. Control [BG87, BC88b, CW80, FP80, GLR81, Gar85b, Gup89, Hal80, Lub81, PR87b, SW89, Shu86, SE86, VM88, Yan84]. Controls [Ros88b]. Convergence [AK85]. Convex [PTF89]. Convolutions [LR81]. Cornfield [Lev84]. Correct [Dri88, Kok81]. Correcting [Hal85]. Correction [HW87, Kot88, Mil84]. Corrections [iA83, Ano81a, AO83, Azz87, Bro88, Che84b, Cor87, CP91, CS82, Dav90, DA87, FL84, Fra88, Gar87, HW84a, HR83, Har87, Hol82, HP76, Hos82, Hou88a, Hou88b, Joh82, Kee84, Ken82a, Kot87, LSW89, LSG83, Mar88, Mat88b, MP84, MK89, O’B82, Pos90, PS84, Pre83, Pre84b, Ros88a, RO89, Sch87b, Sch89, Shi82, SL84, TA83, TKK91, TJ89, TP84, Ton83a, Tyl84, Wei83, WFBH88, WL84, Yan80, AR81, BR81, Fin90, Fuj86, Hal90, HP80b, JW81, JM81, Roc90, Sol86]. Correlated [Bho84, CL88, Con89, Cur88, Fri80, GS85, Har85b, Kun85, Kun87a, Mar82c, MB89, MS89, Nag82, OP84, OV81, RGL85, WR82]. Correlation [BB81, Bha80, BVJM84, BJV88, BFR89, Che84a, CS88b, DK80, Dur80a, FL83a, FL86, Gil84, HK84b, JM80, Jup87, Ken83, Mar86, MW80, O’B80, Sch88b, SK88, You88, JM81, Mar88, O’B82]. Correlations [Ber89, BVJM84, BE86, Ren88, Roy89, Sri84]. Correspondence [Gre88]. Costs [GP82]. Count [BP88b]. Counts [Hal85, Moo86, Zeg88]. Covariance [Bel86, CF84, CF85, GU83, Gol87, GT84b, HP85, Jey82, MR87, MM84, Por85a, Roy89, RS82, Sol85, WR87]. Covariances [DM88b]. Covariate [Pre82, Pre84b, Sch87a]. Covariates [Fra84, GWP84, Hil81, Lag88, Lus80, Pat86, RO88, RO89, TT88, TRT85, Ver88, ZLS85]. Cox [AA86, GJ82, SP82, Swe84]. Cramér [SL84, Bro82b, Mar82b, SL82]. Criteria [CL87, HP85, Sch84b]. Criterion [Atk80, Dav80, Hur88, SH80, Tay87, Wel83]. Cross [AK87, Bed83, Bow84, BHT84, BJV88, De 88a, FS84, Gre85, Hal82a, KA89, Le84a, Shu86, Bur89]. Cross-Classified [Bed83, BJV88]. Cross-Over [Le84a]. Cross-Product [Shu86]. Cross-Validation [AK87, Bow84, BHT84, De 88a, Gre85, Hal82a, KA89, Bur89]. Cross-Validatory [FS84]. Crossings [CL80, KS81, KR86]. Crossover [Mat87, Mat89, Pat85b, PR87b, Mat88b]. Cubic [MJ89]. Cumulant [MB88, Tan82]. Cumulants [McC84b, PGT86]. Cumulative [DDS89, Hir82, Hir86, Kha81, Pet80b, Wor83]. Curious [Bes89]. Curve [HW80, Kuo88, Las82, Ver86, VV88, Whe80]. Curves [Cha87, FD83, Jen84, NH80, SR83, TJ82, WS81]. Cusum [Gol89a]. Cyclic [DK87b, IJ85]. cycloid [Hol81, Hol82].

D [Roc90]. D. [MOL86]. Daniels [Wil84]. Data [Ait83, AK83, AO81, BP88b, Bed83, BLS^{+}82, Ber89, Bho84, Blæ81, Bow80, Bre81, BC88b, Bur82, Cam81, CD86, Cha87, Che84a, CP81, CL88, Cox82, Cuz82, Dav87a, DP88, DL86, DCS87, Don87, DJRT85, DM87, Ekb82, Far82, FL83a, GAR85c, Gof87, Gre88, HW80, HWC87, HF82, Hur88, IR85, Jam86, Jew85, JW79, JM80, KL87, KO88, KA83, Koz80, LM85a, LR81, LBD88, LL83, LM84, Leu83b, Leu87, Lia85, LS85a, LZ86, LS85b, Lub81, MN81, MV82, MD89, Mar81, Mat89, MMM82, Meh81, MH82, Moo80, MSS87, Nai81, Oak86, Pat85b, Pet83, Por87, Pre78, PWP81, Pre84a, Reg80, RRK87, Sch85, SW86, ST87, She82, Sir82, Sol84, Sri84, SP82, SF89, Sun86, STV84, Swe87, TG80, TJ83, Tsa88, TRT85, TMKG87, Tur89, WS81]. Data [Wei82, Wel85, WGJJ89, Wil83, ZLS85, Zhe88, AO83, Jam87, JW81, JM81, MB88, Pet84, Pre83, Sch87b, Sol86, TJ89, Wei83]. Death [GLR81, GG81]. Decaying [Ber89]. Decision [Ton82, Ton83a]. Decomposition [FL89, Gla86b]. Deconvolution [Var89]. Decreasing [Var89]. Defining [Hir83]. Definition [HLH88, Kra81]. Degeneration [Bre84]. Degrees [Cox84, Kai83]. Delay [ZF89]. Delayed [DH85, Ton83b]. Delta [Par83]. Densities [BHT84, But89, Dur80b, TKK89, TKK91]. Density [BZ88, Bow84, ESSAH89, EN88, Hal82a, HWC87, HW88, HH88, Las82, Lig88, Lig89, Ng80, SvW86, Tan82, Tay89, Tho86, Tut86, Var89]. Departure [MI81]. Dependence [ĆGKP82, FS85, IS86, KO88, Kie82, KG87, PS81, TA83, WJ80]. Dependencies [MS88]. Dependent [Don87, HO81, Jew85, Lia85, Lus80, RZ87, SK81, Saw83, SS89b, Sir82, SR83, WJ85]. Derivation [AK82]. Derivatives [MSS87]. Derived [Bur82, Col89, Hou86b, Hou88b]. Describing [Hou84, She82]. Description [Fin80]. Deshpande [BB89a]. Design [CS86a, CT80, Cox88, FM82, FS80, FTW85, KD87, Mar82c, Mar86, MH88, Pat85b, Pre86, Sam81, SS84a, SW83, Tam88a, VS80, WGPB89, Wel83, Wu85, Mar88]. Design-Unbiased [Tam88a]. Designs [Afs83, AP82, Atk82, AD89, AG87, BT83, CW81a, Che86, Con89, FK86, Fle87, GS85, GV83, GJ83, Gup89, HJ81, IJ85, Kun85, Kun87a, Le84a, Mat87, ML88b, MH85, MK85b, OS88, Pat86, Pat83a, Pat83b, PR87b, RL84, Sch85, Sin87, SV80, SE86, Sre89, Ste85b, SC85, SS84b, TSE88, Ver88, Wel87, WW87, WR80a, WL81, Yeh86, FR87, GS89, Jac80, JE86, Mat88b, Sch87b]. Destructive [BT83]. Detect [WFBH87, WFBH88]. Detectable [DCS88]. Detecting [AR88, MS88, PS85, Tam82]. Detection [Bri89, CB88]. Determination [Hal80, Pos87, Yan78, Yan80, Jac80, Pos90]. Determined [Sas80]. Determining [PT87]. Determinism [BB87]. Deterministic [FH83]. Deviance [DG89]. Diagnosis [Lau81]. Diagnostic [BD87, Lau83, Lju86, Spi83, WGJJ89]. Diagnostics [BS89, CW83, Fow87, Hin85]. Diagrams [SG85]. Did [TE87]. Difference [HM84, Wei81, Wei82, Wei83]. Differences [DP88, HW88, MB89]. Differencing [Hos81]. Different [FS84]. Differential [FR88]. Differing [FL83b, FL84]. Difficulties [JR81]. Diffusion [AD80a, PS85]. Dimensional [Blæ81, DF85, Kün87b, Las82, Vec85, Wyn84, Guy82]. Dimensionality [Sch84b]. Direct [God84b, Lau83]. Direction [FL83b, Upt86, FL84]. Directional [DJRT85, DM87, JM80, Pre84a, WS81, JM81]. Directions [Jup87]. Dirichlet [CJ88]. disaggregations [Zid84]. Discontinuous [Ton82, Ton83a]. Discounted [Lüt82]. Discovered [TE87]. Discrete [Bau86, Hir82, LD83, Tut86, WV81, Fre81, FJ89]. Discriminant [CF84, Dav87c, Hal81b, Tut86]. Discriminating [She82]. Discrimination [AB82a, BMW86, CF85, CFR88, FK86, Hal81a, HW88, MS86, Per84]. Discs [Dup80]. Disease [AW80, Bec81, BH83, Pre86, WFBH87, WFBH88]. Diseases [WH89]. Dispersion [Mat89, SO82]. Dispersions [FL83b, FL84]. Displays [Atk81]. Distance [GW80, Krz83, MK85a, MK89]. Distributed [AW80, Fin86, Fin90]. Distribution [Ahn88, AH89, Ali84, AG83, AR72, AR88, Azz81, BDK89, BN83, Blæ81, Bre82, Bur88, CD86, Chi80, Col89, Cox88, CS89d, D’E81, DG82, Daw81, Dur80a, Gab82, GW80, Gol89a, GT84b, HD82, Har85a, Hin80, Hor82, Hos80, Hos84, Joh80, Koc81, Kou86, Kuk88, Leu83a, LM89, Li85, LS85c, LKP86, Mor80, MT80, MT81, Özt86, PW80a, PKL83, Phi82, Por86, Pre84a, PW80b, RW87, SO82, SK85, Sha85, Sic82, Ste82, Swe84, Tsi81, Tyl87, Upt86, AR81, Fre81, Har87, Hos82, Joh82, LU80, Smi80a]. Distribution-Free [HD82, Koc81, Leu83a, Pre84a, SO82, Smi80a]. Distributional [Spi83]. Distributions [AS80, AD80b, Blæ81, Bro82b, DK87b, ESP82, FL83b, Fuj80a, Gen87, HJ82, HP85, Hou84, Hou86a, Hou86b, Hou88a, HK84b, Kee82, Kee84, Kim85, Kou82, ME82, Meh81, MK85a, MW80, Nag80, RW83, Sch81b, Sch84b, SG85, Wah80, WV81, Wan87, Wes87, DG85, FH82, FL84, Fuj86, Hou88b, MK89]. Divisible [Sin87]. Dolby [Gle85]. Domain [Mil84, PN85]. Dose [SM86]. Dose-Response [SM86]. Doses [SM86]. Double [AR88, Pal87]. Downturns [SM86]. Drift [PS85]. Dual [DF89, Sri80]. Duality [FK86, McL84, Wol88]. Durbin [Fra88, Ali84, Bha89, Fra85, JR80, MJ89, Ton88]. Dynamic [Bol88].

Early [GP82]. Edge [DK87a, Hal85]. Edgeworth [BZ88]. Effect [BK87a, CS86a, DH87, EN88, GWP84, GTP88, GPT85, PS81, RZ87, Sol84, Sol86, Yan84, DG85]. Effective [BLS^{+}82, Cox84]. Effects [Cur88, DK87a, Dav80, Hal85, KL83, Lon87, Mat87, Pet86, RR83, Rus80, Ste85a, Wer89, Mat88b]. Efficiencies [Sic82, dBP81]. Efficiency [Bas81a, BKB89, CA83, Fir87, GS85, GL88, Hil81, Lag88, Lia84, MV82, Mor81a, Nur81, Pat83a, Pat83b, PW86, Sam81, Shi84, Slu82, SP82, TT88, Tur89, Tyl83, WGPB89, Tyl84, Wil84]. Efficient [AK87, Bas81b, BT83, DK87a, DHS86, Gup89, Hal89c, HP80a, Har85a, Hir82, Jen87, KA85, New80b, RS80, Sch84a, Tar85, Tsi81, Har87]. Efron [Ste80]. Egon [BT81]. Eliminating [SM89]. Elimination [SC85]. Elliptic [MK85a, MK89]. Elliptical [BB86, HP85, Kou86]. Emphasis [FP80]. Empirical [CS85, Dav85a, Dea80, DR89, DHR89, ESP82, EP83, Feu89, GPT85, HW83, Jon83, Kou80, Kou82, Lub81, MT81, Owe88, HW84a]. Enforcing [Mon84]. Entropy [Lar83]. Entry [Gar87]. Enzyme [Eav83]. Epidemic [CS88a]. Epidemiologic [Pre86]. Epidemiological [Sam81]. Equal [BVJM84, CF85]. Equality [Bar83, Bho84, Che85, FD83, GT84b, Koc81, MM89, Nag80, Sch88b, WH89]. Equation [Lee85]. Equations [Dan83, Fer82, GT84a, Jam86, Mor81a, Mor81b, Hol86]. Equilibrium [DK87b]. Equioverlapping [Kam89]. Equireplicate [GJ83]. Equivalence [Gar87, Kai83, LM85a, New80a]. Equivalent [AL89]. Error [AK86, BNH88, BB87, Ber87, Cha82a, Efr81, Gla86a, Gla88, Hur88, KD87, LG84, Lig88, Sch87a, Sch81a, Ste85a, SC87, Toy82a, Toy82b, Wel83, Won89]. Errors [AG87, BB89b, CSL^{+}84, CM84, CS79, Gla88, Hal89a, Kun87a, Mar82c, Mat87, ML88b, OV81, Pet86, PG82, Pre82, RGL85, Tam88b, WF82, Zam89, CS82, Hal90, Mat88b, Pre84b, WR82]. Errors-in-Variables [CSL^{+}84, Zam89]. Estimate [MH85]. Estimated [AA86, AK86, CC88, CS89c, CF84, HM84, Toy82a, Toy82b]. Estimates [AA84, AD80b, Bha80, BF88, Bow84, DD87, DM88b, Efr81, GWP84, Gla86a, GL88, Lei88, MV82, Mar82b, Mas87, MJ83, Naz87, RS82, SH82, Smi80b, TP83, Tut86, Tyl82, LM86a, SD86, TP84]. Estimating [BE89, CD86, Cro87, Dan83, EN88, Fer82, FL83b, FL84, FS80, FS84, GT84a, Hal85, Jam86, LR81, Lia87a, ML88a, Mat89, Mor81a, Mor81b, Nay88, Por86, Rei81, Rit89, Ruk88, SSW89, SP82, ZF89]. Estimation [AD80a, AB82a, Ano81a, Azz81, Azz84, Azz87, BB80, BB87, BK87b, BHT84, Bre82, Bre84, BES88, Bur88, BO78, Cam81, CS89a, CH81a, CM83, Coo80, CW81b, CS79, CS85, CF85, CFR88, CL80, DK87a, DP88, DK87b, DiC84, Dup80, DF89, EM81, ESSAH89, EN88, FK86, Fir87, Fra80, GPT85, GT84a, God85, GD82, Gol89b, GG81, GN87, Hal82a, HJ82, HWC87, HH88, HR82, HK84a, HKM86, HT81a, HT81b, HS81, IR85, Joe87, Jør83, KP81, Kam89, Kan87, Kim88, KF83, KA83, KA85, Kou82, Kra81, Kuk88, KiA84, LG84, Le84b, LOdBP88, Lig88, Lig89, LS85b, Lon87, LKP86, LT87, MP89, Mak81, Mak85, Mal86, MM84, Mar80, Mat81, MT83, Mil84, MH88, MS89, New80b, Ng80, NRW84, OV81, PW80a, PW82, PTF89, PKL83, PS82, Pre82]. Estimation [Pre84a, PW80b, PN85, RL84, RGL85, RR88, SK81, SW86, Sic82, Smi85, Spr80, Sri84, SK88, Ste85a, Ste87, SvW86, Tam84, Tam88b, Tan82, Taw88, Tay89, Tho86, TJ83, Tsa88, TMKG87, Van80, Var89, Wal88, WR87, Wei81, Wei82, WSM89, Whi86, Wil80, WF82, Won89, Wu82, Zam89, Azz83, CS82, Gar85a, Guy82, HR83, KR87, Le86, PS84, Pre84b, TJ89, WR82, Wei83].

Estimator [AJ89, BN83, BNJL89, CS88a, CL81, DGS87, DF85, GU83, God84b, Gol80, Hal81b, HD82, Kot88, Kum87, MI81, PS80, SSW89, Shu86, Sir82, Sri80, STV84, Toy82b, TJW87, Wu82, ZH89, Wil84]. Estimators [BW81, Bas81a, Bre81, Bro81, CA83, Dur80b, Gla86a, JR81, KY83, Kuo88, Mau83, MH80, Moo86, Mor87a, Nur81, Pal87, Par83, RZ87, RS80, RW87, Ruk88, Sch84a, SS89a, SS80, SJ81, TGH83, WV81, WS81, Whe80, ZH89, DG85, FH82]. Evaluate [SS82]. Evaluation [GU83, KY83, Lag81, Lub81]. Even [LKP86]. Events [Bar83]. Evidence [Gro80, Gün87]. Exact [AK83, AD89, Dan80, Jen84, LOdBP88, McA83, MPW88, Pea80, PR87a, RT88, Wei88, Wyn84]. Example [Eav83]. Exciting [PD86]. Exclusive [Bar83]. Existence [AA84, SD86]. Expansion [Har85a, Kai83, Toy82b, Har87]. Expansions [Fuj80a, Jar84, KF83, Sch81a, Win80, Fuj86]. Expectancy [Kum87]. Expectation [MP83, MP84]. expectations [LU80]. Expected [DG82, Kok81, LSW84, LSW89]. Experiment [SF81]. Experimental [AD89, Mar82c, VS80]. Experiments [ABS84, Bas81b, Bel84, BT83, BKB89, DW85, Fle87, GWP84, Hug89, Lei88, Mar86, Por86, RO88, RO89, SC84, Wel83, Wil86, ZH89, Mar88]. Expert [Tam88b]. Explanatory [KL87, Lag81, Pal81]. Exponential [BMW86, Bas81a, Che82, Che84b, CV89, CP89, Cox88, Har89, HT81b, IS86, Kim85, Mal88, Meh81, Nag80, Reg80, Wor86, CP91]. Exponentiality [BB89a, Des83, SL82, SL84]. Extended [NP87]. Extending [Fer81]. Extension [MJ89, VV88]. Extensions [Rya81]. Extra [Mor87b]. Extra-Poisson [Mor87b]. Extraction [KA87, KA89]. Extreme [AR72, HS80, Hos84, MT83, Özt86, PW80b, Taw88, AR81]. Extreme- [Hos84]. Extreme-Value [PW80b].

Factor [Bar84, Bro87b, IO80, Pat83b, Yan84, Bro88]. Factorial [Fle87, SC85]. Factorially [O’N82]. Factorially-Structure [O’N82]. Factorials [Ste85b]. Factors [BJV88, HW87, PW86, Atk82]. Failure [Che85, CJ88, Des83, Hou86a, Hou88a, Koc81, Lou81, PW80a, PWP81, Pre82, Tsa88, Wil83, Pre84b]. Familial [MN81, Sri84]. Families [AO81, Bas81a, ESP82, Har89, Pes84, RS80, RB86, AO83]. Family [CR81, CP89, Cox88, Gen87, GNK82, WGJJ89, Wor86, CP91]. Farlie [HK84b]. Fast [EH85, KR86, KA86, KA89, Lon87]. Favour [Gün87]. Few [HJ81]. Field [DF85, Dup80, Gre85, Wil86, ZH89]. Field-Plot [ZH89]. Fields [DK87a]. Filter [De 88a]. Filtering [KA84, KA86]. Finding [RS82]. Finite [Bre84, CS89a, God85, Hor82, KA87, Kot86, Kot88, LOdBP88, MS89, RP82, SSW89, Kot87]. First [AT84, BHS85, CL81, Kan87, Tsa86a]. First- [CL81]. First-Order [AT84, BHS85, Kan87, Tsa86a]. Fish [Lei88]. Fisher [Fen83, Fer81, FR82, God84a, Hos88, Jar84, Pal81, Smi80a]. Fit [Che84a, CS89c, Col89, Har89, Hir86, KS81, Kou80, Koz80, Li88, LS85c, Mar81, Mil81, Nai81, Oja81, SL80, Sch80, Tsi80, LU80, MB88]. Fitted [Jon83, PT80]. Fitting [DCS87, Kem84, Sti84]. Fixed [AK82, Jor87, KA83]. fold [Bur89]. Following [Kim88, RT88, Whi86]. Forensic [Gro80]. Form [BM80]. Forms [JW79, MMM82, JW81]. Formula [BN83, Bes89]. Foundations [Bar84, God85]. Fourier [Dur80a]. Fourth [Tan82]. Fractional [Hos81, LM86b]. Fragmentary [Wei81]. Frank [Gen87]. Free [BDK89, HD82, Koc81, Leu83a, Oja81, Pre84a, SO82, Smi80a]. Freedom [Cox84, Kai83]. Freeman [Law84]. Frequency [CL80, EN88, MP89, Mil84, PN85, RR88, TJ82, MB88]. Frequentist [Pet80a]. Front [Ano80d, Ano80e, Ano80f, Ano81e, Ano81f, Ano81g, Ano82d, Ano82e, Ano82f, Ano83d, Ano83e, Ano83f, Ano84d, Ano84e, Ano84f, Ano85d, Ano85e, Ano85f, Ano86d, Ano86e, Ano86f, Ano87e, Ano87f, Ano87g, Ano87h, Ano88e, Ano88f, Ano88g, Ano88h, Ano89e, Ano89f, Ano89g, Ano89h]. Full [BN86]. Function [Azz81, Bha80, Bur88, CF84, DA85, ESP82, EP83, FS80, HW83, KD87, Kou80, Kou82, Lju82, Min83, Mor80, MT81, NP87, PW80a, Pat85a, Phi82, SvW86, Tsa88, DA87, HW84a]. Functional [CM83, CM84, DCS87, Lee85, Owe88]. Functions [BMW86, CD86, Cro87, DCS88, Dav87c, DHR89, DF89, Gla80, Kuk88, Lia87a, Lin82, Mat81, Mat88a, Nag82, Ng80, RC85, Ren88, SCR86, TKK89, VM88, WR87, TKK91]. Further [GW80]. Future [AR72, AR81].

G. [Gle85]. Gain [IS86, Ken83]. Gamma [AW80, BB80, Ber81, D’E81, Kou86, SH88]. Gaussian [Fin90, Fin86, GD82, GU83, God84b, GT84b, HLP88, Lju82, Sic82, SS83b, Tam87, Tyl87, Wal88, WR87]. General [BF88, Cha82b, DF89, JM80, Ken83, MS89, Nur81, SSW89, Vec85, JM81]. Generalization [Ber88]. Generalizations [Flu87]. Generalized [AK87, Atk80, BC89, CS89a, FR82, GAR85c, Gol86, Gol89b, Gor80, Hos84, IJ85, Jør83, Kou86, LZ86, Mor87b, Por85b, PW80b, Sch87a, SCR86, SC87, TC89, TT88, Toy82b]. Generated [TJ82]. Generating [ESP82]. Geometric [Gol89a, MJ83]. Geometrical [iA82, iA83, AK82, Böh89]. Global [Dal84]. Good [Lam86]. Goodness [Che84a, CS89c, Col89, Hir86, KS81, Kou80, Koz80, Li88, Nai81, Oja81, SL80, Sch80, Tsi80, LU80, MB88]. Goodness-Of-Fit [CS89c, Che84a, Col89, Kou80, Koz80, Li88, Oja81, SL80, Sch80, Tsi80, LU80, MB88]. Graphical [Atk81, BE89, DDS89, EK83, FL89, Lag81, WL83]. Greatest [BB86]. Green [CH81b]. Group [DW82, DW80, GP82, Jen87, KD87, RT88, Sin87, TRT85, ZF89]. Grouped [Bur82, Pet84, SP82, Tur89]. Groups [Pet82, Sch88a, Sch89]. Growth [Bol88, Ver86, VV88]. Gumbel [HK84b, Kim85].

H [Sti80]. H. [Wil84]. Haenszel [LSG83, LSG79, Man85]. Hannan [Pos90, Pos87]. Hazard [DD87, KP81, KY83, Lag81, MF85, NRW84, SK86]. Hazards [DDS89, GS87, Kuk84, Lag88, MOL86, RC85, Sch80, Sch82, SS83a, Tsi81]. Hermite [PKL83]. Heterogeneity [Hou84, SC85, Tar85]. Heterogeneous [Hou84, Hou86b, Hou88b]. Heteroscedastic [CM84, DM88a]. Heteroscedasticity [CW83, Tsa86a]. Hettmansperger [Smi80a]. Hidden [TG80]. Hierarchical [Raf88, SV80]. High [Pea80, SM86]. High-Order [Pea80]. Higher [KS81, KR86, PGT86]. Highly [Sch84a, ST87]. Histogram [Tay87]. history [Hol81, Hol82, Hol86, Sti80, Sti84]. Histospline [Ban88]. Hommel [Fal89]. Homogeneity [BBS84, Chi80, EM84, Fir89, Har85b, Kee82, Kee84, LS85a, Lia87b, Nag84, O’N82, Shu82]. Hormone [Bol88]. Horvitz [Kot88, PS80]. Households [Bec81, BH83]. Hurst [DH87]. Hyperbolic [Blæ81]. Hypergeometric [Lev84]. Hypersphere [FL86]. Hypotheses [Kou80, Sas80, Tho85, Wol88, Zhe88]. Hypothesis [Azz84, Dav87b, DW80, Gün87, Ken86, Por87, SW89, Azz87].

I. [WR82]. Identical [BFR89]. Identically [Fin86, Fin90]. Identifiability [HH89]. Identification [CH81a, TT85]. Ignorable [SS84b]. Ignoring [Gof87]. Illustration [CRC89]. Implications [HLH88]. Importance [HS89]. Improper [CV89]. Improve [Win80]. Improved [CP89, CP91, Kim88, Mat81, Por85a, Por85b, Sim86, Wor82, FJ89, Law84]. Improvement [Sko86]. Improvements [HW84b, Lev84]. Improving [Hal82b]. Incident [Yan78, Yan80]. Inclusion [BS84, GNK82, SS81]. Incomplete [AP82, Bho84, CW81a, Che86, Ekb82, Gla80, HJ81, Pat85b, Pat83a, Pat83b, SE86, Sre89, Wei82, WJ85, Wei83]. Incomplete-Block [Che86]. Incorporating [GP82]. Increased [RW82]. Increasing [Des83, PW80a, SM89, Tsa88]. Indecision [O’H81]. Independence [KG87, Por87, Shi81b, MB88]. Independent [Fin86, ZLS85, Fin90]. Index [BB87, Li85, Rei83]. Individualized [BG84]. Induced [Kou86]. Inefficiency [BB81, TGH83]. Inequalities [Sas80]. Inequality [Fal89, Sha85, Wor82]. Infection [CS88a]. Infectious [AW80]. Infectiousness [Bec81, BH83]. Inference [AT84, iA82, BN86, BB89b, Böh89, Bro87a, Bro87b, CS89d, DiC88, Don87, FTW85, FR88, Fri80, Jør83, Lar83, Lia85, Mar84, Muk88, Oak86, OP84, PR87a, Raf88, RW83, RP82, SC84, SK85, SS84a, Ski88, SV80, SS84b, Wah80, WT85, Wu85, iA83, Bro88]. Inferences [Ros87, Slu82, Ros88a]. Influence [Cha87, Cri85, Joh85, KTK89, KA89, RC85, TC89, VS80, WS81]. Influential [Atk81]. Information [Ano80g, Ano81h, Ano82g, Ano83g, Ano84g, Ano85g, Ano86g, Ano87i, Ano87j, Ano88i, Ano88j, Ano89i, Ano89j, Atk80, BM86, Bro82a, Cur88, Fer81, God84a, Hos88, IS86, Jar84, JR80, Ken83, LM84, Lia83, Pal81, Pat85a, SJ81, Tay87]. Informative [HHP82, SS84b]. Inhomogeneous [Ber81]. Initial [GG81]. Innovation [AK85, PN85]. Inspection [CS85]. Integrals [Tan82, Wan87]. Intensity [Hal85]. Interaction [DW85, DGS87, FL89, Gar85b]. Interactive [KL83]. Interclass [Sri84, SK88]. intercropping [FR87]. interested [Sti80]. Intergroup [Kra81]. Interlaboratory [Roc83]. Interpoint [GW80]. Interpolation [BB87]. Interpret [Ren88]. Interpretation [Gro80, Hen86, Mak85]. Interval [AK82, CT80, CF85, CFR88, Dav87c, DiC84, IR85, KA83, Pre84a, Tab87, Wei82, Wei83, WSM89, Gar85a]. Intervals [DD87, Efr85, Hal82b, JT85, Mat88a, MSS87, Owe88, RT88, Upt86, Win80, Sie80]. Intervention [Abr80]. Intraclass [DK80, Shi81b]. Intrinsic [Kün87b]. Invariant [Koz82, VM88]. Inverse [AL84, Bha80, Hos80, PS82, Sic82, Hos82, PS84, Sär80]. Invertibility [BM80]. investigations [FR87]. Involving [CP81, Gla80]. Isotonic [NS88, SIK83]. Items [Lig88, FR87]. Iterated [HK84b]. Iteration [HM88]. Iterations [Jor87]. Iterative [Gol86, Gol89b].

J. [SD86]. Jackknife [Efr81, Fra80, FS84, HW84b, KF83, Par83]. Jackknifing [Jor87, SS89a]. Jacobi [Ber88]. James [Joh82, Joh80, STV84]. James-Type [STV84]. Johannsen [Blæ81]. Johnson [Whe80]. Joint [Fli85, Hil86, Tsi81]. Judgement [Tam88b]. Judges [Pet82].

Kalman [KA84]. Kaplan [Mau83]. Kendall [KG87]. Kernel [Azz81, Bow80, HWC87, Lau83, Tay89, Tut86]. Kernel-Based [Tut86]. Knowledge [Fer82]. Known [Jar84]. Kolmogorov [SL80]. Koziol [CH81b]. Kruskal [Fli85]. Kurtosis [BB86, GPT85, AG83].

Lag [Wal88]. Lagrangian [CFR88]. Large [BBS84, BB89b, BCE83, BE86, CS86b, CS88b, Dav85b, DK80, God84b, Jør83, KiA84, Mak81, Win80]. Large-Sample [BBS84, BB89b, God84b, Jør83, Mak81]. Largest [CT80]. Latencies [TMKG87]. Latent [AW80, EGG89, Gil84, HP85, Lau81, Sch88b]. Latin [Fir89]. Lattice [Kün87b, WR80a, Guy82]. Law [Kou86]. Layout [BP88b]. Layouts [Bai87]. Leading [TA83]. learning [Bur89]. learning-testing [Bur89]. Least [CC88, Fin80, GPT85, Gol86, Gol89b, Gre88, Jew85, Kem84, Kun85, Mau83, RZ87, Toy82b]. Least-Squares [CC88, Gol89b, Gre88, RZ87]. Left [SF89, TJW87, Tsa88]. Length [Böh89, DG82, Gol89a]. Level [BNH88, Ber87, Bol88, CL80, RW83, Ste85b, WT85]. Level-Error [BNH88]. Levels [Fal89]. Levinson [Fra88, Fra85, Ton88]. Life [BW81, BT83, DD87, HP75, Hou84, KL83, Kum87, HP76, HP80b]. Lifetime [Law86]. Likelihood [AD80a, AA84, AB82a, AK83, BN83, BN86, BNH88, Bas81a, BF88, Böh89, CS88a, Cor87, CP89, Cox84, CFR88, CRC89, CL81, Dav80, Dav86, De 88b, DiC88, DR89, DHR89, DF89, FP80, Fir87, FR89, FL89, Gab82, GLR81, GD82, GU83, God84b, Gro80, GT84b, Hal87, Har86, HP85, HA87, Hin80, Jen86, Jør83, Ken82b, KS89, Lar83, LOdBP88, Lia84, Lia87a, LS85b, Lju82, Lon87, Mal86, MR87, MM84, MW89, MF85, Mat88a, MJ83, MS86, Min83, MT80, NP87, O’B80, OP84, OV81, Owe88, PW80a, PTF89, Pat85a, Pea80, Por85a, Por85b, PW80b, RS82, SH80, Slu82, Smi85, Spr80, TGG89, WR87, Whi86, Won89, Wor83, Azz83, CP91, Dav90, FJ89, Ken82a, O’B82, SD86, WR82, Wil84]. Likelihood-Based [Hal87, Mat88a]. Likelihoods [Bur82, CS89b, HA87, Lub81]. Limit [AR72, HW84b, TJW87, AR81]. Limiting [Vas81]. Limits [Bas81a, Fuj80b].

Line [Dar80, Las82]. Linear [BB87, Bol88, CS89b, CH81a, Cha82a, CM83, CM84, Coo80, CF84, Cro87, Dav85a, Dav87c, DiC88, FL81, GAR85c, Gol80, Gol86, GG81, Gre85, HKM86, Hin85, IO80, Jey82, Joh80, Jør83, JL86, KS89, Kuo88, LG84, Leu87, LZ86, MN81, Mak81, MW89, Mar84, MMM82, Min83, Mor87b, Nag82, Naz87, Pal81, Per84, Por85b, PT87, Pre78, Pre83, Sas80, Sch87a, SW86, Sen82, SV80, SCR86, SC87, Swe84, Tam82, TC89, TT88, Toy82a, Wel85, Wer89, Won89, You82, Joh82, Sär80, Sti84]. Linear-Model [MN81]. Linearity [LST88]. Local [Cox80, Dal84, DG82, HP80a, McC84a, Rya81, Ton88]. Locally [Ede89, Gor80, Lia87b]. Location [Ber89, BGW82, CS89a, Che82, Che84b, GL88, HJ82, HS81, Kou82, Oja81, RZ87, RB86, Sir82, Wei81, Wei82, DG85, FH82, Jam87, LM86a, Wei83]. Location- [GL88]. Location-Scale [RB86]. Log [ABS84, AH89, BN86, Bro84, Dav87c, Fra84, Jen83, Jen86, JW79, Man85, Min83, Naz87, Pal81, Por85b, ST87, SH88, JW81]. Logarithm [PGT86]. Logistic [AS80, AA84, AB82a, BG84, BC88b, CL88, Dav85a, Fow87, Fra84, Joh85, KL87, Le84a, RRK87, SCR86, TJ82, Tsi80, SD86]. Logistic-Normal [AS80]. Logit [Dav85a, GPT85, MT83]. Longitudinal [LZ86, ZLS85]. Look [Fli85]. Loss [Bro82a, Lag88, LT87, Ruk88, SF81, Smi80b]. Lower [BB86]. Lowered [Cox82].

M [Roc90]. M. [WR82]. Mahalanobis [MK89, MK85a]. Major [Cla80]. Malin [Smi80a]. Mann [PS81]. Manpower [Vas81]. Mantel [LSG83, LSG79, Man85]. manuscript [Sti84]. Many [Lia87b, Mat88a, ML88b, SS82, Tib89]. Marginal [Fir89, Kee82, Kee84, TKK89, TKK91]. Mark [SF81]. Mark-Recapture [SF81]. Marked [SK81]. Markers [WGJJ89]. Markov [DK87b, DL86, LM84, SS85, SS89b, Swa83, Vas81, WJ80]. Markovian [CW81b, Sun86]. Masking [Atk86, BK87a]. Matched [GLR81, Gar85b, Ros87, Shu86, Wel87, Wil80, Wil83, WL80, Ros88a, WL84]. Matched-Pairs [Wel87]. Matching [EM81, EM84, Le84b, Ros88b, Sam81, Le86]. Matrices [BVJM84, BJV88, Gla86b, GT84b, MR87, RS82, Tyl83, Tyl84]. Matrix [Daw81, GU83, HS80, HP85, Jey82, Pal81, Sch88b, Toy82b]. Matrix-Variate [Daw81]. Matter [Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h]. Maximal [Zid84]. Maximization [BFR89]. Maximized [Pat85a]. Maximum [AD80a, AA84, AB82a, AR72, Azz83, BN83, Bas81a, BF88, CS88a, CL81, DF89, FL89, GD82, GU83, God84b, HM84, Joe87, Jør83, LOdBP88, LS85b, Lon87, Mal86, MM84, MJ83, OV81, PW80a, PTF89, PW80b, RS82, Smi85, Spr80, Whi86, AR81, SD86, Wil84, WR82]. Mean [AK86, BW81, CT80, DD87, FL83b, Gar87, HS81, HS87, HP75, Hur88, Lev84, Sas80, SJ81, STV84, TGG89, Toy82b, Upt86, Wel83, FL84, HP76, HP80b]. Mean-Shift [HS87]. Means [BK83, ML88a, Mat81, MM89, Sch81a, KR87]. Measure [Fer81, Ken83]. Measurement [Pre82, RGL85, Sch87a, Ste85a, SC87, Pre84b]. Measurement- [SC87]. Measurements [Afs83, AG87, CT80, Gea88, Kun85, Lig88, WJ85]. Measures [Dau80, DG82, FL81, FL82, Gat86, IS86, Joh85, KO88, Muk88, Pat86, Pet86, Ver88]. Median [DD87, DM87, FR82, JT85, LM85b, Rei81, Wel87]. Medians [Kuk88, Sie82]. Medical [CP81]. Meier [Mau83]. Method [Azz81, Bow80, Bow84, BE89, Che86, God80b, HK84a, ML88a, Mal86, Par83, Pre84a, She82, Sin87, SM89, Ste80]. Methodology [BG87]. Methods [AP82, BM86, DH88, DW80, Efr81, EGG89, FS84, GP82, HW84b, HS87, Hou84, KTK89, Lia84, Mar82a, MS86, PKL83, Pet83, RT88, Sic82, Bur89, MB88]. Midzuno [CA83, PS80]. minimal [FR87]. Minimax [MH85, Mur83]. Minimizing [HM84]. Minimum [DCS88, GW80, HS81, LSW84, Pat83b, LSW89]. Misclassification [ML88a]. Mises [SL84, Bro82b, KF83, LS85c, SL82, Ste82, Upt86]. Misidentification [DP88]. Missing [AK83, Bha89, Gof87, KA83, LS85b, Lju82, Mar80, TJ83, TJ89]. Missing-Data [TJ83, TJ89]. Misspecification [Gla88, HS81, Sol84, Sol86]. Misspecified [SK86]. Misspecifying [Lag88]. Mixed [FL89, GAR85c, Gol86, GN87, HR82, Krz83, LS85b, Lon87, Pet86, Pos87, HR83, Pos90]. Mixture [Bas81b, SF89]. Mixtures [ABS84, Bas81a, Böh89, DW85, KL83, Wes87, FR87]. Model [AT84, AA86, Atk80, ABH89, BDK89, BMW86, BB89b, Bol88, CA83, CW81b, CL81, DD87, De 88b, EH85, FK86, Flu87, Fra84, GAR85c, Gla86a, Gol86, HO81, HS81, HH89, HS87, HT89, IO80, JL86, KY83, KL87, KA84, Kot88, Kuk84, Lou81, Lüt82, MN81, Mal88, MW89, Mar81, MF85, McA83, Mor87b, Muk88, MS89, New80a, Oak86, OP84, PT80, Pre82, RW89, RGL85, Sch80, Sch82, SS83a, Ski88, SV80, Ste85b, SP82, SF89, Swe84, Tam84, Tam86, Tam87, TT88, Toy82a, Tsa86a, TT85, Tsi80, Tsi81, Vas81, Ver86, VV88, Wil86, WF82, Won89, Zam89, Zeg88, CH81b, Gle85, Pre84b]. Model-Based [Kot88, Ski88, Tam86]. Modelling [CJ88, HO81, LM86b, Li88, LH89, Ton82, Ton83a]. Models [Ahn88, ABS84, AA84, AK86, AK87, BBS84, BD87, Bro87b, CSL^{+}84, CL88, CV89, CP89, DM88b, De 88a, DiC88, DL86, EK83, FH83, FL89, Gil84, Gol87, Goo81, GL88, Gre85, GJ82, GT84b, Hin85, Hir82, Hou86b, IS86, Jey82, Jon83, Jør83, KS81, Kem84, Kie82, KA83, KA85, KA86, KA89, Kün87b, LM85a, Lag81, Lag88, Lau81, Law86, Le84a, Lee85, Leu87, Li85, LZ86, Lju86, Lon87, LST88, Mal86, MM84, Mar84, McL84, Mil81, Mil84, Min83, Mon84, MS88, Nai86, NRW84, Pal81, Per84, Pet84, Pet86, PG82, Pie85, Por85a, Por85b, PT83, PT87, Rei83, SD84, SC84, Sch87a, SW86, Sen82, She82, SH82, Sol84, SCR86, SC87, SK86, Taw88, TC89, TG80, Vec85, VRW86, WJ80, WL83, Wer89, Wil83, Yam81, Bro88, CP91]. models [FR87, Hou88b, LM86a, dBP81, SD86, Sol86]. Moderate [Dav80]. Moderating [Wer89]. Modification [AB82b, Dav85a, FR82, Moo80, PS80]. Modified [BGW82, Hom88, Hom89, JW79, LL83, Pos87, ST87, JW81, Pos90]. Modulated [Ber81]. Moment [ESP82, Moo86, O’H81, Sha85]. Moments [Dav87c, DA85, DA87, Jar84, LSW84, LSW89]. Monitoring [DW82]. Monotone [Böh89, Las82]. Monotonic [Bri89, WR80b]. Monte [BC89]. Mood [FR82]. Moran [CS89c]. Morgenstern [D’E81, HK84b]. Morgenstern-Type [D’E81]. Mortality [BNJL89, Lei88]. Most [Ede89, Gor80, Lia87b]. Motion [PS85]. Moving [AK83, AK85, CW81b, CL81, Fin86, Fra85, God80b, GD82, GU83, God84b, Hal89a, HLP88, HR82, HK84a, Hos80, KA84, KA86, Lju82, Mas87, McL84, Mon84, New80b, Pea80, PT80, Pos87, SD84, TG80, DR83, Fin90, Fra88, Hal90, HR83, Hos82, LM86a, Pos90, Yam81]. MR [AR81, BR81, CS82, Hol82, HP80b, Hos82, Joh82, JW81, JM81, O’B82, dBP81]. MR0497517 [Pre83]. MR0570503 [Fuj86]. MR0601097 [WL84]. MR0626394 [AO83]. MR0637784 [WR82]. MR0655666 [iA83]. MR0655673 [HR83]. MR0655691 [Wei83]. MR0655696 [Ton83a]. MR0655698 [PS84]. MR0671971 [Pre84b]. MR0695210 [SL84]. MR0712022 [FL84]. MR0712026 [TP84]. MR0712037 [HW84a]. MR0712046 [MP84]. MR0725375 [TJ89]. MR0738329 [Azz87]. MR0738342 [Le86]. MR0743002 [SS85]. MR0767157 [Sol86]. MR0775397 [LSW89]. MR0817571 [Fra88]. MR0817575 [Sch87b]. MR0817580 [Har87]. MR0817585 [DA87]. MR0836445 [Roc90]. MR0855887 [Mar88]. MR0855892 [Dav90]. MR0855898 [Hou88b]. MR0855914 [Fin90]. MR0885915 [Ros88a]. MR0903131 [Mat88b]. MR0903138 [Bro88]. MR0909368 [WFBH88]. MR0919846 [Pos90]. MR0946048 [Sch89]. MR0991422 [Hal90]. MR0991426 [CP91]. MR1040637 [TKK91]. Multidimensional [EH85, SL80]. Multilevel [Gol86, Gol87]. Multimodality [MW89]. Multinomial [BK87b, MJ83, SC84]. multinomials [Law84]. Multiple [Abr80, Cas81, CW80, EM81, EM84, HLH88, Hoc88, Hom88, Le84b, LH89, Lüt82, NS88, Ros88b, Sim86, VRW86, Le86]. Multiple-Recapture [Cas81]. Multiplicative [McL84, Var89]. Multiply [Gar85b]. Multiply-Matched [Gar85b]. Multiprocess [Bol88]. Multiresponse [MBBW81, Ste87]. Multistage [RL84]. Multistate [HCT83]. Multivariate [Ahn88, AH89, BS89, CM83, CF85, Dav80, Gre88, GT84b, Hai87, Hal81a, Hou86a, Hou88a, Kie82, Kon81, Koz82, Li85, Mac83, MT80, MW80, New80b, O’B80, PWP81, Rei83, Roy89, Sas80, SIK83, Ski83, Wol88, Yam81, O’B82]. Mutual [Por87]. Mutually [Bar83].

Natural [Har89, Lei88]. Near [Hal81b, Mar84]. Near-Neighbour [Mar84]. Nearest [GS85, You82]. Nearly [AT84, CW81a, SO82]. Negative [BP88b, Chi80]. Neighbour [GS85, Hal81b, Kun87a, Mar84, WW87, Wil86, You82]. Nested [AP82, Che86, HS89, IJ85, Lon87, Mor87b, Sre89]. Newly [TE87]. Newly-Discovered [TE87]. Neyman [RB86]. Neyman-Type [RB86]. No [GTP88, iA83, AO83, Azz87, BR81, Bro88, CP91, CS82, Dav90, DA87, Fin90, FL84, Fra88, Fuj86, HW84a, Hal90, HR83, Har87, Hol82, HP80b, Hos82, Hou88b, Joh82, JW81, JM81, Kot87, Le86, LSW89, Mar88, Mat88b, MP84, O’B82, dBP81, Pos90, PS84, Pre83, Pre84b, Roc90, Ros88a, Sch87b, SS85, Sch89, SL84, Sol86, TKK91, TJ89, TP84, Ton83a, Tyl84, WR82, Wei83, WFBH88, WL84]. Noise [Fin80, KR86, Res89]. Noisy [MSS87, Rit89, TMKG87]. Non [Fin86, HLP88, Wal88, Fin90]. Non-Gaussian [Fin86, HLP88, Wal88, Fin90]. Nonadditivity [Kee85]. Nonadditivity-Type [Kee85]. Noncentral [CF84, Kou86, Lev84]. Nonconsecutively [RW87]. Nonconvergence [Mur83]. Nonequivalence [HM85]. Nonhomogeneous [Vas81]. Noninformative [Ste87, Tib89]. Noniterative [Nur81, TGH83]. Nonlinear [CT85, CTW86, Coo87, CP89, Eav83, FS80, GWP84, GSJS86, GH88, HO81, Ham86, HW87, Jør83, Mor87a, TKK89, Wu85, ZH89, CP91, TKK91]. Nonlinearity [Kee85, PD86, Tsa86b]. Nonnegative [Hil86, TGG89]. Nonnested [Ken86, McA83]. Nonnormal [Min83, MW80]. Nonnormality [Dav80, Fuj80a, VS80, Fuj86]. Nonparametric [ABH89, BHT84, Bre82, Bre84, BES88, BE89, Cam81, DP88, DGS87, DF89, Efr81, FL81, FL82, GG80, Hal81a, HW88, HH88, HM85, Jup87, LBD88, Lou81, MBBW81, Sch81b, SM86, Var89, WGJJ89, Zhe88, Ede86]. Nonregular [BF88, Smi85]. Nonresolvable [HJ81]. Nonresponse [SS84a]. Nonstandard [BF88]. Nonstationary [AT84, KA87]. Normal [AS80, AH89, BK87a, CF85, DG89, Goo81, GG80, Hal82b, IS86, ML88b, Mor80, MT80, Nag82, Per84, Pet86, Sas80, Sch84a, SG85, TGG89, Wan87, Wes87, AG83]. Normality [EP83, HW83, Koz82, LM80, Mac83, MI81, Oja83, Per81, PG82, Pie85, Spi80, HW84a]. Normalizing [Kon81]. Notation [McC84b]. Notational [Daw81]. Note [AL84, AA86, Atk80, Azz81, BB89a, BNJL89, BB80, Bow80, Cox88, DL86, Far82, HS80, Har86, Jen83, Kha81, KA84, Kum87, Law86, LOdBP88, Mar81, Mau83, MP83, Mon84, Nag82, Par83, Por85b, SV80, Tam82, Ton83b, Ton88, TJW87, Tsi80, Van80, WR80a, You88, Gle85, MP84, SD86, Sti84, Wil84]. Notion [CS89a]. Nuisance [Dav87b, God80a, God84a, HA87, KiA84, Lia83, Rit89, SM89, Spr80]. Null [Ali84, Gün87, Har85a, SK85, Wes88, Har87]. Number [Dav85b, KiA84, Shi84]. Numbers [CS86b, CS88b, GN87, Jac80].

Observational [RR83, Ros87, Ros88a]. Observations [AR72, Atk81, Bha89, Con89, DH85, GS85, Joe87, Kun85, Leu83a, Lju82, Mar80, PW80a, Pet80b, Pet86, TA83, WJ80, AR81]. Observed [Dav85b, Gei81, Pal81, RW87]. Obtaining [JR81]. Odds [Bre81, Bro87a, Dav85b, Dav87c, Don87, HT81a, Lia85, LS85a, PT83, TGH83, WH89]. Odds-Ratio [WH89]. Olkin [WR82]. Omitted [GWP84]. Omnibus [Spi80]. One [Bai87, BP88b, DW80, DF85, Hal82b, OV81, Pet80b, RW82, Tib89, Wyn84, Law84, SH82, WR82]. One- [DF85, SH82]. One-Dimensional [Wyn84]. One-Sided [DW80, Hal82b]. One-Way [Bai87, BP88b, Law84]. Only [Dav87b]. Onset [Por86]. Optimal [AD80b, AG87, Bel84, Bro87a, CT80, Coo80, FK86, GV83, Hal81b, Kun85, Mat87, Mat88b, MH88, MS89, Sch84b, Shi81a, SE86, SC87, Tam84, Tho85, GS89, Jac80, Shi82]. Optimality [CL87, Lin82, Tam87, Yeh86]. Optimally [SCR86]. Optimization [Jon83, Mar82a]. Optimum [Atk82, AD89]. Order [AT84, BHS85, BS84, CL81, DG82, DR83, Dea80, EN88, Fra80, God80b, HR82, Hor82, Hos88, Kan87, KS81, KR86, MH85, Muk88, Mur83, Pea80, Pos87, RW83, SD84, SK85, SW87, Sko86, Tan82, Toy82b, Tsa86a, WT85, Azz83, HR83, Pos90]. Ordered [Bar83, Dal84, Goo81, Hir82, KY82, MM89, Nag82, Nai86, Pet84, RW82, Wah80]. Ordering [BDK89, DF89, WT85]. Ordinal [Far82, LM85a]. Ordinary [LM86a, Bur89]. Origin [SF89, VM88]. Origin-Invariant [VM88]. Orthant [Hor82]. Orthogonal [GN87]. Orthogonality [Ton88]. Other [Böh89, Bre84, Efr81, MW80, ZH89, FR87, McL84]. Outlier [CB88, Roc86, Roc90, Tam82, DG85]. Outliers [BK87a, CH81a, JL86]. Outlying [Atk81]. Overdispersed [Moo86]. Overdispersion [Cox83]. Overlapping [MV82, DR83].

Page [Kha81]. Paired [CP81, Dav87a, DP88, Ekb82, LL83, MB89, SO82, WGJJ89]. Paired-Comparison [Dav87a]. Pairs [Wel87, Wil80, WL80, WL84]. Pairwise [DGS87, Fli85]. Pairwise-Interaction [DGS87]. Panel [Sun86]. Papadakis [DF85, ZH89]. paper [WR82]. Parallel [Dar80]. Parameter [AT84, BB86, BGW82, CS89a, Col89, Coo87, DK87a, Dav87b, DiC84, Dup80, FK86, FR88, Gla86a, God80a, God84a, Guy82, Ham86, Hos84, KY83, KiA84, Lia83, Pre82, Raf88, RW87, Ruk88, Ste85a, Ste87, Tay89, Tib89, Ton88, Pre84b]. Parameterization [Ahn88]. Parameters [AK86, BN86, BG84, Che82, Che84b, CS89c, CW81b, GD82, Hal82b, HS81, HA87, JR81, Kan87, KiA84, LG84, Mas87, Mat88a, Mat89, PW82, PW80b, Rit89, SM89, Spr80, SP82, Toy82a, Toy82b, Whe80]. Parametric [DHR89, Efr85, ESSAH89, FS80, Ng80, Pes84, Pet82]. Partial [AB82b, BN86, Dau80, Dur80a, FP80, Sch82, Slu82, ZF89]. Partially [CW80, DL86, Gei81, LR81, Sun86]. Particular [CA83]. Pattern [GSJS86, Hir83]. Patterned [RS82]. Peak [EN88]. Peaks [LT87]. Pearson [LSW89, BT81, LSW84, SG85]. Penalized [AB82a]. Percentage [HS80]. Period [Pat85b]. Periodic [TG80]. Periods [AW80]. Permutation [Bro82b, Jen83, MBBW81, Ros87, Wei88, Ros88a]. Perturbations [Bol88]. Petersen [SF81]. Phase [ST89, Wal88]. Phase-Lag [Wal88]. Pivotal [CS89b, Hin80]. Pivotals [CS89b]. Pivots [But89, Mor81a]. Plackett [Wah80]. Plan [Cha82b, Sen89, Zhe88]. Plans [Yeh88]. Play [Wei88]. Play-the-Winner [Wei88]. Plot [Col89, Kim85, Mic83, ZH89]. Plots [Coo87, DG89, FS85, Nai81, SS82]. Plus [Fin80]. Poem [TE87]. Point [DGS87, JJS87, Joh85, KS89, LKP86, MF85, NRW84, Pet80b, Pet81, RA86, Wor86]. Points [HS80, Jor87, LM85b]. Poisson [AH89, BLE88, Dup80, Hal82b, LKP86, Mor87b, RA86, SC84, Sic82]. Poisson-Log [AH89]. Polychotomous [BG84]. Polynomial [Dea80, Wyn84]. Polynomials [Lüt82, McC84b]. Population [Ano81a, BK87b, BO78, Cas81, DP88, HP85, Kot86, Kot88, LOdBP88, MS89, Nay88, PR87a, RP82, SSW89, SJ81, Kot87]. Populations [DDS89, Hou84, Hou86b, Krz83, Lam86, MW80, O’N82, SK81, Hou88b]. Portfolio [PT87]. Portion [Gei81]. Portmanteau [PD86]. Positive [KG87]. Possibly [Cox82, SK81]. Posterior [Pet81, PT83]. Potency [Kuo88]. Potential [TMKG87]. Power [Ali84, CR81, HP80a, Pes84, SW87, SvW86, Tay85, Wes88, Wor83]. Powerful [Ede89, Gor80, Lia87b]. Precision [Pal87, RW82]. Prediction [AK86, Bes89, BS89, CR81, Gei81, IO80, Kem84, KA85, Ski83, Toy82a]. Predictions [Yam81]. Predictive [AJ89, AD80b, But89, Dav86, Har89, Lar83, Dav90]. Predictor [Tam87]. Predictors [Tam86, Tam88a]. Prepivoting [Ber87]. Preposterior [BW81]. Presence [BR75, Fer82, God80a, God84a, Hal89a, Jen84, KiA84, Lia83, Mat87, Spr80, BR81, Hal90, Mat88b]. Present [Bel84, Dav87b]. Prevalent [BG87]. Prevention [Pre86]. Primary [GN87]. Principal [Ait83, Cri85, DM85, Flu87, Sch88a, Sch89]. Principle [Lar83]. Prior [BM86, Fer82, Gol80, Ste87, Swe84]. Priors [CV89, Tib89]. Probabilistic [BM80]. Probabilities [Ano81a, Bar83, BS84, BHT84, BO78, Hor82, Kuk88, Lau83, MJ83, Pal87, Pet81, RW83, SK81, SS81, WT85]. Probability [AA86, BHT84, Cha82b, Col89, DP88, Dri88, Gab81, Gab84, GNK82, Kim85, Kok81, Lig88, Mic83, SA81, SSS86, Sen89, ST89, Wan87, Wor83, Hol81, Hol82, Hol86, Sär80, Sti80, Sti84]. Probit [Kie82, OP84]. Problem [Fen83, FR82, HLH88, HT81a, MS86, NS88, PW82, Pet80b, RW89]. Problems [Bre84, ĆGKP82, Efr85, HHP82, JM80, Lom87, Lun85, TJ83, WR87, JM81, TJ89]. Procedure [AB82b, AR88, EH85, Gab81, Hoc88, Hom88, Kha81, Lam86, Per84, RW89, SA81, Shi84, Sim86, SS81]. Procedures [Don87, FD83, Gei81, HF82, Hom89, Kem84, Koz82, MMM82, MBBW81, MW80, Tho85, TJ83, TJ89]. Process [AD80a, AK83, AK85, CJ88, God80b, GD82, GU83, God84b, Gof87, GG81, Hos80, Lju82, Pea80, PS85, RA86, Ton83b, WJ80, Guy82, Hos82]. Processes [BB80, Bas81b, BHS85, Ber81, DK87b, DGS87, Fra85, God85, Las82, SS83b, Ton82, Var89, Vec85, WR87, Azz83, Fra88, KR87, Ton83a]. Produce [BFR89]. Product [AJ89, RS80, Shu86, TJW87]. Product-Limit [TJW87]. Product-Type [RS80]. Products [MP83, MP84]. Profile [CFR88, FR89]. Prognostic [Atk82]. Programming [MJ83]. Projection [SM89]. proof [Wil84]. Propensity [RR83]. proper [LM86a]. Properties [AS80, Bur82, CL81, Ken82b, LSG79, LSG83, McL84, Moo86, Pet80a, Sch81b, Tyl83, Ken82a, Tyl84]. Proportion [FS84]. Proportional [DD87, GS87, GNK82, KY83, Kuk84, Lag81, Lag88, MOL86, RC85, Sch80, Sch82, SS83a, SS81, ST89, SK86, Tsi81]. Proportions [Moo86, SS89b, Ste82]. Propriety [Man85]. Providing [SS80]. Pure [GG81]. Purpose [Cha82b].

Quadrant [KG87]. Quadratic [Cro87, Min83, WF82]. Qualitative [Dau80]. Quantal [MH80]. Quantile [Bro81, HD82, Whe80]. Quantiles [Azz81, Joe87, MT83, BC88a]. Quasi [Fir87, NP87]. Quasi-Likelihood [Fir87, NP87]. Quasisymmetry [McC82].

R [Gle85]. R. [Sti80]. Radial [Tyl82]. Random [Bur88, Coo80, CF84, CH81b, DK87a, Fin86, FL86, HO81, Jup87, Kum87, Leu87, Lev84, Lon87, Mal86, Nag82, Van80, Vec85, Wor86, Yan78, Yan80, Fin90, Hol86, KR87]. Randomization [Bai83, Bel84, Bel86, MPW88, SW83, WSM89]. Randomized [GWP84, GTP88, Sch85, Wei88, WL81, Sch87b]. Randomly [Cam81, Cha87, Che84a, Koz80, Mar80, Nai81]. Randomness [GG80]. Range [CS89a, Che82, Che84b]. Rank [Bro84, BE86, Cuz82, Gor80, HF82, Jen83, JW79, LL83, Leu83b, Lom87, Man85, MMM82, Pre78, Sch85, ST87, Shi81b, SH88, VRW86, WL80, WL81, WL84, JW81, Mar84, Pre83, Sch87b]. Ranking [Dav87a, Fli85, Mak85]. Ranks [Hen86, Pet81, Pet83]. Rao [Mar82b]. Rare [CS79, Tam88b, CS82]. Rate [AK85, CS88a, Des83, KD87, MF85, NRW84, PW80a, PTF89, Tsa88]. Rates [Che85, Koc81, Sch81a]. Ratio [BN86, BNH88, Bre81, Bro87a, CA83, Cor87, CP89, Cox84, Dav80, Dav87c, DR89, Don87, Gro80, GT84b, Har86, HP85, HT81a, Jen86, KP81, Ken82b, KS89, LG84, Lia85, LS85a, MR87, MT80, Owe88, PW82, Por85a, Por85b, RS80, SH80, Sri80, TGG89, TGH83, Wes88, WH89, Wor83, Wu82, CP91, FJ89, Ken82a]. Ratio-Type [CA83]. Ratios [Dav85b, HW87]. Rayleigh [Moo80]. Real [Bol88, FR88]. really [FJ89]. Recapture [Cas81, Nay88, SC84, SF81]. Record [Fin80]. Recursion [Fra85, Fra88]. Recursive [HR82, HR83, HKM86, SM86, TJ83, TJ89, WL83]. Reduce [Ber87, Dav85a]. Reduced [Cla80, VRW86]. Redundancy [BFR89]. Reference [KR86, Pat85b]. Region [DM88a]. Regions [Hal87, Ham86, Krz89, Wor86]. Regression [AA84, AA86, AB82a, Atk81, AL89, ABH89, BP88a, BB89b, BG84, BC88b, BES88, BE89, CSL^{+}84, CH81a, CL88, CV89, CW83, CT85, CTW86, Coo87, Dau80, Dav85a, Dea80, DiC88, Dur80a, Eav83, Far82, Fow87, GSJS86, Gla86a, Gol80, GL88, Hai87, Ham86, HW87, HM85, Hil86, Hin88, HT89, Jew85, Joh80, Joh85, Jør83, JM80, KY83, KL87, Kem84, KS89, KA85, Kuk84, Lag81, Lag88, Law86, Law87, Mal86, MM84, Mar84, Mat81, MH82, Mor87a, MS88, NS88, OP84, Pet83, PP88, PG82, PWP81, Pre82, QJ86, RW89, RC85, RRK87, SSW89, SIK83, Sch80, Sch82, Shi81a, Shi84, Sie82, SS80, SS84a, SH88, Sol84, SCR86, SP82, Tam84, Tam82, TC89, Tsi80, VS80, Won89, Wyn84, Zeg88, Joh82, JM81, dBP81, Pre84b, SD86]. regression [Shi82, Sol86]. Regressions [BG84, DM88a, GWP84]. Regular [Sin87]. Rejective [Hom88]. Related [Blæ81, Bro81, Bro87b, JM80, Kün87b, WJ80, Bro88, JM81, Sie80]. Relation [Yan78, Yan80, Sti84]. Relationship [CM84, Don87, Kee82, Kee84, MMM82, Mor81b]. Relationships [Cha82a, CM83, DCS87, LG84, SM86]. Relative [BP88a, CA83, EM81, EM84, Hil81, Le84b, Nur81, TGH83, TT88, Tur89, VM88, Gar85a, Le86]. Relevance [BM80]. Reliability [Win80]. Remarks [CV89, Cox83, SS85]. Renewal [Var89, KR87]. Repeated [Afs83, AG87, CT80, FM82, Gea88, JT85, Kun85, Muk88, Pat86, Pet86, Sie82, Ver88, WJ85, Bur89]. Repeats [WR80a]. Replacement [BS84, Gab81, Gab84]. Replicate [Lig88]. Replicated [DCS87]. Replicates [HJ81, Jac80]. Replication [CC88]. Replications [KY82]. Representation [AK85, BZ88, CW81b, God84b]. Representations [Fin86, HJ82, HLP88, Fin90]. Rerandomizing [Wel87]. Resampling [DH88, HM88, Hal89b]. Residual [Ahn88, BES88, BE89, CB88, GSJS86, HP75, Joh80, Li85, MM84, SSW89, HP76, HP80b, Joh82]. Residuals [BP88a, CT85, DG89, Dur80a, Sch82, SH88, Tam82]. Resilient [Gla88]. Resistance [Roc86, Roc90]. Resistant [Bes81]. Resolutions [BN80]. Resolvable [MK85b]. Response [AO81, AE83, AD89, Cox88, FR87, GJ82, MH85, SM86, Ste85b, AO83]. Responses [Dal84, HM84]. Restricted [Bai83, Gol89b, MPW88, Muk88, RW83, SK85, SW87, SW83, WSM89, WT85]. Restrictions [Dea80, Hil86]. Result [Bes89]. Results [BFR89, Kot86, Lin82, Mak81, Rei83, Kot87]. Retrospective [Hal80]. Reuse [Gei81]. Reversibility [HLP88]. Ridge [Hai87]. Right [Cuz82, IR85, Jen84, PW82, Pre78, TJW87, Tsa88, Pre83]. Right-Censored [IR85]. Risk [BP88a, EM81, EM84, Le84b, Nur81, TGH83, VM88, Gar85a, Le86]. Risks [BDK89, HH89, SR83]. Rissanen [Pos90, Pos87]. Robin [Rus80]. Robust [BHS85, BKB89, Con89, GT84a, HJ82, Ken82a, Ken82b, Li88, LH89, Mar82b, Mas87, MH80, Roc83, RP82, SH80, SH82, Sie82, Ste85b, Zam89]. Robustness [Bro87b, Bro88, Tyl83, Tyl84, VS80, Wes88, Sti80]. Rocke [Roc90]. Rodent [BKB89, RO88, RO89]. Role [Cox82, DF85, RR83]. Root [DR89, Hal89a, PP88, Hal90]. Roots [Fuj80a, HS80, HP85, SD84, Sch88b, Fuj86]. Rotating [ME82]. Round [Rus80]. Row [IJ85, JE86]. Row-Column [IJ85, JE86]. Rows [AP82, Che86, Sre89]. Royall [Tam87]. Rule [Wei88]. Rules [Bin81, MPW88, Pet80a, Rob83, WSM89]. Run [Gol89a].

Sacrifice [BT83]. Saddlepoint [Dan80, Dan83, DH88, Feu89]. Sampford [Gab81]. Sample [BBS84, BB89b, BLS^{+}82, Bre84, CS86a, Cla80, CL81, DK80, FR88, Fuj80a, Gei81, God85, God84b, Hal80, Hos80, Jør83, Jup87, LSG79, LSG83, Mak81, MR87, PW82, PS81, RRK87, SH82, SS84a, Sri80, ST89, Tab87, Wei88, Win80, Yan78, Yan80, FH82, Fuj86, Hos82, KR87]. Sampled [DA85, DA87]. Samples [BK87a, CP80, CS86b, CS88b, HT89, Jew85, Kam89, Lig89, QJ86, RZ87, RW82, Roc86, RP82, SO82, Sch84a, SL82, Ski83, Spi80, Spr80, Wei81, AG83, DR83, Jam87, Law84, Roc90, SL84]. Sampling [BR75, BS84, BD87, Bri80, Cas81, Cha82b, CL87, CS79, CS85, Gab81, Gab84, GNK82, HS89, HHP82, Kot86, Kuk88, Nay88, Pal87, Por87, PS80, PS82, RP82, SA81, SSS86, Sen89, SS80, Ski88, SJ81, SS81, ST89, SS84b, Tam84, Tam86, Tam88a, Yeh88, BR81, CS82, Kot87, KR87, PS84, Sär80]. Scale [CS89a, GL88, HS81, Kou82, MI81, Oja81, PW82, RB86, SH82, Wes87]. Scale-Free [Oja81]. Scaling [Ste85b]. Scatter [Tyl83, Tyl84]. Scheme [PS80, SSS86]. Schemes [GNK82, Por87, SS80]. Schoenfeld [MOL86]. Score [BKB89, Har85a, Hin88, Law87, Lin82, RR83, SCR86, TT88, Tsa86a, Har87]. Scores [DG89, HP80a, Hir82, SC87, Tar85, Tsi81]. Scoring [Lon87]. Screening [BD87, RO88, RO89]. Search [EH85]. Seasonal [FH83, KA84, KA86, McL84]. Second [BS84, Fra80, Hos88, MH85, Toy82b]. Second- [MH85]. Second-Order [BS84, Fra80, Hos88, Toy82b]. Segment [Hal85, Las82]. Selected [Ski83]. Selecting [Cur88, Lam86]. Selection [Dri88, HHP82, HT89, Kem84, Por85a, RW89, Shi81a, Shi84, ST89, TSE88, Shi82]. Selections [GN87]. Selectors [HM85]. Self [PD86]. Self-Exciting [PD86]. Semi [DK87b, DL86]. Semi-Markov [DK87b, DL86]. Semiparametric [BNJL89, Oak86]. Sen [CA83]. Sensitivity [Cox82, KTK89, LG84, Ros87, Ros88a, Ros88b, SM89]. Separate [ESP82, Pes84, dBP81]. Sequence [Wor86]. Sequential [AB82b, AR88, Atk82, Bau86, BT83, BS84, CP81, CW80, DW82, DW80, Ede89, FTW85, Gea88, GP82, HT81b, Jen87, JW79, JW81, KD87, Kim88, LD83, LF80, Meh81, NH80, Rob83, RT88, SS83a, Sie80, Ste80, Swa83, Tab87, TRT85, Whi86, Wu85, Zhe88, SS85]. Sequentially [FS80]. Serial [Ber89, BT83, CS88b, Dur80a, O’B80, Roy89, TA83, O’B82]. Serially [Con89, Gla86b, MB89]. Serially-Structured [Gla86b]. Series [Abr80, Azz84, BZ88, BBS84, BHS85, Bri89, CMH87, CRC89, De 88a, Dur80a, Fin86, FH83, HO81, HLP88, Hur88, HT89, KS81, Kee85, KA85, KA87, LM86b, Li88, LH89, Lju86, Lüt82, Mar80, MB89, Mil81, Mil84, New80a, PD86, PP88, PT83, PT87, PN85, RW87, Roy89, TG80, TP83, Ton82, TT85, Tsa86b, VRW86, Wal88, Zeg88, Azz87, Fin90, TP84, Ton83a]. Set [Yan78, Yan80]. Sets [Ber87, BCE83, BE86]. Several [CA83, Che82, Che84b, FL83b, Lig88, Lig89, Nag80, WH89, FL84]. Shakespeare [TE87]. Shape [DM85, Hos84, MD89, MT81, Spi83]. Sharpe [BT81]. Sharper [Hoc88, PS82, PS84]. Shift [HS87]. Sided [DW80, Hal82b, Kha81]. Sighting [SK81]. Signal [Fin80, KA87, KA89, Rit89]. Signals [TMKG87]. Signed [BN86, DR89]. Significance [FM82, Hoc88, JR80, MPW88, Sim86, SW83, Tho85, BC89]. Similar [Jen86]. Similarities [Gil84]. Simple [CS88a, Dav85a, GS87, KS89, Kou80, Lev84, LM80, Mar84, Pet80b, SA81, Sch84a, Tab87, Tho85, Won89, WT85]. Simplified [RL84]. Simpson [Sti84, Zid84]. Simpson-disaggregations [Zid84]. Simulation [DHS86, Hal89c]. Simultaneous [NP84]. Simultaneously [SS82]. Single [Owe88, Yeh88, KR87]. Singularity [MF85]. Sister [ĆGKP82]. Situation [Wu85]. Size [Ano81a, BLS^{+}82, BK87b, BO78, GG81, GNK82, Hal80, LOdBP88, Nay88, PR87a, SS81, ST89, Yan78, Yan80]. Sizes [FS84, GJ83, Jen87]. Skewness [GPT85]. Slippage [SW87]. Slope [CH81a, IR85, MH85]. Slowly [Ber89]. Small [CP80, CS86b, CS88b, CL81, FH82, HT89, LSG79, Lig89, Roc86, Spi80, Spr80, LU80, Law84, Roc90, LSG83]. Small-Sample [CL81, FH82]. Smallest [Sch88b]. Smith [Sti80]. Smooth [LST88, MS88, RB86, SY87, WV81]. Smoothed [DM88b]. Smoothing [AK82, Ban88, Bow84, Fow87, Gre85, Tay89, Tut86]. Sobolev [HR85]. Sojourn [DK87b]. Some [AP82, AS80, AG87, Bur82, CV89, Cox83, Daw81, Fow87, Gab82, Gil84, Gla86b, Gor80, HLH88, HM85, HCT83, Jen84, KA83, Kon81, Kot86, Kot87, LM89, Lin82, Mar81, Mar82c, McC82, PKL83, Rei83, SK85, Sch84b, SW87, Sre89, TP83, TP84]. Space [AK86, AK87, DA85, DM88b, De 88b, FR88, HJ82, KA83, KA89, DA87]. Space-Time [DA85, DA87]. Spaced [Fal89]. Spacings [DG82]. Sparse [Bre81, LS85a]. Spatial [DM87, MM84, MW89, Mar81, Mar86, SS83b, WR87, Mar88]. Special [KR86]. Specification [PT80]. Specified [LM85b]. Spectra [NP84]. Spectral [EN88, LT87, SvW86, Tan82]. Spending [KD87]. Sphere [FL86, GW80, Tyl87]. Spherical [FL83b, HWC87, FL84]. Sphericity [Tyl82]. Spline [MJ89]. Splines [WR80b]. Spread [AW80]. Squared [AK86, Bed83, HS81, Hir86, Hur88, Kou86, LSW84, Sch80, TA83, Toy82b, Wel83, LSW89]. Squares [CC88, Fin80, Fir89, GPT85, Gol86, Gol89b, Gre88, Jew85, Joh80, Kem84, Kun85, Mau83, RZ87, Toy82b, Joh82]. Stabilization [Tib88]. Stabilized [Col89, Kim85, Mic83]. Stabilizing [BLE88]. Stable [BB80, Hou86b, Hou88b]. Stage [BC88b, TSE88]. Stagewise [Hom88]. Standard [Efr81, Gla88, Per84]. Standardized [BN86, Jen86]. State [AK86, AK87, DM88b, De 88b, KA83, KA89]. States [Sun86]. Stationarity [Mon84]. Stationary [Bri89, DK87a, Fin86, HLP88, Lju82, PN85, Vec85, Azz83, Fin90, Guy82]. Statistic [Ali84, BNH88, Bha89, CS89a, Che82, Che84b, CS89c, DR89, Fli85, Gla80, GT84b, HP80a, Har85a, Hir86, Jen86, Law87, LSW84, MT80, Pet80b, PS81, SK85, SL82, You82, AG83, Fre81, FJ89, Har87, LU80, LSW89, SL84]. Statistical [Bes81, Bro87b, ĆGKP82, Hug89, Jon83, Lau81, MD89, Roc83, SC84, Tyl87, Bro88]. Statistics [AL89, Cor87, CP89, DG82, Hor82, HCT83, JR80, Kon81, LM89, Lun85, Mac83, Mar84, Por85a, Por85b, Sha85, TA83, WGJJ89, CP91, DR83, Hol81, Hol82, Hol86, Law84, Sti80, Sti84]. Step [Böh89]. Step-Length [Böh89]. Stimuli [KL83]. Stochastic [BDK89, CS88a, DH85, DF89, God85, HH88, Azz83]. Stopping [Pet80a].

Strata [Lia87b, Mor87b]. Strategy [CA83, Pos87, Pos90]. Stratified [Jew85, Kot86, QJ86, RZ87, ST87, Kot87]. Stratum [GN87]. Strengths [Leu83b]. Strong [CFR88]. Strongly [RW89]. Structural [Cha82a, KiA84, LG84, Lee85, RGL85]. Structure [HK84b, Ken86, O’N82, Roy89, Sun86]. Structured [Ahn88, Gla86b]. Structures [Sha85]. Studentized [MT81]. Studies [BG87, FP80, GLR81, Gar85b, Hal80, Hol81, Hol82, Hol86, Pre86, Roc83, RR83, Ros87, Sam81, SW89, Shu86, Sti80, Sti84, VM88, Yan84, Ros88a]. study [Bur89]. Subclass [SE86]. Subgroup [ML88a]. Subgroups [Wer89]. Subject [SM86]. Subsets [Ham86, Kuk84]. Subspaces [Sch88a, Sch89]. Successive [Sko86]. Sufficiency [God80a, Lar83, McC84a]. Sufficient [Dur80b, Gab84, Jey82]. Sum [Hor82, Joh80, Kha81, LL83, Pet80b, Wor83, Joh82]. Summary [SR83]. Superiority [Gab84]. Superpopulation [Tam84, Tam87]. Superposed [Bri89]. Superposition [Fir89]. Surface [AD89, MH85, Ste85b]. Surrogate [TT88]. Survey [Bed83, CD86, CL87, RRK87, Ski88, SJ81, SS84b, Tam84, Tam86, Tam88a, Yan78, Yan80]. Surveys [Sri80, ST89]. Survival [AA86, BLS^{+}82, DF89, FD83, Fra84, GLR81, HW80, HF82, Hil81, Hou86b, Hou88b, HCT83, JT85, Jen84, KO88, LBD88, Las82, Lus80, Mal88, Meh81, Nag80, NH80, Oak86, Rei81, SK81, Sch81b, SR83, Sol84, SP82, Swe87, Tsa88, TRT85, Tur89, VM88, Sol86]. Swamy [SS85]. Switch [OS88]. Switch-Back [OS88]. Sylvester [Hol86]. Symmetric [Bre82, Bro81, Hor82, DG85]. Symmetry [BGW82, CH87, LM85b, Tay85]. Synthetic [Leu87]. System [Bre84, Win80]. Systematic [BR75, BR81, Kot86, Kot87]. Systems [CS89b, HKM86, TJ82, Vas81].

Table [Fri80, Gar87, Hou84, KL83, Wan87, Ede86]. Tables [AE83, Bed83, Bri80, CP80, Dav85b, EK83, EH85, EGG89, Gil84, Goo81, Hir83, KY82, LSW84, Naz87, TA83, Wah80, WL83, Gar85a, LSW89, MB88, Zid84]. Tag [SF81]. Tagging [Lei88]. Tapering [Kan87]. Tau [BCE83, KG87]. Technique [SSW89]. Techniques [Bes81]. Tensor [McC84b]. Test [AL89, BMW86, Ber89, BCE83, Bro84, Che84a, CS89c, Cox84, DM88a, Ede89, ESP82, EP83, FM82, FR82, GS87, Gup89, HW83, HF82, HP80a, HR85, Hom88, HCT83, Jen83, Jup87, Kai83, KR86, Kee85, KS89, Koc81, KG87, Kou80, LL83, LM89, LSG79, Li88, Lia87b, LM80, LM85b, Man85, MI81, Mil81, Moo80, MJ89, MT81, Nag80, Nag84, NS88, O’B80, Özt86, PD86, Reg80, Res89, SO82, Sas80, SL80, Sch85, ST87, Sch84b, Sha85, SS83b, Spi80, TGG89, Tsa86a, Tsi80, Tsi81, Tyl82, VS80, Whi86, WFBH87, HW84a, Law84, O’B82, Sch87b, Smi80a, WFBH88, LSG83]. Testing [Azz84, Bar83, BT83, Bho84, Chi80, CS88b, CH87, Dav87b, EM84, FD83, FH83, Gar85b, Gea88, God80b, GT84b, Hal89a, Hal90, Har85b, HP85, Hir82, Hir86, Hos84, Kie82, KG87, KY82, Lee85, Lju86, LST88, Mac83, ML88b, MOL86, Nai86, PP88, PG82, Pie85, PT80, RW82, SD84, Sch88b, SW89, Shu82, SM86, SS89b, SL82, TSE88, TMKG87, WH89, Wol88, Yan78, Yan80, Zhe88, Azz87, Bur89, SL84]. Tests [BDK89, BB89a, Bas81b, BBS84, BHS85, Bed83, BK87a, BLS^{+}82, BGW82, Bro82b, BKB89, Che85, Col89, Cuz82, DH87, DP88, Des83, FM82, Fen83, GTP88, Gat86, Gor80, GG80, Har86, Hil81, Hil86, Hin88, Hoc88, HP75, HP76, HP80b, JJS87, Jam87, Jen87, Jen84, Jen86, Jey82, JW79, JL86, Kee82, Kee84, Ken82b, Ken86, KD87, Kim88, Kim85, Koz80, Leu83a, Leu83b, LS85a, LF80, LS85c, Lom87, MR87, McA83, MPW88, MBBW81, MM89, Nai81, New80a, O’N82, Oja81, Oja83, dBP81, Pes84, Pet82, Pre78, RB86, Sch80, Sch81b, Sch84b, SS82, Sen82, Shi81b, SH82, Sim86, SW87, Spi83, Sun86, Tar85, TT88, Tsa86b, TRT85, WJ85, Wei88, Wes88, WL80, WL81, Wor83, Wor86, BC89, Gar85a, JW81, Ken82a, Pre83, Sie80, SW83, WL84]. Theorem [Ber88]. Theory [iA82, BB86, CC88, CFR88, Daw81, Mak85, Phi87, Taw88, iA83]. there [SS84a]. Third [MH85]. Third-Order [MH85]. Thompson [Kot88, PS80]. Three [DW85, Leu83b, Naz87, Pet80a, Sic82]. Three-Component [DW85]. Three-Way [Naz87]. Threshold [PD86, Ton82, Ton83a]. Tied [GLR81, Pet84]. Time [Abr80, Azz84, BBS84, BHS85, Bol88, Bri89, CMH87, CJ88, CRC89, DA85, De 88a, DK87b, Fin86, Fra84, FH83, HO81, HLP88, Hou86a, Hou88a, Hur88, HT89, JT85, KS81, Kee85, KA85, KA87, LM86b, Li88, LH89, Lju86, Lou81, Lus80, Lüt82, Mar80, MB89, Mil81, Mil84, New80a, PD86, PP88, PT83, PT87, PWP81, Pre82, PN85, Reg80, Rei81, RW87, Roy89, SF89, TG80, TP83, Ton82, Ton83b, TT85, Tsa86b, Tsi81, VRW86, Wal88, Wil83, ZLS85, Zeg88, Zhe88, Azz87, DA87, Fin90, Pre84b, TP84, Ton83a]. Time-Censored [Reg80]. Time-Dependent [Lus80]. Time-Independent [ZLS85]. Time-Reversibility [HLP88]. Times [GLR81]. Tool [Hir86]. Total [MS89, SSW89]. Tournaments [Rus80]. Transformation [AL89, CR81, GJ82, Hin85, Hin88, Law87, Mak81]. Transformations [AO81, BLE88, DiC84, KL87, Kon81, Per81, Sol85, TJ82, Tay85, WR80b, AO83]. Transformed [Swe84]. Transition [LST88]. Treatment [Cur88, GWP84, GTP88, HA87, Rob83, SE86, ZH89]. Treatments [CW80, Gup89, Mat87, PR87b, Mat88b]. Tree [WT85]. Trend [BR75, Bri89, FH83, BR81]. Trial [RT88]. Trials [Atk82, CP81, DW82, DW80, DF85, GTP88, Gea88, GP82, Gre85, Gün87, HT81b, LD83, Mat89, Pat85b, Pat86, Pet80a, Pre86, SS89b, Ste80, TGG89, TSE88]. Triangles [PW86]. Trimmed [BK83, LF80, Mar84]. True [Phi82]. Truncated [AB82b, Cha87, LBD88, SG85, Tsa88]. Truncation [SF89, TJW87]. Tukey [Kee85, Law84]. Tumour [Por86]. Tumourigenicity [BKB89]. Two [AO81, Atk81, Bai87, BGW82, Blæ81, BC88b, BJV88, Col89, CF84, DDS89, Don87, DF85, EGG89, Flu87, Gar87, Gil84, HT81a, HT81b, Hir83, Hom89, Hor82, HM84, JL86, Jup87, Kee82, Kha81, Koc81, Kün87b, LM85a, LSW84, Mac83, Mat87, MMM82, Meh81, New80a, Oja81, OV81, PW82, Pal87, Pat85b, Pet84, Sch88a, Shu82, ST89, Ste85b, SC85, Swa83, Tab87, TSE88, Vec85, Wei88, AO83, Kee84, Las82, LSW89, Mat88b, SS85, Sch89, SH82, WR82]. Two-Armed [HT81a, HT81b]. Two-Dimensional [Blæ81, DF85, Kün87b, Vec85, Las82]. Two-Level [Ste85b]. Two-Parameter [Col89]. Two-Period [Pat85b]. Two-Phase [ST89]. Two-Sample [Jup87, PW82, Tab87, Wei88, SH82]. Two-Sided [Kha81]. Two-Stage [BC88b, TSE88]. Two-Way [Bai87, EGG89, Gil84, Hir83, Kee82, Kee84, LSW84, OV81, Shu82, SC85, LSW89, WR82]. Type [AR81, CA83, D’E81, Fal89, Kee85, PD86, Pet80b, RS80, RB86, STV84, AR72, KD87].

Ultrastructural [Mor81b, Gle85]. Unbalanced [Dav87a, Lon87]. Unbiased [Gol89b, KR87, PS82, RL84, SS80, Tam88a, PS84, Sär80]. Unbiasedness [ZH89]. Uncertain [SF89]. Uncertainty [AA86]. Unconditional [Lub81]. Underlying [Ken86, Sun86]. Unequal [BS84, Cha82b, Gab81, Gab84, GJ83, GN87, Jen83, Jen84, Kuk88, SA81, SSS86, Sen89, Jac80]. Unequally [DCS87, Fal89]. Unidentified [DP88]. Unified [Phi87]. Uniformity [HR85, Mar81]. Unimodality [Bur82, Gab82]. Uniqueness [Fin86, HLP88, Fin90]. Unit [FL86, Hal89a, PP88, SD84, Saw83, Hal90]. Univariate [Lju86, PN85]. Universal [Yeh86]. Unknown [BGW82, Cha82a, CL80, SD84]. Unmasked [Atk86]. Unobserved [Gei81]. Unpaired [DCS87, WGJJ89]. Unpredictable [Jen87]. unreplicated [Gle85]. Unsigned [Pre84a]. Updating [BS89]. Upon [SH80]. Upper [AR72, Pat83b, PS82, AR81, PS84]. Use [ABH89, Bha89, CV89, Cur88, Fri80, GJ82, Hal81b, Hir82, Lub81, Mat81, Mor81a, Ren88, RO88, RO89, Ste82, Tam82, TT85, HCT83, LU80]. Used [SK85, TP83, TP84]. Useful [HJ82, FR87]. Uses [AS80]. Using [BG84, BM86, CS89c, FM82, Gol86, Gro80, HO81, HW88, KF83, Kou82, Krz83, LZ86, MR87, Mar84, Pet81, Pet83, Pet86, Sie82, SV80, SJ81, Tam88b, Wah80, WR80b].

Vaeth [WR82]. Valid [Jey82]. Validation [AK87, Bow84, BHT84, De 88a, Gre85, Hal82a, KA89, Bur89]. Validatory [FS84]. Value [AR72, Hos84, LSW84, Özt86, PW80b, Taw88, AR81, LSW89]. Values [BCE83, FM82, JR80, LS85b, SS82]. Variability [CS86b, Far82]. Variable [Jew85, Kem84, Lau81, Lev84, RZ87, TMKG87, Yeh88]. Variable-Selection [Kem84]. Variables [AE83, BJV88, CSL^{+}84, Coo80, CF84, Fin86, FL86, Har85b, KL87, Krz83, Lag81, Nag82, Pal81, Shi81a, Shi84, TJ82, Van80, WF82, Won89, Wor86, Yan78, Yan80, Zam89, Fin90, Shi82]. Variance [BLE88, BB87, Bro84, BES88, BE89, CF84, DCS88, DK80, Fra80, GPT85, GSJS86, Gla86b, Gla88, GN87, Har85b, HM84, JR81, KR86, KF83, KY82, Man85, MT80, MK85b, MH88, Nai86, OV81, Par83, PS82, PN85, RL84, SSW89, SH80, Sol85, Tib88, WGPB89, Wes88, Wil80, Wu82, PS84, WR82]. Variance-Balanced [MK85b]. Variances [Bho84, BK83, Cha82a, DG82, Ekb82, Gla86a, LG84, Nag84, Sch81a, Shu82]. Variate [BLE88, Daw81, Krz89, PGT86]. Variates [Bho84]. Variation [BP88b, DG82, Mor87b]. Vary [Ano81a, BO78]. Vector [AK83, Gei81, Moo80, TGG89]. Vectors [HP85, Jup87, Saw83]. Version [SIK83]. Versus [CW80, Dal84, Fli85, Gab81, Sär80]. via [Fow87, MJ83]. Vibrations [HO81]. Victorian [Sti80]. View [Joh85]. Vocabulary [BK87b]. Volume [Ano80g, Ano81h, Ano82g, Ano83g, Ano84g, Ano85g, Ano86g, Ano87i, Ano87j, Ano88i, Ano88j, Ano89i, Ano89j]. Vulnerability [Ste80].

Wallis [Fli85]. Waring [Hol86]. Watson [Ali84, Bha89, Fre81, MJ89]. Way [Bai87, BP88b, EGG89, Gil84, Hir83, Kee82, Kee84, LSW84, Naz87, OV81, Shu82, SC85, Tab87, Law84, LSW89, MB88, WR82]. Weak [Gol80]. Weaknesses [Leu83b]. Weibull [Col89, Kim85, SG85]. Weight [Gün87]. Weighted [CC88, Dav85b, GPT85, Gre88, Joh80, Kun85, ME82, SSW89, Joh82]. weighting [Sär80]. Weights [CC88]. Welch [Joh82, Joh80]. Whether [Hos84]. Which [BFR89, Bel84, HHP82]. White [KR86, Res89]. Whitney [PS81]. Wilcoxon [JW81, BGW82, JW79, LL83, PS81]. Wilks [Dav80]. Williams [Law84]. Window [Bha80]. Windows [Hur88]. Winner [Wei88]. Wishart [HS80]. Within [Bec81, PS81]. Within-Sample [PS81]. Without [BS84, Gab84, KY82]. Write [TE87].

XL [Sti84]. XLI [Hol86]. XXXIX [Hol82, Hol81]. XXXVIII [Sti80].

Zero [Hil86, Hos84, Pet80b]. Zero-One [Pet80b].

References

Albert:1984:EML

[AA84] A. Albert and J. A. Anderson. On the existence of maximum likelihood estimates in logistic regression models. Biometrika, 71(1): 1–10, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

          3510 (electronic). URL http://www.jstor.org/stable/2336390.

Altman:1986:NUS

[AA86] Douglas G. Altman and Per Kragh Andersen. A note on the uncertainty of a survival probability estimated from Cox’s regression model. Biometrika, 73(3):722–724, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

          http://www.jstor.org/stable/2336538.

Anderson:1982:PML

[AB82a] J. A. Anderson and V. Blair. Penalized maximum likelihood estimation in logistic regression and discrimination. Biometrika, 69 (1):123–136, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335860.

Arghami:1982:MTP

[AB82b] N. R. Arghami and L. Billard. A modification of a truncated partial sequential procedure. Biometrika, 69(3):613–618, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335996.

Azzalini:1989:UNR

[ABH89] A. Azzalini, A. W. Bowman, and W. Härdle. On the use of nonparametric regression for model checking. Biometrika, 76(1):1–11, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336363.

Abraham:1980:IAM

[Abr80] Bovas Abraham. Intervention analysis and multiple time series. Biometrika, 67(1):73–78, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335318.

Aitchison:1984:LCM

[ABS84] J. Aitchison and J. Bacon-Shone. Log contrast models for experiments with mixtures. Biometrika, 71(2):323–330, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336249.

Adke:1980:MLE

[AD80a] S. R. Adke and S. R. Dharmadhikari. The maximum likelihood estimation of coefficient of diffusion in a birth and diffusion process. Biometrika, 67(3):571–576, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335125.

Amaral:1980:OEP

[AD80b] Maria Antonia Amaral and Ian R. Dunsmore. Optimal estimates of predictive distributions. Biometrika, 67(3):685–689, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335141.

Atkinson:1989:CEO

[AD89] A. C. Atkinson and A. N. Donev. The construction of exact D-optimum experimental designs with application to blocking response surface designs. Biometrika, 76(3):515–526, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336117.

Asmussen:1983:CRV

[AE83] Søren Asmussen and David Edwards. Collapsibility and response variables in contingency tables. Biometrika, 70(3):567–578, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336492.

Afsarinejad:1983:BRM

[Afs83]  K. Afsarinejad.  Balanced repeated measurements designs.

Biometrika, 70(1):199–204, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335957.

Anscombe:1983:DKS

[AG83] F. J. Anscombe and William J. Glynn. Distribution of the kurtosis statistic b_{2} for normal samples. Biometrika, 70(1):227–234, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335960.

Azzalini:1987:SOD

[AG87] Adelchi Azzalini and Alessandra Giovagnoli. Some optimal designs for repeated measurements with autoregressive errors. Biometrika, 74(4):725–734, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336466.

Aitchison:1989:MPL

[AH89] J. Aitchison and Chih-Hsiang H. Ho. The multivariate Poisson-log normal distribution. Biometrika, 76(4):643–653, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336624.

Ahn:1988:DRA

[Ahn88] S. K. Ahn. Distribution for residual autocovariances in multivariate autoregressive models with structured parameterization. Biometrika, 75(3):590–593, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336612.

Aitchison:1983:PCA

[Ait83] J. Aitchison. Principal component analysis of compositional data. Biometrika, 70(1):57–65, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335943.

Agrawal:1989:NPP

[AJ89] M. C. Agrawal and Nirmal Jain. A new predictive product estimator. Biometrika, 76(4):822–823, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336647.

Ansley:1982:GDF

[AK82] Craig F. Ansley and Robert Kohn. A geometrical derivation of the fixed interval smoothing algorithm. Biometrika, 69(2):486–487, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335428.

Ansley:1983:ELV

[AK83] C. F. Ansley and R. Kohn. Exact likelihood of vector autoregressive-moving average process with missing or aggregated data. Biometrika, 70(1):275–278, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335967.

Ansley:1985:RCI

[AK85] Craig F. Ansley and Robert Kohn. On the rate of convergence of the innovation representation of a moving average process. Biometrika, 72(2):325–330, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336084.

Ansley:1986:PMS

[AK86] Craig F. Ansley and Robert Kohn. Prediction mean squared error for state space models with estimated parameters. Biometrika, 73(2):467–473, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336224.

Ansley:1987:EGC

[AK87] Craig F. Ansley and Robert Kohn. Efficient generalized crossvalidation for state space models. Biometrika, 74(1):139–148, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336028.

Abraham:1984:NIA

[AL84] Bovas Abraham and Johannes Ledolter. A note on inverse autocorrelations. Biometrika, 71(3):609–614, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336571.

Atkinson:1989:CAE

[AL89] A. C. Atkinson and A. J. Lawrance. A comparison of asymptotically equivalent test statistics for regression transformation. Biometrika, 76(2):223–229, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336654.

Ali:1984:AND

[Ali84] Mukhtar M. Ali. An approximation to the null distribution and power of the Durbin–Watson statistic. Biometrika, 71(2):253–261, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336241.

Anonymous:1980:BMa

[Ano80a] Anonymous. Back matter. Biometrika, 67(1):??, April 1980. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335352.

Anonymous:1980:BMb

[Ano80b] Anonymous. Back matter. Biometrika, 67(2):??, August 1980. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335505.

Anonymous:1980:BMc

[Ano80c] Anonymous. Back matter. Biometrika, 67(3):??, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335155.

Anonymous:1980:FMa

[Ano80d] Anonymous. Front matter. Biometrika, 67(1):??, April 1980. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335309.

Anonymous:1980:FMb

[Ano80e] Anonymous. Front matter. Biometrika, 67(2):??, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335469.

Anonymous:1980:FMc

[Ano80f] Anonymous. Front matter. Biometrika, 67(3):??, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335118.

Anonymous:1980:VI

[Ano80g] Anonymous. Volume information. Biometrika, 67(3):i–ix, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335117.

Anonymous:1981:ACE

[Ano81a] Anonymous. Amendments and corrections: “Estimation of the Size of a Closed Population When Capture Probabilities Vary Among Animals”. Biometrika, 68(1):345, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335844. See [BO78].

Anonymous:1981:BMa

[Ano81b] Anonymous. Back matter. Biometrika, 68(1):??, April 1981. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335846.

Anonymous:1981:BMb

[Ano81c] Anonymous. Back matter. Biometrika, 68(2):??, August 1981. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335611.

Anonymous:1981:BMc

[Ano81d] Anonymous. Back matter. Biometrika, 68(3):??, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335468.

Anonymous:1981:FMa

[Ano81e] Anonymous. Front matter. Biometrika, 68(1):??, April 1981. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335799.

Anonymous:1981:FMb

[Ano81f] Anonymous. Front matter. Biometrika, 68(2):??, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335578.

Anonymous:1981:FMc

[Ano81g] Anonymous. Front matter. Biometrika, 68(3):??, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335439.

Anonymous:1981:VI

[Ano81h] Anonymous. Volume information. Biometrika, 68(3):i–vii, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335438.

Anonymous:1982:BMa

[Ano82a] Anonymous. Back matter. Biometrika, 69(1):??, April 1982. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335890.

Anonymous:1982:BMb

[Ano82b] Anonymous. Back matter. Biometrika, 69(2):??, August 1982. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335437.

Anonymous:1982:BMc

[Ano82c] Anonymous. Back matter. Biometrika, 69(3):??, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336013.

Anonymous:1982:FMa

[Ano82d] Anonymous. Front matter. Biometrika, 69(1):??, April 1982. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335847.

Anonymous:1982:FMb

[Ano82e] Anonymous. Front matter. Biometrika, 69(2):??, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335400.

Anonymous:1982:FMc

[Ano82f] Anonymous. Front matter. Biometrika, 69(3):??, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335983.

Anonymous:1982:VI

[Ano82g] Anonymous. Volume information. Biometrika, 69(3):i–ix, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335982.

Anonymous:1983:BMa

[Ano83a] Anonymous. Back matter. Biometrika, 70(1):??, April 1983. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335981.

Anonymous:1983:BMb

[Ano83b] Anonymous. Back matter. Biometrika, 70(2):??, August 1983. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335577.

Anonymous:1983:BMc

[Ano83c] Anonymous. Back matter. Biometrika, 70(3):x–xi, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336514.

Anonymous:1983:FMa

[Ano83d] Anonymous. Front matter. Biometrika, 70(1):??, April 1983. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335937.

Anonymous:1983:FMb

[Ano83e] Anonymous. Front matter. Biometrika, 70(2):??, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335544.

Anonymous:1983:FMc

[Ano83f] Anonymous. Front matter. Biometrika, 70(3):??, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336489.

Anonymous:1983:VI

[Ano83g] Anonymous. Volume information. Biometrika, 70(3):i–ix, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336488.

Anonymous:1984:BMa

[Ano84a] Anonymous. Back matter. Biometrika, 71(1):??, April 1984. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336425.

Anonymous:1984:BMb

[Ano84b] Anonymous. Back matter. Biometrika, 71(2):??, August 1984. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336264.

Anonymous:1984:BMc

[Ano84c] Anonymous. Back matter. Biometrika, 71(3):x–xii, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336589.

Anonymous:1984:FMa

[Ano84d] Anonymous. Front matter. Biometrika, 71(1):??, April 1984. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336389.

Anonymous:1984:FMb

[Ano84e] Anonymous. Front matter. Biometrika, 71(2):??, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336237.

Anonymous:1984:FMc

[Ano84f] Anonymous. Front matter. Biometrika, 71(3):??, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336551.

Anonymous:1984:VI

[Ano84g] Anonymous. Volume information. Biometrika, 71(3):i–ix, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336550.

Anonymous:1985:BMa

[Ano85a] Anonymous. Back matter. Biometrika, 72(1):??, April 1985. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336360.

Anonymous:1985:BMb

[Ano85b] Anonymous. Back matter. Biometrika, 72(2):??, August 1985. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336106.

Anonymous:1985:BMc

[Ano85c] Anonymous. Back matter. Biometrika, 72(3):x–xi, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336751.

Anonymous:1985:FMa

[Ano85d] Anonymous. Front matter. Biometrika, 72(1):??, April 1985. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336328.

Anonymous:1985:FMb

[Ano85e] Anonymous. Front matter. Biometrika, 72(2):??, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336075.

Anonymous:1985:FMc

[Ano85f] Anonymous. Front matter. Biometrika, 72(3):??, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336720.

Anonymous:1985:VI

[Ano85g] Anonymous. Volume information. Biometrika, 72(3):i–ix, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336719.

Anonymous:1986:BMa

[Ano86a] Anonymous. Back matter. Biometrika, 73(1):??, April 1986. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336300.

Anonymous:1986:BMb

[Ano86b] Anonymous. Back matter. Biometrika, 73(2):??, August 1986. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336236.

Anonymous:1986:BMc

[Ano86c] Anonymous. Back matter. Biometrika, 73(3):xi–xii, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336549.

Anonymous:1986:FMa

[Ano86d] Anonymous. Front matter. Biometrika, 73(1):??, April 1986. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336265.

Anonymous:1986:FMb

[Ano86e] Anonymous. Front matter. Biometrika, 73(2):??, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336202.

Anonymous:1986:FMc

[Ano86f] Anonymous. Front matter. Biometrika, 73(3):??, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336516.

Anonymous:1986:VI

[Ano86g] Anonymous. Volume information. Biometrika, 73(3):i–x, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336515.

Anonymous:1987:BMa

[Ano87a] Anonymous. Back matter. Biometrika, 74(1):??, March 1987. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336041.

Anonymous:1987:BMb

[Ano87b] Anonymous. Back matter. Biometrika, 74(2):??, June 1987. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336164.

Anonymous:1987:BMc

[Ano87c] Anonymous. Back matter. Biometrika, 74(3):??, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336718.

Anonymous:1987:BMd

[Ano87d] Anonymous. Back matter. Biometrika, 74(4):vii–viii, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336487.

Anonymous:1987:FMa

[Ano87e] Anonymous. Front matter. Biometrika, 74(1):??, March 1987. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336015.

Anonymous:1987:FMb

[Ano87f] Anonymous. Front matter. Biometrika, 74(2):??, June 1987. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336135.

Anonymous:1987:FMc

[Ano87g] Anonymous. Front matter. Biometrika, 74(3):??, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336683.

Anonymous:1987:FMd

[Ano87h] Anonymous. Front matter. Biometrika, 74(4):??, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336459.

Anonymous:1987:VIa

[Ano87i] Anonymous. Volume information. Biometrika, 74(1):??, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336014.

Anonymous:1987:VIb

[Ano87j] Anonymous. Volume information. Biometrika, 74(4):i–vi, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336458.

Anonymous:1988:BMa

[Ano88a] Anonymous. Back matter. Biometrika, 75(1):??, March 1988. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336457.

Anonymous:1988:BMb

[Ano88b] Anonymous. Back matter. Biometrika, 75(2):??, June 1988. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336201.

Anonymous:1988:BMc

[Ano88c] Anonymous. Back matter. Biometrika, 75(3):??, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336620.

Anonymous:1988:BMd

[Ano88d] Anonymous. Back matter. Biometrika, 75(4):vi–viii, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336327.

Anonymous:1988:FMa

[Ano88e] Anonymous. Front matter. Biometrika, 75(1):??, March 1988. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336427.

Anonymous:1988:FMb

[Ano88f] Anonymous. Front matter. Biometrika, 75(2):??, June 1988. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336165.

Anonymous:1988:FMc

[Ano88g] Anonymous. Front matter. Biometrika, 75(3):??, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336590.

Anonymous:1988:FMd

[Ano88h] Anonymous. Front matter. Biometrika, 75(4):??, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336302.

Anonymous:1988:VIa

[Ano88i] Anonymous. Volume information. Biometrika, 75(1):??, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336426.

Anonymous:1988:VIb

[Ano88j] Anonymous. Volume information. Biometrika, 75(4):i–v, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336301.

Anonymous:1989:BMa

[Ano89a] Anonymous. Back matter. Biometrika, 76(1):??, March 1989. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336388.

Anonymous:1989:BMb

[Ano89b] Anonymous. Back matter. Biometrika, 76(2):??, June 1989. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336682.

Anonymous:1989:BMc

[Ano89c] Anonymous. Back matter. Biometrika, 76(3):??, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336134.

Anonymous:1989:BMd

[Ano89d] Anonymous. Back matter. Biometrika, 76(4):vii–viii, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336650.

Anonymous:1989:FMa

[Ano89e] Anonymous. Front matter. Biometrika, 76(1):??, March 1989. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336362.

Anonymous:1989:FMb

[Ano89f] Anonymous. Front matter. Biometrika, 76(2):??, June 1989. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336651.

Anonymous:1989:FMc

[Ano89g] Anonymous. Front matter. Biometrika, 76(3):??, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336107.

Anonymous:1989:FMd

[Ano89h] Anonymous. Front matter. Biometrika, 76(4):??, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336622.

Anonymous:1989:VIa

[Ano89i] Anonymous. Volume information. Biometrika, 76(1):??, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336361.

Anonymous:1989:VIb

[Ano89j] Anonymous. Volume information. Biometrika, 76(4):i–vi, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336621.

Aranda-Ordaz:1981:TFT

[AO81] Francisco J. Aranda-Ordaz. On two families of transformations to additivity for binary response data. Biometrika, 68(2):357–363, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335580.

See correction [AO83].

Aranda-Ordaz:1983:ACT

[AO83] Francisco J. Aranda-Ordaz. Amendments and corrections: “On two families of transformations to additivity for binary response data” [Biometrika 68 (1981), no. 2, 357–363; MR0626394 (82h:62046)]. Biometrika, 70(1):303, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335975. See [AO81].

Agrawal:1982:SMC

[AP82] H. L. Agrawal and Jagdish Prasad. Some methods of construction of balanced incomplete block designs with nested rows and columns. Biometrika, 69(2):481–483, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335426.

Antle:1972:UCL

[AR72] Charles E. Antle and Fred Rademaker. An upper confidence limit on the maximum of m future observations from a Type 1 extreme value distribution. Biometrika, 59(2):475–477, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334595. See cor-

rection [AR81].

Antle:1981:ACU

[AR81] Charles E. Antle and Fred Rademaker. Amendments and corrections: “An upper confidence limit on the maximum of m future observations from a Type I extreme value distribution” [Biometrika 59 (1972), 475–477; MR 49 #4143]. Biometrika, 68(3):738, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335466. See

[AR72].

Assaf:1988:DSP

[AR88] David Assaf and Yaakov Ritov. A double sequential procedure for detecting a change in distribution. Biometrika, 75(4):715–722, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336312.

Aitchison:1980:LND

[AS80] J. Aitchison and S. M. Shen. Logistic-normal distributions: Some properties and uses. Biometrika, 67(2):261–272, August 1980. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335470.

Ahtola:1984:PIN

[AT84] J. Ahtola and G. C. Tiao. Parameter inference for a nearly nonstationary first-order autoregressive model. Biometrika, 71 (2):263–272, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336242.

Atkinson:1980:NGI

[Atk80] A. C. Atkinson. A note on the generalized information criterion for choice of a model. Biometrika, 67(2):413–418, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335484.

Atkinson:1981:TGD

[Atk81] A. C. Atkinson. Two graphical displays for outlying and influential observations in regression. Biometrika, 68(1):13–20, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335801.

Atkinson:1982:OBC

[Atk82] A. C. Atkinson. Optimum biased coin designs for sequential clinical trials with prognostic factors. Biometrika, 69(1):61–67, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335853.

Atkinson:1986:MU

[Atk86] A. C. Atkinson. Masking unmasked. Biometrika, 73(3):533–541, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336517.

Anderson:1980:SDG

[AW80] Dorothy Anderson and Ray Watson. On the spread of a disease with gamma distributed latent and infectious periods. Biometrika, 67(1):191–198, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335333.

Azzalini:1981:NED

[Azz81] A. Azzalini. A note on the estimation of a distribution function and quantiles by a kernel method. Biometrika, 68(1):326–328, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335836.

Azzalini:1983:MLE

[Azz83] Adelchi Azzalini. Maximum likelihood estimation of order m for stationary stochastic processes. Biometrika, 70(2):381–387, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335552.

Azzalini:1984:EHT

[Azz84] A. Azzalini. Estimation and hypothesis testing for collections of autoregressive time series. Biometrika, 71(1):85–90, April 1984. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336400. See correction

[Azz87].

Azzalini:1987:ACE

[Azz87] A. Azzalini. Amendments and corrections: “Estimation and hypothesis testing for collections of autoregressive time series” [Biometrika 71 (1984), no. 1, 85–90; MR0738329 (85c:62232)]. Biometrika, 74(3):667, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336713. See [Azz84].

Bailey:1983:RR

[Bai83] R. A. Bailey. Restricted randomization. Biometrika, 70(1):183–198, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335956.

Bailey:1987:OWB

[Bai87] R. A. Bailey. One-way blocks in two-way layouts. Biometrika, 74(1):27–32, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336018.

Banks:1988:HSB

[Ban88] David L. Banks. Histospline smoothing the Bayesian bootstrap. Biometrika, 75(4):673–684, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336308.

Baras:1983:TEP

[Bar83] Mario Baras. Testing equality of probabilities of K mutually exclusive events against ordered alternatives. Biometrika, 70(2):473–478, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335562.

Bartholomew:1984:FFA

[Bar84] D. J. Bartholomew. The foundations of factor analysis. Biometrika, 71(2):221–232, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336238.

Basawa:1981:ECM

[Bas81a] I. V. Basawa. Efficiency of conditional maximum likelihood estimators and confidence limits for mixtures of exponential families. Biometrika, 68(2):515–523, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335600.

Basawa:1981:ECT

[Bas81b] I. V. Basawa. Efficient conditional tests for mixture experiments with applications to the birth and branching processes. Biometrika, 68(1):153–164, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335816.

Bauer:1986:ADS

[Bau86] P. Bauer. Approximation of discrete sequential boundaries. Biometrika, 73(3):759–760, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336547.

Basawa:1980:NEG

[BB80] I. V. Basawa and P. J. Brockwell. A note on estimation for gamma and stable processes. Biometrika, 67(1):234–236, April 1980. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335341.

Bartmann:1981:IC

[BB81] Flavio C. Bartmann and Peter Bloomfield. Inefficiency and correlation. Biometrika, 68(1):67–71, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335806.

Bentler:1986:GLB

[BB86] P. M. Bentler and Maia Berkane. Greatest lower bound to the elliptical theory kurtosis parameter. Biometrika, 73(1):240–241, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336295.

Battaglia:1987:EIE

[BB87] F. Battaglia and R. J. Bhansali. Estimation of the interpolation error variance and an index of linear determinism. Biometrika, 74(4):771–779, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336471.

Bandyopadhyay:1989:NTE

[BB89a] Dipankar Bandyopadhyay and Asit P. Basu. A note on tests for exponentiality by Deshpande. Biometrika, 76(2):403–405, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336676.

Basawa:1989:LSI

[BB89b] I. V. Basawa and L. Billard. Large-sample inference for a regression model with autocorrelated errors. Biometrika, 76(2):283–288, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336661.

Basawa:1984:LST

[BBS84] I. V. Basawa, L. Billard, and R. Srinivasan. Large-sample tests of homogeneity for time series models. Biometrika, 71(1):203–206, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336415.

Breckling:1988:Q

[BC88a] Jens Breckling and Ray Chambers. M-quantiles. Biometrika, 75(4):761–771, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336317.

Breslow:1988:LRT

[BC88b] N. E. Breslow and K. C. Cain. Logistic regression for two-stage case-control data. Biometrika, 75(1):11–20, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336429.

Besag:1989:GMC

[BC89] Julian Besag and Peter Clifford. Generalized Monte Carlo significance tests. Biometrika, 76(4):633–642, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336623.

Best:1983:TCL

[BCE83] D. J. Best, M. A. Cameron, and G. K. Eagleson. A test for comparing large sets of Tau values. Biometrika, 70(2):447–453, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335559.

Boys:1987:DSM

[BD87] R. J. Boys and I. R. Dunsmore. Diagnostic and sampling models in screening. Biometrika, 74(2):365–374, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336151.

Bagai:1989:DFT

[BDK89] Isha Bagai, Jayant V. Deshpande, and Subhash C. Kochar. Distribution free tests for stochastic ordering in the competing risks model. Biometrika, 76(4):775–781, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336637.

Buckley:1986:ALS

[BE86] M. J. Buckley and G. K. Eagleson. Assessing large sets of rank correlations. Biometrika, 73(1):151–157, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336280.

Buckley:1989:GME

[BE89] M. J. Buckley and G. K. Eagleson. A graphical method for estimating the residual variance in nonparametric regression. Biometrika, 76(2):203–210, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336652.

Becker:1981:IDW

[Bec81] Niels Becker. The infectiousness of a disease within households. Biometrika, 68(1):133–141, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335814.

Bedrick:1983:ACS

[Bed83] Edward J. Bedrick. Adjusted chi-squared tests for cross-classified tables of survey data. Biometrika, 70(3):591–595, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336494.

Bellhouse:1984:ORE

[Bel84] D. R. Bellhouse. Optimal randomization for experiments in which autocorrelation is present. Biometrika, 71(1):155–160, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336407.

Bellhouse:1986:RAC

[Bel86] D. R. Bellhouse. Randomization in the analysis of covariance. Biometrika, 73(1):207–211, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336287.

Berman:1981:IMG

[Ber81] Mark Berman. Inhomogeneous and modulated gamma processes. Biometrika, 68(1):143–152, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335815.

Beran:1987:PRL

[Ber87] Rudolf Beran. Prepivoting to reduce level error of confidence sets. Biometrika, 74(3):457–468, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336685.

Berman:1988:TJG

[Ber88] Mark Berman. A theorem of Jacobi and its generalization. Biometrika, 75(4):779–783, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336319.

Beran:1989:TLD

[Ber89] Jan Beran. A test of location for data with slowly decaying serial correlations. Biometrika, 76(2):261–269, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336659.

Besag:1981:RTS

[Bes81] Julian Besag. On resistant techniques and statistical analysis. Biometrika, 68(2):463–469, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335594.

Buckley:1988:ERV

[BES88] M. J. Buckley, G. K. Eagleson, and B. W. Silverman. The estimation of residual variance in nonparametric regression. Biometrika, 75(2):189–199, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336166.

Besag:1989:CFC

[Bes89] Julian Besag. A Candidate’s formula: A curious result in Bayesian prediction. Biometrika, 76(1):183, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336383.

Boente:1988:ABG

[BF88] Graciela Boente and Ricardo Fraiman. On the asymptotic behaviour of general maximum likelihood estimates for the nonregular case under nonstandard conditions. Biometrika, 75(1):45–56, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336433.

Buzas:1989:CUW

[BFR89] Thomas E. Buzas, Claes Fornell, and Byong-Duk Rhee. Conditions under which canonical correlation and redundancy maximization produce identical results. Biometrika, 76(3):618–621, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336129.

Begg:1984:CPL

[BG84] Colin B. Begg and Robert Gray. Calculation of polychotomous logistic regression parameters using individualized regressions. Biometrika, 71(1):11–18, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336391.

Begg:1987:MCC

[BG87] Colin B. Begg and Robert J. Gray. Methodology for case-control studies with prevalent cases. Biometrika, 74(1):191–195, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336034.

Bhattacharya:1982:TMW

[BGW82] P. K. Bhattacharya, J. L. Gastwirth, and A. L. Wright. Two modified Wilcoxon tests for symmetry about an unknown location parameter. Biometrika, 69(2):377–382, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335411.

Becker:1983:IDC

[BH83] Niels Becker and John Llewelyn Hopper. The infectiousness of a disease in a community of households. Biometrika, 70(1):29–39, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335941.

Bhansali:1980:AWE

[Bha80] R. J. Bhansali. Autoregressive and window estimates of the inverse correlation function. Biometrika, 67(3):551–566, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335123.

Bhargava:1989:MOU

[Bha89] Alok Bhargava. Missing observations and the use of the Durbin– Watson statistic. Biometrika, 76(4):828–831, December 1989. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336649.

Bhoj:1984:TEV

[Bho84] Dinesh S. Bhoj. On testing equality of variances of correlated variates with incomplete data. Biometrika, 71(3):639–641, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336579.

Basawa:1985:RTT

[BHS85] I. V. Basawa, R. M. Huggins, and R. G. Staudte. Robust tests for time series with an application to first-order autoregressive processes. Biometrika, 72(3):559–571, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336728.

Bowman:1984:CVN

[BHT84] Adrian W. Bowman, Peter Hall, and D. M. Titterington. Crossvalidation in nonparametric estimation of probabilities and probability densities. Biometrika, 71(2):341–351, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336251.

Binder:1981:ABC

[Bin81] David A. Binder. Approximations to Bayesian clustering rules. Biometrika, 68(1):275–285, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335828.

Brien:1988:ACM

[BJV88] C. J. Brien, A. T. James, and W. N. Venables. An analysis of correlation matrices: Variables cross-classified by two factors. Biometrika, 75(3):469–476, September 1988. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336596.

Boyer:1983:VAT

[BK83] John E. Boyer and Joanna O. Kolson. Variances for adaptive trimmed means. Biometrika, 70(1):97–102, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335946.

Bendre:1987:MET

[BK87a] S. M. Bendre and B. K. Kale. Masking effect on tests for outliers in normal samples. Biometrika, 74(4):891–896, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336486.

Boender:1987:MBA

[BK87b] C. G. E. Boender and A. H. G. Rinnooy Kan. A multinomial Bayesian approach to the estimation of population and vocabulary size. Biometrika, 74(4):849–856, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336479.

Burnett:1989:ERS

[BKB89] R. Burnett, D. Krewski, and S. Bleuer. Efficiency robust score tests for rodent tumourigenicity experiments. Biometrika, 76(2):317–324, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336665.

Blaesild:1981:TDH

[Blæ81] P. Blæsild. The two-dimensional hyperbolic distribution and related distributions, with an application to Johannsen’s bean data. Biometrika, 68(1):251–263, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335826.

Bar-Lev:1988:CCV

[BLE88] Shaul K. Bar-Lev and Peter Enis. On the classical choice of variance stabilizing transformations and an application for a Poisson variate. Biometrika, 75(4):803–804, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336326.

Benedetti:1982:ESS

[BLS^{+}82] Jacqueline K. Benedetti, Ping Yu Liu, Harland N. Sather, Jack Seinfeld, and Michael A. Epton. Effective sample size for tests of censored survival data. Biometrika, 69(2):343–349, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335408.

Babich:1980:RPF

[BM80] G. Babich and D. B. Madan. The relevance of a probabilistic form of invertibility. Biometrika, 67(3):704–705, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335146.

Boos:1986:BMU

[BM86] Dennis D. Boos and John F. Monahan. Bootstrap methods using prior information. Biometrika, 73(1):77–83, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336273.

Bardsley:1986:TMD

[BMW86] W. G. Bardsley, P. B. McGinlay, and A. J. Wright. The F test for model discrimination with exponential functions. Biometrika, 73(2):501–508, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336228.

Barndorff-Nielsen:1980:CR

[BN80] O. Barndorff-Nielsen. Conditionality resolutions. Biometrika, 67 (2):293–310, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335474.

Barndorff-Nielsen:1983:FDM

[BN83] O. Barndorff-Nielsen. On a formula for the distribution of the maximum likelihood estimator. Biometrika, 70(2):343–365, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335549.

Barndorff-Nielsen:1986:IFP

[BN86] O. E. Barndorff-Nielsen. Inference on full or partial parameters based on the standardized signed log likelihood ratio. Biometrika, 73(2):307–322, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336207.

Barndorff-Nielsen:1988:LEA

[BNH88] O. E. Barndorff-Nielsen and Peter Hall. On the level-error after Bartlett adjustment of the likelihood ratio statistic. Biometrika, 75(2):374–378, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336188.

Barndorff-Nielsen:1989:NSE

[BNJL89] O. E. Barndorff-Nielsen, I. R. James, and G. M. Leigh. A note on a semiparametric estimator of mortality. Biometrika, 76(4):803–805, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336642.

Burnham:1978:ESC

[BO78] K. P. Burnham and W. S. Overton. Estimation of the size of a closed population when capture probabilities vary among animals. Biometrika, 65(3):625–633, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335915. See correction [Ano81a].

Bohning:1989:LIM

[Böh89] Dankmar Böhning. Likelihood inference for mixtures: Geometrical and other constructions of monotone step-length algorithms. Biometrika, 76(2):375–383, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336672.

Bolstad:1988:MDL

[Bol88] William M. Bolstad. The multiprocess dynamic linear model with biased perturbations: A real time model for growth hormone level. Biometrika, 75(4):685–692, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336309.

Bowman:1980:NCK

[Bow80] Adrian W. Bowman. A note on consistency of the kernel method for the analysis of categorical data. Biometrika, 67(3):682–684, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335140.

Bowman:1984:AMC

[Bow84] Adrian W. Bowman. An alternative method of cross-validation for the smoothing of density estimates. Biometrika, 71(2):353–360, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336252.

Barlow:1988:RRR

[BP88a] William E. Barlow and Ross L. Prentice. Residuals for relative risk regression. Biometrika, 75(1):65–74, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336435.

Barnwal:1988:AOW

[BP88b] R. K. Barnwal and S. R. Paul. Analysis of one-way layout of count data with negative binomial variation. Biometrika, 75(2):215–222, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336169.

Bellhouse:1975:SSP

[BR75] D. R. Bellhouse and J. N. K. Rao. Systematic sampling in the presence of a trend. Biometrika, 62(3):694–697, December 1975. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335534. See correction

[BR81].

Bellhouse:1981:ACS

[BR81] D. R. Bellhouse and J. N. K. Rao. Amendments and corrections: “Systematic sampling in the presence of a trend” [Biometrika 62 (1975), no. 3, 694–697; MR 52 #15763]. Biometrika, 68(2):576, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335609.

See [BR75].

Breslow:1981:ORE

[Bre81] Norman Breslow. Odds ratio estimators when the data are sparse. Biometrika, 68(1):73–84, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335807.

Breth:1982:NES

[Bre82] M. Breth. Nonparametric estimation for a symmetric distribution. Biometrika, 69(3):625–634, December 1982. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335998.

Breth:1984:SDO

[Bre84] M. Breth. System degeneration and other finite sample nonparametric estimation problems. Biometrika, 71(2):397–401, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336258.

Brier:1980:ACT

[Bri80] Stephen S. Brier. Analysis of contingency tables under cluster sampling. Biometrika, 67(3):591–596, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335128.

Brillinger:1989:CDM

[Bri89] David R. Brillinger. Consistent detection of a monotonic trend superposed on a stationary time series. Biometrika, 76(1):23–30, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336365.

Brown:1981:SQA

[Bro81] B. M. Brown. Symmetric quantile averages and related estimators. Biometrika, 68(1):235–242, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335824.

Brooks:1982:LIT

[Bro82a] R. J. Brooks. On the loss of information through censoring. Biometrika, 69(1):137–144, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335861.

Brown:1982:CMD

[Bro82b] B. M. Brown. Cramér–von Mises distributions and permutation tests. Biometrika, 69(3):619–624, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335997.

Brown:1984:CVL

[Bro84] Mark Brown. On the choice of variance for the log rank test. Biometrika, 71(1):65–74, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336398.

Brooks:1987:OAB

[Bro87a] R. J. Brooks. Optimal allocation for Bayesian inference about an odds ratio. Biometrika, 74(1):196–199, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336035.

Browne:1987:RSI

[Bro87b] M. W. Browne. Robustness of statistical inference in factor analysis and related models. Biometrika, 74(2):375–384, June 1987. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336152. See correction

[Bro88].

Browne:1988:ACR

[Bro88] M. W. Browne. Amendments and corrections: “Robustness of statistical inference in factor analysis and related models” [Biometrika 74 (1987), no. 2, 375–384; MR0903138 (88h:62095)]. Biometrika, 75(2):395, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/

2336193. See [Bro87b].

Bethlehem:1984:SOI

[BS84] Jelke G. Bethlehem and Maarten H. Schuerhoff. Second-order inclusion probabilities in sequential sampling without replacement with unequal probabilities. Biometrika, 71(3):642–644, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336580.

Brown:1989:PDU

[BS89] Philip J. Brown and Rolf Sundberg. Prediction diagnostics and updating in multivariate calibration. Biometrika, 76(2):349–361, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336669.

Bartlett:1981:ESP

[BT81] M. S. Bartlett and L. H. C. Tippett. Egon Sharpe Pearson, 1895– 1980. Biometrika, 68(1):1–11 (1 plate), April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335800.

Bergman:1983:ESD

[BT83] Sten W. Bergman and Bruce W. Turnbull. Efficient sequential designs for destructive life testing with application to animal serial sacrifice experiments. Biometrika, 70(2):305–314, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335545.

Burridge:1982:SUP

[Bur82] J. Burridge. Some unimodality properties of likelihoods derived from grouped data. Biometrika, 69(1):145–151, April 1982. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335862.

Burke:1988:EBD

[Bur88] Murray D. Burke. Estimation of a bivariate distribution function under random censorship. Biometrika, 75(2):379–382, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336189.

Burman:1989:CSO

[Bur89] Prabir Burman. A comparative study of ordinary cross-validation, v-fold cross-validation and the repeated learning-testing methods. Biometrika, 76(3):503–514, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336116.

Butler:1989:APP

[But89] Ronald W. Butler. Approximate predictive pivots and densities. Biometrika, 76(3):489–501, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336115.

Brien:1984:ACM

[BVJM84] C. J. Brien, W. N. Venables, A. T. James, and O. Mayo. An analysis of correlation matrices: Equal correlations. Biometrika, 71(3):545–554, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336563.

Barlow:1981:PAB

[BW81] Richard E. Barlow and Alexander S. Wu. Preposterior analysis of Bayes estimators of mean life. Biometrika, 68(2):403–410, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335585.

Balitskaya:1988:RDE

[BZ88] E. O. Balitskaya and L. A. Zolotuhina. On the representation of a density by an Edgeworth series. Biometrika, 75(1):185–187, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336456.

Chaudhuri:1983:EMS

[CA83] Arijit Chaudhuri and Arun Kumar Adhikary. On the efficiency of Midzuno and Sen’s strategy relative to several ratio-type estimators under a particular model. Biometrika, 70(3):689–693, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336506.

Campbell:1981:NBE

[Cam81] Gregory Campbell. Nonparametric bivariate estimation with randomly censored data. Biometrika, 68(2):417–422, August 1981. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335587.

Castledine:1981:BAM

[Cas81] B. J. Castledine. A Bayesian analysis of multiple-recapture sampling for a closed population. Biometrika, 68(1):197–210, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335820.

Chaloner:1988:BAO

[CB88] Kathryn Chaloner and Rollin Brant. A Bayesian approach to outlier detection and residual analysis. Biometrika, 75(4):651–659, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336306.

Carroll:1988:ATW

[CC88] Raymond J. Carroll and Daren B. H. Cline. An asymptotic theory for weighted least-squares with weights estimated by replication. Biometrika, 75(1):35–43, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336432.

Chambers:1986:EDF

[CD86] R. L. Chambers and R. Dunstan. Estimating distribution functions from survey data. Biometrika, 73(3):597–604, December 1986. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336524.

Critchley:1984:CTN

[CF84] Frank Critchley and Ian Ford. On the covariance of two noncentral F random variables and the variance of the estimated linear discriminant function. Biometrika, 71(3):637–638, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336578.

Critchley:1985:IED

[CF85] Frank Critchley and Ian Ford. Interval estimation in discrimination: The multivariate normal equal covariance case. Biometrika, 72(1):109–116, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336340.

Critchley:1988:IEB

[CFR88] Frank Critchley, Ian Ford, and Omar Rijal. Interval estimation based on the profile likelihood: Strong Lagrangian theory with applications to discrimination. Biometrika, 75(1):21–28, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336430.

Cwik:1982:CSP

[ĆGKP82] J. Ćwik, M. Gołembiewska, T. Kowalczyk, and E. Pleszczyńska. Conceptual and statistical problems of sister dependence. Biometrika, 69(3):513–520, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335986.

Chambers:1981:ESI

[CH81a] R. L. Chambers and C. R. Heathcote. On the estimation of slope and the identification of outliers in linear regression. Biometrika, 68(1):21–33, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335802.

Csorgo:1981:KGM

[CH81b] Sándor Csörgő and Lajos Horváth. On the Koziol–Green model for random censorship. Biometrika, 68(2):391–401, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335584.

Csorgo:1987:TS

[CH87] Sȧndor Csörgő and C. R. Heathcote. Testing for symmetry. Biometrika, 74(1):177–184, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336032.

Chan:1982:LSR

[Cha82a] N. N. Chan. Linear structural relationships with unknown error variances. Biometrika, 69(1):277–279, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335886.

Chao:1982:GPU

[Cha82b] M. T. Chao. A general purpose unequal probability sampling plan. Biometrika, 69(3):653–656, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336002.

Chao:1987:ICR

[Cha87] Min Te Chao. Influence curves for randomly truncated data. Biometrika, 74(2):426–429, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336158.

Chen:1982:NRS

[Che82] Hubert J. Chen. A new range statistic for comparisons of several exponential location parameters. Biometrika, 69(1):257–260, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335881. See

correction [Che84b].

Chen:1984:CGF

[Che84a] Chen-Hsin Chen. A correlation goodness-of-fit test for randomly censored data. Biometrika, 71(2):315–322, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336248.

Chen:1984:ACN

[Che84b] Hubert J. Chen. Amendments and corrections: “A New Range Statistic for Comparisons of Several Exponential Location Parameters”. Biometrika, 71(1):219, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336419. See [Che82].

Cheng:1985:TEF

[Che85] K. F. Cheng. Tests for the equality of failure rates. Biometrika, 72(1):211–215, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336354.

Cheng:1986:MCB

[Che86] Ching-Shui Cheng. A method for constructing balanced incompleteblock designs with nested rows and columns. Biometrika, 73 (3):695–700, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336534.

Chi:1980:THN

[Chi80] P. Y. Chi. Testing for homogeneity: The negative binomial distribution. Biometrika, 67(1):252–254, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335348.

Christensen:1988:MAF

[CJ88] Ronald Christensen and Wesley Johnson. Modelling accelerated failure time with a Dirichlet process. Biometrika, 75(4):693–704, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336310.

Cuzick:1980:FEC

[CL80] Jack Cuzick and Georg Lindgren. Frequency estimation from crossings of an unknown level. Biometrika, 67(1):65–72, April 1980. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335317.

Cryer:1981:SSP

[CL81] Jonathan D. Cryer and Johannes Ledolter. Small-sample properties of the maximum likelihood estimator in the first- order moving average model. Biometrika, 68(3):691–694, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335454.

Cheng:1987:OCS

[CL87] Ching-Shui Cheng and Ker-Chau Li. Optimality criteria in survey sampling. Biometrika, 74(2):337–345, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336148.

Connolly:1988:CLR

[CL88] Margaret A. Connolly and Kung-Yee Liang. Conditional logistic regression models for correlated binary data. Biometrika, 75 (3):501–506, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336600.

Clarke:1980:RMA

[Cla80] M. R. B. Clarke. The reduced major axis of a bivariate sample. Biometrika, 67(2):441–446, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335487.

Chan:1983:EML

[CM83] N. N. Chan and T. K. Mak. Estimation of multivariate linear functional relationships. Biometrika, 70(1):263–267, April 1983. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335965.

Chan:1984:HEL

[CM84] N. N. Chan and T. K. Mak. Heteroscedastic errors in a linear functional relationship. Biometrika, 71(1):212–215, April 1984. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336417.

Camacho:1987:CBT

[CMH87] F. Camacho, A. I. McLeod, and K. W. Hipel. Contemporaneous bivariate time series. Biometrika, 74(1):103–113, March 1987. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336025.

Coles:1989:GFT

[Col89] Stuart G. Coles. On goodness-of-fit tests for the two-parameter Weibull distribution derived from the stabilized probability plot.

Biometrika, 76(3):593–598, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336124.

Constantine:1989:RDS

[Con89] Gregory M. Constantine. Robust designs for serially correlated observations. Biometrika, 76(2):245–251, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336657.

Cooke:1980:OLE

[Coo80] Peter Cooke. Optimal linear estimation of bounds of random variables. Biometrika, 67(1):257–258, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335350.

Cook:1987:PPN

[Coo87] R. Dennis Cook. Parameter plots in nonlinear regression. Biometrika, 74(4):669–677, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336460.

Cordeiro:1987:CLR

[Cor87] Gauss M. Cordeiro. On the corrections to the likelihood ratio statistics. Biometrika, 74(2):265–274, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336140.

Cox:1980:LA

[Cox80] D. R. Cox. Local ancillarity. Biometrika, 67(2):279–286, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335472.

Cox:1982:RDP

[Cox82] D. R. Cox. On the role of data of possibly lowered sensitivity. Biometrika, 69(1):215–219, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335871.

Cox:1983:SRO

[Cox83] D. R. Cox. Some remarks on overdispersion. Biometrika, 70(1):269– 274, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335966.

Cox:1984:EDF

[Cox84] D. R. Cox. Effective degrees of freedom and the likelihood ratio test. Biometrika, 71(3):487–493, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336557.

Cox:1988:NDW

[Cox88] D. R. Cox. A note on design when response has an exponential family distribution. Biometrika, 75(1):161–164, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336449.

Cox:1980:SSC

[CP80] M. A. A. Cox and R. L. Plackett. Small samples in contingency tables. Biometrika, 67(1):1–13, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335310.

Chernoff:1981:SMT

[CP81] Herman Chernoff and A. John Petkau. Sequential medical trials involving paired data. Biometrika, 68(1):119–132, April 1981. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335813.

Cordeiro:1989:ILR

[CP89] Gauss M. Cordeiro and Gilberto A. Paula. Improved likelihood ratio statistics for exponential family nonlinear models. Biometrika, 76(1):93–100, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2336372. See correction [CP91].

Cordeiro:1991:ACI

[CP91] Gauss M. Cordeiro and Gilberto A. Paula. Amendments and corrections: “Improved likelihood ratio statistics for exponential family nonlinear models” [Biometrika 76 (1989), no. 1, 93–100; MR0991426 (90f:62067)]. Biometrika, 78(4):935, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336949. See [CP89].

Carroll:1981:PPT

[CR81] R. J. Carroll and David Ruppert. On prediction and the power transformation family. Biometrika, 68(3):609–615, December 1981.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335443.

Cruddas:1989:TSI

[CRC89] A. M. Cruddas, N. Reid, and D. R. Cox. A time series illustration of approximate conditional likelihood. Biometrika, 76(2):231–237, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336655.

Critchley:1985:IPC

[Cri85] Frank Critchley. Influence in principal components analysis. Biometrika, 72(3):627–636, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336735.

Crowder:1987:LQE

[Cro87] Martin Crowder. On linear and quadratic estimating functions. Biometrika, 74(3):591–597, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336698.

Cox:1979:SER

[CS79] D. R. Cox and E. J. Snell. On sampling and the estimation of rare errors. Biometrika, 66(1):125–132, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335251. See correction [CS82].

Cox:1982:ACS

[CS82] D. R. Cox and E. J. Snell. Amendments and corrections: “On sampling and the estimation of rare errors” [Biometrika 66 (1979), no. 1, 125–132; MR 81f:62015]. Biometrika, 69(2):491, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335430. See [CS79].

Cressie:1985:EBE

[CS85] Noel Cressie and Allan Seheult. Empirical Bayes estimation in sampling inspection. Biometrika, 72(2):451–458, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336097.

C:1986:ESD

[CS86a] E. A. Molina C. and T. M. F. Smith. The effect of sample design on the comparison of associations. Biometrika, 73(1):23–33, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336268.

Cox:1986:AVL

[CS86b] D. R. Cox and P. J. Solomon. Analysis of variability with large numbers of small samples. Biometrika, 73(3):543–554, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336518.

Choi:1988:AML

[CS88a] Youn J. Choi and Norman C. Severo. An approximation for the maximum likelihood estimator of the infection rate in the simple stochastic epidemic. Biometrika, 75(2):392–394, June 1988. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336192.

Cox:1988:TSC

[CS88b] D. R. Cox and P. J. Solomon. On testing for serial correlation in large numbers of small samples. Biometrika, 75(1):145–148, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336445.

Chamberlin:1989:ELP

[CS89a] S. R. Chamberlin and D. A. Sprott. The estimation of a location parameter when the scale parameter is confined to a finite range: The notion of a generalized ancillary statistic. Biometrika, 76(3):609–612, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336127.

Chamberlin:1989:LSP

[CS89b] S. R. Chamberlin and D. A. Sprott. Linear systems of pivotals and associated pivotal likelihoods with applications. Biometrika, 76(4):685–691, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336628.

Cheng:1989:GFT

[CS89c] R. C. H. Cheng and M. A. Stephens. A goodness-of-fit test using Moran’s statistic with estimated parameters. Biometrika, 76 (2):385–392, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336673.

Crowder:1989:BIB

[CS89d] Martin Crowder and Trevor Sweeting. Bayesian inference for a bivariate binomial distribution. Biometrika, 76(3):599–603, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336125.

Carroll:1984:EVB

[CSL^{+}84] Raymond J. Carroll, Clifford H. Spiegelman, K. K. Gordon Lan, Kent T. Bailey, and Robert D. Abbott. On errors-in-variables for binary regression models. Biometrika, 71(1):19–25, April 1984. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336392.

Chen:1980:OCI

[CT80] Hubert J. Chen and Paul J. Tsai. Optimal confidence interval for the largest mean in repeated measurements design. Biometrika, 67(1):119–126, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335324.

Cook:1985:RNR

[CT85] R. D. Cook and Chih-Ling Tsai. Residuals in nonlinear regression. Biometrika, 72(1):23–29, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336331.

Cook:1986:BNR

[CTW86] R. D. Cook, C.-L. Tsai, and B. C. Wei. Bias in nonlinear regression. Biometrika, 73(3):615–623, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336526.

Curnow:1988:UCI

[Cur88] R. N. Curnow. The use of correlated information on treatment effects when selecting the best treatment. Biometrika, 75(2):287– 293, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336176.

Cuzick:1982:RTA

[Cuz82] Jack Cuzick. Rank tests for association with right censored data. Biometrika, 69(2):351–364, August 1982. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335409.

Consonni:1989:SRU

[CV89] Guido Consonni and Piero Veronese. Some remarks on the use of improper priors for the analysis of exponential regression models. Biometrika, 76(1):101–106, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336373.

Costello:1980:PST

[CW80] P. Costello and D. A. Wolfe. Partially sequential treatments versus control multiple comparisons. Biometrika, 67(2):403–412, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335483.

Cheng:1981:NBI

[CW81a] Ch’ing Shui Chêng and Chien-Fu Wu. Nearly balanced incomplete block designs. Biometrika, 68(2):493–500, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335598.

Cooper:1981:EPM

[CW81b] D. M. Cooper and E. F. Wood. Estimation of the parameters of the Markovian representation of the autoregressive-moving average model. Biometrika, 68(1):320–322, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335834.

Cook:1983:DHR

[CW83] R. Dennis Cook and Sanford Weisberg. Diagnostics for heteroscedasticity in regression. Biometrika, 70(1):1–10, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335938.

DeGooijer:1985:MSS

[DA85] J. G. De Gooijer and O. D. Anderson. Moments of the sampled space-time autocovariance and autocorrelation function. Biometrika, 72(3):689–693, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336746. See corrections [DA87].

DeGooijer:1987:ACM

[DA87] J. G. De Gooijer and O. D. Anderson. Amendments and corrections: “Moments of the sampled space-time autocovariance and autocorrelation function” [Biometrika 72 (1985), no. 3, 689–693; MR0817585 (87b:62118)]. Biometrika, 74(3):667, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336714. See [DA85].

Dale:1984:LVG

[Dal84] Jocelyn R. Dale. Local versus global association for bivariate ordered responses. Biometrika, 71(3):507–514, December 1984. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336559.

Daniels:1980:ESA

[Dan80] H. E. Daniels. Exact saddlepoint approximations. Biometrika, 67(1):59–63, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335316.

Daniels:1983:SAE

[Dan83] H. E. Daniels. Saddlepoint approximations for estimating equations. Biometrika, 70(1):89–96, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335945.

Darby:1980:BAP

[Dar80] S. C. Darby. A Bayesian approach to parallel line bioassay. Biometrika, 67(3):607–612, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335130.

Daudin:1980:PAM

[Dau80] J. J. Daudin. Partial association measures and an application to qualitative regression. Biometrika, 67(3):581–590, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335127.

Davis:1980:EMM

[Dav80] A. W. Davis. On the effects of moderate multivariate nonnormality on Wilks’s likelihood ratio criterion. Biometrika, 67(2):419–427, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335485.

Davis:1985:MEL

[Dav85a] Linda June Davis. Modification of the empirical logit to reduce bias in simple linear logistic regression. Biometrika, 72(1):199–202, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336350.

Davis:1985:WAO

[Dav85b] Linda June Davis. Weighted averages of the observed odds ratios when the number of tables is large. Biometrika, 72(1):203–205, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336351.

Davison:1986:APL

[Dav86] A. C. Davison. Approximate predictive likelihood. Biometrika, 73(2):323–332, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2336208. See correction [Dav90].

David:1987:RUP

[Dav87a] H. A. David. Ranking from unbalanced paired-comparison data. Biometrika, 74(2):432–436, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336160.

Davies:1987:HTW

[Dav87b] Robert B. Davies. Hypothesis testing when a nuisance parameter is present only under the alternative. Biometrika, 74(1):33–43, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336019.

Davis:1987:MLD

[Dav87c] A. W. Davis. Moments of linear discriminant functions and an asymptotic confidence interval for the log odds ratio. Biometrika, 74(4):829–840, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336477.

Davison:1990:ACA

[Dav90] A. C. Davison. Amendments and corrections: “Approximate predictive likelihood” [Biometrika 73 (1986), no. 2, 323–332; MR0855892 (87i:62014)]. Biometrika, 77(3):667, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337009. See [Dav86].

Dawid:1981:SMV

[Daw81] A. P. Dawid. Some matrix-variate distribution theory: Notational considerations and a Bayesian application. Biometrika, 68(1):265– 274, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335827.

Pereira:1981:ATE

[dBP81] B. de B. Pereira. Addendum: “Tests and efficiencies of separate regression models” [Biometrika 65 (1978), no. 2, 319–337; MR 80h:62017]. Biometrika, 68(1):345, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335845.

Dolby:1987:FBF

[DCS87] G. R. Dolby, R. M. Cormack, and D. F. Sinclair. On fitting bivariate functional relationships to unpaired and unequally replicated data. Biometrika, 74(2):393–399, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336154.

Davidian:1988:VFM

[DCS88] M. Davidian, R. J. Carroll, and W. Smith. Variance functions and the minimum detectable concentration in assays. Biometrika, 75(3):549–556, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336606.

Dabrowska:1987:ECI

[DD87] Dorota M. Da̧browska and Kjell A. Doksum. Estimates and confidence intervals for median and mean life in the proportional hazard model. Biometrika, 74(4):799–807, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336474.

Dabrowska:1989:GCC

[DDS89] Dorota M. Da̧browska, Kjell A. Doksum, and Jae-Kee Song. Graphical comparison of cumulative hazards for two populations. Biometrika, 76(4):763–773, December 1989. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336636.

DeJong:1988:CVF

[De 88a] Piet De Jong. A cross-validation filter for time series models. Biometrika, 75(3):594–600, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336613.

DeJong:1988:LSS

[De 88b] Piet De Jong. The likelihood for a state space model. Biometrika, 75(1):165–169, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336450.

DEste:1981:MTB

[D’E81] G. M. D’Este. A Morgenstern-type bivariate gamma distribution. Biometrika, 68(1):339–340, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335841.

Deaton:1980:EBA

[Dea80] Leonard W. Deaton. An empirical Bayes approach to polynomial regression under order restrictions. Biometrika, 67(1):111–117, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335323.

Deshpande:1983:CTE

[Des83] Jayant V. Deshpande. A class of tests for exponentiality against increasing failure rate average alternatives. Biometrika, 70(2):514– 518, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2335571.

Draper:1985:RPE

[DF85] N. R. Draper and D. Faraggi. Role of the Papadakis estimator in one- and two-dimensional field trials. Biometrika, 72(1):223–226, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336357.

Dykstra:1989:NML

[DF89] Richard L. Dykstra and Carol J. Feltz. Nonparametric maximum likelihood estimation of survival functions with a general stochastic ordering and its dual. Biometrika, 76(2):331–341, June 1989. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336667.

David:1982:MLV

[DG82] H. A. David and R. A. Groeneveld. Measures of local variation in a distribution: Expected length of spacings and variances of order statistics. Biometrika, 69(1):227–232, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335873.

David:1985:EOE

[DG85] H. A. David and J. K. Ghosh. The effect of an outlier on Lestimators of location in symmetric distributions. Biometrika, 72 (1):216–218, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336355.

Davison:1989:DRN

[DG89] A. C. Davison and A. Gigli. Deviance residuals and normal scores plots. Biometrika, 76(2):211–221, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336653.

Diggle:1987:NEP

[DGS87] Peter J. Diggle, David J. Gates, and Alyson Stibbard. A nonparametric estimator for pairwise-interaction point processes. Biometrika, 74(4):763–770, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336470.

Dupac:1985:SAD

[DH85] Václav Dupač and Ulrich Herkenrath. Stochastic approximation with delayed observations. Biometrika, 72(3):683–685, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336744.

Davies:1987:THE

[DH87] R. B. Davies and D. S. Harte. Tests for Hurst effect. Biometrika, 74(1):95–101, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336024.

Davison:1988:SAR

[DH88] Anthony C. Davison and David V. Hinkley. Saddlepoint approximations in resampling methods. Biometrika, 75(3):417–431, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336592.

DiCiccio:1989:CPE

[DHR89] Thomas J. DiCiccio, Peter Hall, and Joseph P. Romano. Comparison of parametric and empirical likelihood functions. Biometrika, 76(3):465–476, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336113.

Davison:1986:EBS

[DHS86] A. C. Davison, D. V. Hinkley, and E. Schechtman. Efficient bootstrap simulation. Biometrika, 73(3):555–566, December 1986. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336519.

DiCiccio:1984:PTI

[DiC84] T. J. DiCiccio. On parameter transformations and interval estimation. Biometrika, 71(3):477–485, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336556.

DiCiccio:1988:LIL

[DiC88] T. J. DiCiccio. Likelihood inference for linear regression models. Biometrika, 75(1):29–34, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336431.

Ducharme:1985:BCC

[DJRT85] Gilles R. Ducharme, Myoungshic Jhun, Joseph P. Romano, and Kinh N. Truong. Bootstrap confidence cones for directional data. Biometrika, 72(3):637–645, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336736.

Donner:1980:LSV

[DK80] Allan Donner and John J. Koval. The large sample variance of an intraclass correlation. Biometrika, 67(3):719–722, December 1980.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335151.

Dahlhaus:1987:EEE

[DK87a] R. Dahlhaus and H. Künsch. Edge effects and efficient parameter estimation for stationary random fields. Biometrika, 74(4):877–882, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336483.

Dewanji:1987:EST

[DK87b] Anup Dewanji and John D. Kalbfleisch. Estimation of sojourn time distributions for cyclic semi-Markov processes in equilibrium. Biometrika, 74(2):281–288, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336142.

Dinse:1986:NSM

[DL86] Gregg E. Dinse and Martin G. Larson. A note on semi-Markov models for partially censored data. Biometrika, 73(2):379–386, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336214.

Darroch:1985:CPC

[DM85] John N. Darroch and James E. Mosimann. Canonical and principal components of shape. Biometrika, 72(2):241–252, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336077.

Ducharme:1987:SMD

[DM87] Gilles R. Ducharme and Philip Milasevic. Spatial median and directional data. Biometrika, 74(1):212–215, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336038.

Dalal:1988:CTC

[DM88a] Siddhartha R. Dalal and Govind S. Mudholkar. A conservative test and confidence region for comparing heteroscedastic regressions. Biometrika, 75(1):149–152, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336446.

DeJong:1988:CSE

[DM88b] Piet De Jong and Murray J. Mackinnon. Covariances for smoothed estimates in state space models. Biometrika, 75(3):601–602, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336614.

Donner:1987:ORI

[Don87] Allan Donner. Odds ratio inference with dependent data: A relationship between two procedures. Biometrika, 74(1):220, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336040.

Davies:1988:NTP

[DP88] P. Davies and A. J. Phillips. Nonparametric tests of population differences and estimation of the probability of misidentification with unidentified paired data. Biometrika, 75(4):753–760, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336316.

David:1983:OSO

[DR83] H. A. David and M. P. Rogers. Order statistics in overlapping samples, moving order statistics and U-statistics. Biometrika, 70 (1):245–249, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335962.

DiCiccio:1989:ABS

[DR89] Thomas J. DiCiccio and Joseph P. Romano. On adjustments based on the signed root of the empirical likelihood ratio statistic. Biometrika, 76(3):447–456, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336111.

Driessen:1988:PCS

[Dri88] S. G. A. J. Driessen. The probability of correct selection. Biometrika, 75(3):618–619, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336619.

Dupac:1980:PEP

[Dup80] V. Dupac. Parameter estimation in the Poisson field of discs. Biometrika, 67(1):187–190, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335332.

Durbin:1980:ADP

[Dur80a] J. Durbin. The approximate distribution of partial serial correlation coefficients calculated from residuals from regression on Fourier series. Biometrika, 67(2):335–349, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335476.

Durbin:1980:ADS

[Dur80b] J. Durbin. Approximations for densities of sufficient estimators. Biometrika, 67(2):311–333, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335475.

Demets:1980:GSM

[DW80] David L. Demets and James H. Ware. Group sequential methods for clinical trials with a one-sided hypothesis. Biometrika, 67(3):651–660, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335135.

DeMets:1982:AGS

[DW82] David L. DeMets and James H. Ware. Asymmetric group sequential boundaries for monitoring clinical trials. Biometrika, 69 (3):661–663, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336004.

Darroch:1985:AIT

[DW85] J. N. Darroch and J. Waller. Additivity and interaction in threecomponent experiments with mixtures. Biometrika, 72(1):153–163, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336345.

Eaves:1983:BNR

[Eav83] David M. Eaves. On Bayesian nonlinear regression with an enzyme example. Biometrika, 70(2):373–379, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335551.

Edelman:1986:BNT

[Ede86] David Edelman. Bounds for a nonparametric t table. Biometrika, 73(1):242–243, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336296.

Edelman:1989:CLM

[Ede89] David Edelman. A candidata for locally most powerful sequential t test. Biometrika, 76(1):197–201, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336387.

Efron:1981:NES

[Efr81] Bradley Efron. Nonparametric estimates of standard error: The jackknife, the bootstrap and other methods. Biometrika, 68(3): 589–599, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335441.

Efron:1985:BCI

[Efr85] Bradley Efron. Bootstrap confidence intervals for a class of parametric problems. Biometrika, 72(1):45–58, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336334.

Evans:1989:LCA

[EGG89] Michael J. Evans, Zvi Gilula, and Irwin Guttman. Latent class analysis of two-way contingency tables by Bayesian methods. Biometrika, 76(3):557–563, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336120.

Edwards:1985:FPM

[EH85] David Edwards and Tomáš Havránek. A fast procedure for model search in multidimensional contingency tables. Biometrika, 72 (2):339–351, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336086.

Edwards:1983:ACT

[EK83] David Edwards and Svend Kreiner. The analysis of contingency tables by graphical models. Biometrika, 70(3):553–565, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336491.

Ekbohm:1982:CVP

[Ekb82] Gunnar Ekbohm. On comparing variances in the paired case with incomplete data. Biometrika, 69(3):670–673, December 1982. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336007.

Ejigou:1981:RRE

[EM81] Ayenew Ejigou and Richard McHugh. Relative risk estimation under multiple matching. Biometrika, 68(1):85–91, April 1981. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335808.

Ejigou:1984:THR

[EM84] Ayenew Ejigou and Richard McHugh. Testing the homogeneity of the relative risk under multiple matching. Biometrika, 71 (2):408–411, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336261.

Ensor:1988:EOE

[EN88] Katherine B. Ensor and H. Joseph Newton. The effect of order estimation on estimating the peak frequency of an autoregressive spectral density. Biometrika, 75(3):587–589, September 1988. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336611.

Epps:1983:TNB

[EP83] T. W. Epps and Lawrence B. Pulley. A test for normality based on the empirical characteristic function. Biometrika, 70(3):723–726, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336512.

Epps:1982:TSF

[ESP82] T. W. Epps, K. J. Singleton, and L. B. Pulley. A test of separate families of distributions based on the empirical moment generating function. Biometrika, 69(2):391–399, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335413.

El-Sayyad:1989:PDE

[ESSAH89] G. M. El-Sayyad, M. Samiuddin, and A. A. Al-Harbey. On parametric density estimation. Biometrika, 76(2):343–348, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336668.

Falk:1989:HBT

[Fal89] R. W. Falk. Hommel’s Bonferroni-type inequality for unequally spaced levels. Biometrika, 76(1):189–191, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336385.

Farewell:1982:NRA

[Far82] V. T. Farewell. A note on regression analysis of ordinal data with variability of classification. Biometrika, 69(3):533–538, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335988.

Farewell:1983:CPT

[FD83] V. T. Farewell and S. Dahlberg. On the comparison of procedures for testing the equality of survival curves. Biometrika, 70 (3):707–709, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336508.

Fenstad:1983:CBT

[Fen83] G. U. Fenstad. A comparison between the U and V tests in the Behrens–Fisher problem. Biometrika, 70(1):300–302, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335973.

Ferreira:1981:EFM

[Fer81] Pedro E. Ferreira. Extending Fisher’s measure of information. Biometrika, 68(3):695–698, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335455.

Ferreira:1982:EEP

[Fer82] P. E. Ferreira. Estimating equations in the presence of prior knowledge. Biometrika, 69(3):667–669, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336006.

Feuerverger:1989:ESA

[Feu89] Andrey Feuerverger. On the empirical saddlepoint approximation. Biometrika, 76(3):457–464, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336112.

Field:1982:SSA

[FH82] Christopher A. Field and Frank R. Hampel. Small-sample asymptotic distributions of M-estimators of location. Biometrika, 69(1): 29–46, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2335850.

Franzini:1983:TDT

[FH83] L. Franzini and A. C. Harvey. Testing for deterministic trend and seasonal components in time series models. Biometrika, 70 (3):673–682, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336504.

Finch:1980:LSD

[Fin80] Peter D. Finch. Least squares description of a record as signal plus noise. Biometrika, 67(3):539–550, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335122.

Findley:1986:UMA

[Fin86] David F. Findley. The uniqueness of moving average representations with independent and identically distributed random variables for non-Gaussian stationary time series. Biometrika, 73(2):520–521, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336232.

See correction [Fin90].

Findley:1990:ACU

[Fin90] David F. Findley. Amendments and corrections: “The uniqueness of moving average representations with independent and identically distributed random variables for non-Gaussian stationary time series” [Biometrika 73 (1986), no. 2, 520–521; MR0855914 (88e:62210)]. Biometrika, 77(1):235, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336071. See [Fin86].

Firth:1987:EQL

[Fir87] David Firth. On the efficiency of quasi-likelihood estimation. Biometrika, 74(2):233–245, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336137.

Firth:1989:MHS

[Fir89] David Firth. Marginal homogeneity and the superposition of Latin squares. Biometrika, 76(1):179–182, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336382.

Frydenberg:1989:ILR

[FJ89] Morten Frydenberg and Jens Ledet Jensen. Is the “improved likelihood ratio statistic” really improved in the discrete case? Biometrika, 76(4):655–661, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336625.

Fedorov:1986:DOD

[FK86] V. Fedorov and V. Khabarov. Duality of optimal designs for model discrimination and parameter estimation. Biometrika, 73(1):183– 190, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336284.

Fisher:1981:NMA

[FL81] N. I. Fisher and A. J. Lee. Nonparametric measures of angularlinear association. Biometrika, 68(3):629–636, December 1981. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335445.

Fisher:1982:NMA

[FL82] N. I. Fisher and A. J. Lee. Nonparametric measures of angularangular association. Biometrika, 69(2):315–321, August 1982. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335405.

Fisher:1983:CCC

[FL83a] N. I. Fisher and A. J. Lee. A correlation coefficient for circular data. Biometrika, 70(2):327–332, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335547.

Fisher:1983:ECM

[FL83b] Nicholas I. Fisher and Toby Lewis. Estimating the common mean direction of several circular or spherical distributions with differing dispersions. Biometrika, 70(2):333–341, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335548. See correction [FL84].

Fisher:1984:ACE

[FL84] Nicholas I. Fisher and Toby Lewis. Amendments and corrections: “Estimating the common mean direction of several circular or spherical distributions with differing dispersions” [Biometrika 70 (1983), no. 2, 333–341; MR0712022 (85a:62074)]. Biometrika, 71(3):655, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336584. See [FL83b].

Fisher:1986:CCR

[FL86] N. I. Fisher and A. J. Lee. Correlation coefficients for random variables on a unit sphere or hypersphere. Biometrika, 73(1):159– 164, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336281.

Frydenberg:1989:DML

[FL89] Morten Frydenberg and Steffen L. Lauritzen. Decomposition of maximum likelihood in mixed graphical interaction models. Biometrika, 76(3):539–555, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336119.

Fletcher:1987:NCC

[Fle87] David J. Fletcher. A new class of change-over designs for factorial experiments. Biometrika, 74(3):649–654, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336708.

Fligner:1985:PVJ

[Fli85] Michael A. Fligner. Pairwise versus joint ranking: Another look at the Kruskal–Wallis statistic. Biometrika, 72(3):705–709, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336750.

Flury:1987:TGC

[Flu87] Bernhard K. Flury. Two generalizations of the common principal component model. Biometrika, 74(1):59–69, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336021.

Ferreira:1981:CA

[FM81] P. E. Ferreira and Ch. E. Minder. On a concept of ancillarity. Biometrika, 68(1):344, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335843.

Fairbanks:1982:VTU

[FM82] Kenneth Fairbanks and Richard Madsen. P values for tests using a repeated significance test design. Biometrika, 69(1):69–74, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335854.

Fowlkes:1987:SDB

[Fow87] Edward B. Fowlkes. Some diagnostics for binary logistic regression via smoothing. Biometrika, 74(3):503–515, September 1987. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336689.

Farewell:1980:APL

[FP80] V. T. Farewell and R. L. Prentice. The approximation of partial likelihood with emphasis on case-control studies. Biometrika, 67(2):273–278, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335471.

Fligner:1982:MMM

[FR82] Michael A. Fligner and Steven W. Rust. A modification of Mood’s median test for the generalized Behrens–Fisher problem. Biometrika, 69(1):221–226, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335872.

Federer:1987:RMM

[FR87] W. T. Federer and D. Raghavarao. Response models and minimal designs for mixtures of n of m items useful for intercropping and other investigations. Biometrika, 74(3):571–577, September 1987.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336696.

Fraser:1988:CIR

[FR88] D. A. S. Fraser and N. Reid. On conditional inference for a real parameter: A differential approach on the sample space. Biometrika, 75(2):251–264, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336173.

Fraser:1989:APL

[FR89] D. A. S. Fraser and N. Reid. Adjustments to profile likelihood. Biometrika, 76(3):477–488, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336114.

Frangos:1980:VES

[Fra80] Christos C. Frangos. Variance estimation for the second-order jackknife. Biometrika, 67(3):715–718, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335150.

Franco:1984:LLM

[Fra84] M. A. Paiva Franco. A log logistic model for survival time with covariates. Biometrika, 71(3):621–623, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336573.

Franke:1985:LDR

[Fra85] J. Franke. A Levinson–Durbin recursion for autoregressive-moving average processes. Biometrika, 72(3):573–581, December 1985. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336729. See correction

[Fra88].

Franke:1988:ACL

[Fra88] J. Franke. Amendments and corrections: “A Levinson–Durbin recursion for autoregressive-moving average processes” [Biometrika 72 (1985), no. 3, 573–581; MR0817571 (87k:62150)]. Biometrika, 75(2):395, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/

2336194. See [Fra85].

Freedman:1981:WSD

[Fre81] L. S. Freedman. Watson’s U_{N}^{2} statistic for a discrete distribution. Biometrika, 68(3):708–711, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335458.

Frisen:1980:CUC

[Fri80] M. Frisén. Consequences of the use of conditional inference in the analysis of a correlated contingency table. Biometrika, 67(1):23–30, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335312.

Ford:1980:SCD

[FS80] I. Ford and S. D. Silvey. A sequentially constructed design for estimating a nonlinear parametric function. Biometrika, 67(2):381–388, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335481.

Frangos:1984:JCV

[FS84] C. C. Frangos and M. Stone. On jackknife, cross-validatory and classical methods of estimating a proportion with batches of different sizes. Biometrika, 71(2):361–366, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336253.

Fisher:1985:CPA

[FS85] N. I. Fisher and P. Switzer. Chi-plots for assessing dependence. Biometrika, 72(2):253–265, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336078.

Ford:1985:ISD

[FTW85] I. Ford, D. M. Titterington, and C.-F. J. Wu. Inference and sequential design. Biometrika, 72(3):545–551, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336726.

Fujikoshi:1980:AED

[Fuj80a] Yasunori Fujikoshi. Asymptotic expansions for the distributions of the sample roots under nonnormality. Biometrika, 67(1):45–51, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335314. See

correction [Fuj86].

Fujino:1980:ABC

[Fuj80b] Yoritake Fujino. Approximate binomial confidence limits. Biometrika, 67(3):677–681, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335139.

Fujikoshi:1986:ACA

[Fuj86] Y. Fujikoshi. Amendments and corrections: “Asymptotic expansions for the distributions of the sample roots under nonnormality” [Biometrika 67 (1980), no. 1, 45–51; MR0570503 (81h:62036)]. Biometrika, 73(1):245, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336297. See [Fuj80a].

Gabler:1981:CSS

[Gab81] S. Gabler. A comparison of Sampford’s sampling procedure versus unequal probability sampling with replacement. Biometrika, 68(3):725–727, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335462.

Gabrielsen:1982:ULC

[Gab82] Gorm Gabrielsen. On the unimodality of the likelihood for the Cauchy distribution: Some comments. Biometrika, 69(3):677–678, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336009.

Gabler:1984:UPS

[Gab84] S. Gabler. On unequal probability sampling: Sufficient conditions for the superiority of sampling without replacement. Biometrika, 71(1):171–175, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336409.

Gart:1985:ATI

[Gar85a] John J. Gart. Approximate tests and interval estimation of the common relative risk in the combination of 2×2 tables. Biometrika, 72(3):673–677, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336742.

Gart:1985:TIM

[Gar85b] John J. Gart. Testing for interaction in multiply-matched casecontrol studies. Biometrika, 72(2):468–470, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336100.

Gilmour:1985:ABD

[GAR85c] A. R. Gilmour, R. D. Anderson, and A. L. Rae. The analysis of binomial data by a generalized linear mixed model. Biometrika, 72(3):593–599, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336731.

Gart:1987:ETC

[Gar87] John J. Gart. The equivalence of two corrections to the approximate mean of an entry in a contingency table. Biometrika, 74(3):661–663, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336711.

Gates:1986:MTA

[Gat86] J. Gates. Measures and tests of alignment. Biometrika, 73(3): 731–734, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336540.

Godolphin:1982:MLE

[GD82] E. J. Godolphin and J. G. De Gooijer. On the maximum likelihood estimation of the parameters of a Gaussian moving average process. Biometrika, 69(2):443–451, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335420.

Geary:1988:STC

[Gea88] D. N. Geary. Sequential testing in clinical trials with repeated measurements. Biometrika, 75(2):311–318, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336179.

Geisser:1981:SRP

[Gei81] Seymour Geisser. Sample reuse procedures for prediction of the unobserved portion of a partially observed vector. Biometrika, 68(1):243–250, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335825.

Genest:1987:FFB

[Gen87] Christian Genest. Frank’s family of bivariate distributions. Biometrika, 74(3):549–555, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336693.

Gupta:1980:NTR

[GG80] G. D. Gupta and Z. Govindarajulu. Nonparametric tests of randomness against autocorrelated normal alternatives. Biometrika, 67(2):375–379, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335480.

Goudie:1981:ISE

[GG81] I. B. J. Goudie and Charles M. Goldie. Initial size estimation for the linear pure death process. Biometrika, 68(2):543–550, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335603.

Gower:1988:NB

[GH88] J. C. Gower and S. A. Harding. Nonlinear biplots. Biometrika, 75(3):445–455, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336594.

Gilula:1984:SSB

[Gil84] Zvi Gilula. On some similarities between canonical correlation models and latent class models for two-way contingency tables. Biometrika, 71(3):523–529, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336561.

Guerrero:1982:UBC

[GJ82] Victor M. Guerrero and Richard A. Johnson. Use of the Box– Cox transformation with binary response models. Biometrika, 69(2):309–314, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335404.

Gupta:1983:EBB

[GJ83] S. C. Gupta and B. Jones. Equireplicate balanced block designs with unequal block sizes. Biometrika, 70(2):433–440, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335557.

Gould:1988:CER

[GL88] Ann Gould and J. F. Lawless. Consistency and efficiency of regression coefficient estimates in location- scale models. Biometrika, 75(3):535–540, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336604.

Glaser:1980:CBS

[Gla80] Ronald E. Glaser. A characterization of Bartlett’s statistic involving incomplete beta functions. Biometrika, 67(1):53–58, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335315.

Glasbey:1986:CEV

[Gla86a] C. A. Glasbey. Conservative estimates of the variances of regression parameter estimators for classes of error model. Biometrika, 73(3):746–750, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336544.

Glasbey:1986:DSS

[Gla86b] C. A. Glasbey. A decomposition of some serially-structured variance matrices. Biometrika, 73(2):447–453, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336221.

Glasbey:1988:SER

[Gla88] C. A. Glasbey. Standard errors resilient to error variance misspecification. Biometrika, 75(2):201–206, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336167.

Gleser:1985:NGR

[Gle85] Leon Jay Gleser. A note on G. R. Dolby’s unreplicated ultrastructural model. Biometrika, 72(1):117–124, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336341.

Gail:1981:LCM

[GLR81] Mitchell H. Gail, Jay H. Lubin, and Lawrence V. Rubinstein. Likelihood calculations for matched case-control studies and survival studies with tied death times. Biometrika, 68(3):703–707, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335457.

Gupta:1987:MOA

[GN87] V. K. Gupta and A. K. Nigam. Mixed orthogonal arrays for variance estimation with unequal numbers of primary selections per stratum. Biometrika, 74(4):735–742, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336467.

Gupta:1982:FSS

[GNK82] V. K. Gupta, A. K. Nigam, and Pranesh Kumar. On a family of sampling schemes with inclusion probability proportional to size. Biometrika, 69(1):191–196, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335868.

Godambe:1980:SAP

[God80a] V. P. Godambe. On sufficiency and ancillarity in the presence of a nuisance parameter. Biometrika, 67(1):155–162, April 1980. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335328.

Godolphin:1980:MTO

[God80b] E. J. Godolphin. A method for testing the order of an autoregressive-moving average process. Biometrika, 67(3):699–703, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335145.

Godambe:1984:AFI

[God84a] V. P. Godambe. On ancillarity and Fisher information in the presence of a nuisance parameter. Biometrika, 71(3):626–629, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336575.

Godolphin:1984:DRL

[God84b] E. J. Godolphin. A direct representation for the large-sample maximum likelihood estimator of a Gaussian autoregressive-moving average process. Biometrika, 71(2):281–289, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336244.

Godambe:1985:FFS

[God85] V. P. Godambe. The foundations of finite sample estimation in stochastic processes. Biometrika, 72(2):419–428, August 1985. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336094.

Goffinet:1987:ACI

[Gof87] B. Goffinet. Alternative conditions for ignoring the process that causes missing data. Biometrika, 74(2):437–439, June 1987. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336161.

Goldstein:1980:LBR

[Gol80] Michael Goldstein. The linear Bayes regression estimator under weak prior assumptions. Biometrika, 67(3):621–628, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335132.

Goldstein:1986:MML

[Gol86] H. Goldstein. Multilevel mixed linear model analysis using iterative generalized least squares. Biometrika, 73(1):43–56, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336270.

Goldstein:1987:MCC

[Gol87] H. Goldstein. Multilevel covariance component models. Biometrika, 74(2):430–431, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336159.

Gold:1989:GAC

[Gol89a] Morris S. Gold. The geometric approximation to the Cusum run length distribution. Biometrika, 76(4):725–733, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336632.

Goldstein:1989:RUI

[Gol89b] Harvey Goldstein. Restricted unbiased iterative generalized leastsquares estimation. Biometrika, 76(3):622–623, September 1989.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336130.

Goodman:1981:AMB

[Goo81] Leo A. Goodman. Association models and the bivariate normal for contingency tables with ordered categories. Biometrika, 68(2):347–355, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335579.

Goria:1980:SLM

[Gor80] M. N. Goria. Some locally most powerful generalized rank tests. Biometrika, 67(2):497–500, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335499.

Gould:1982:GSM

[GP82] A. Lawrence Gould and Victor J. Pecore. Group sequential methods for clinical trials allowing early acceptance of H_{o} and incorporating costs. Biometrika, 69(1):75–80, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335855.

Gart:1985:EBV

[GPT85] John J. Gart, Hugh M. Pettigrew, and Donald G. Thomas. The effect of bias, variance estimation, skewness and kurtosis of the empirical logit on weighted least squares analyses. Biometrika, 72(1):179– 190, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336348.

Green:1985:LMF

[Gre85] Peter J. Green. Linear models for field trials, smoothing and crossvalidation. Biometrika, 72(3):527–537, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336724.

Greenacre:1988:CAM

[Gre88] Michael J. Greenacre. Correspondence analysis of multivariate categorical data by weighted least-squares. Biometrika, 75(3):457–467, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336595.

Grove:1980:IFE

[Gro80] D. M. Grove. The interpretation of forensic evidence using a likelihood ratio. Biometrika, 67(1):243–246, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335345.

Gill:1985:ENN

[GS85] P. S. Gill and G. K. Shukla. Efficiency of nearest neighbour balanced block designs for correlated observations. Biometrika, 72 (3):539–544, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336725.

Gill:1987:STP

[GS87] Richard Gill and Martin Schumacher. A simple test of the proportional hazards assumption. Biometrika, 74(2):289–300, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336143.

Gupta:1989:OBD

[GS89] V. K. Gupta and Rajeshwar Singh. On E-optimal block designs. Biometrika, 76(1):184–188, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336384.

Gasser:1986:RVR

[GSJS86] Theo Gasser, Lothar Sroka, and Christine Jennen-Steinmetz. Residual variance and residual pattern in nonlinear regression. Biometrika, 73(3):625–633, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336527.

Godambe:1984:RET

[GT84a] V. P. Godambe and M. E. Thompson. Robust estimation through estimating equations. Biometrika, 71(1):115–125, April 1984. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336403.

Gupta:1984:DLR

[GT84b] A. K. Gupta and J. Tang. Distribution of likelihood ratio statistic for testing equality of covariance matrices of multivariate Gaussian models. Biometrika, 71(3):555–559, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336564.

Gail:1988:TNT

[GTP88] M. H. Gail, W. Y. Tan, and S. Piantadosi. Tests for no treatment effect in randomized clinical trials. Biometrika, 75(1):57–64, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336434.

Godolphin:1983:ECM

[GU83] E. J. Godolphin and J. M. Unwin. Evaluation of the covariance matrix for the maximum likelihood estimator of a Gaussian autoregressive-moving average process. Biometrika, 70(1):279–284, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335968.

Gunel:1987:BTW

[Gün87] Erdoḡan Günel. Bernoulli trials and the weight of evidence in favour of a null hypothesis. Biometrika, 74(2):440–442, June 1987. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336162.

Gupta:1989:EDC

[Gup89] Sudhir Gupta. Efficient designs for comparing test treatments with a control. Biometrika, 76(4):783–787, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336638.

Guyon:1982:PES

[Guy82] Xavier Guyon. Parameter estimation for a stationary process on a d-dimensional lattice. Biometrika, 69(1):95–105, April 1982. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335857.

Giovagnoli:1983:BOO

[GV83] Alessandra Giovagnoli and Isabella Verdinelli. Bayes D-optimal and E-optimal block designs. Biometrika, 70(3):695–706, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336507.

Gates:1980:FBD

[GW80] David J. Gates and Mark Westcott. Further bounds for the distribution of minimum interpoint distance on a sphere. Biometrika, 67(2):466–469, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335491.

Gail:1984:BET

[GWP84] M. H. Gail, S. Wieand, and S. Piantadosi. Biased estimates of treatment effect in randomized experiments with nonlinear regressions and omitted covariates. Biometrika, 71(3):431–444, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336553.

Hinde:1987:CLN

[HA87] John Hinde and Murray Aitkin. Canonical likelihoods: A new likelihood treatment of nuisance parameters. Biometrika, 74(1):45–58, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336020.

Haitovsky:1987:MRR

[Hai87] Yoel Haitovsky. On multivariate ridge regression. Biometrika, 74(3):563–570, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336695.

Halperin:1980:SSD

[Hal80] Max Halperin. Sample size determination in retrospective casecontrol studies. Biometrika, 67(3):577–580, December 1980. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335126.

Hall:1981:NMB

[Hal81a] Peter Hall. On nonparametric multivariate binary discrimination. Biometrika, 68(1):287–294, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335829.

Hall:1981:ONN

[Hal81b] Peter Hall. Optimal near neighbour estimator for use in discriminant analysis. Biometrika, 68(2):572–575, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335608.

Hall:1982:CVD

[Hal82a] Peter Hall. Cross-validation in density estimation. Biometrika, 69(2):383–390, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335412.

Hall:1982:INA

[Hal82b] Peter Hall. Improving the normal approximation when constructing one-sided confidence intervals for binomial or Poisson parameters. Biometrika, 69(3):647–652, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336001.

Hall:1985:CSC

[Hal85] Peter Hall. Correcting segment counts for edge effects when estimating intensity. Biometrika, 72(2):459–463, August 1985. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336098.

Hall:1987:BLB

[Hal87] Peter Hall. On the bootstrap and likelihood-based confidence regions. Biometrika, 74(3):481–493, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336687.

Hall:1989:TUR

[Hal89a] Alastair Hall. Testing for a unit root in the presence of moving average errors. Biometrika, 76(1):49–56, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336368. See correction [Hal90].

Hall:1989:ARB

[Hal89b] Peter Hall. Antithetic resampling for the bootstrap. Biometrika, 76(4):713–724, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336631.

Hall:1989:EBS

[Hal89c] Peter Hall. On efficient bootstrap simulation. Biometrika, 76 (3):613–617, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336128.

Hall:1990:ACT

[Hal90] Alastair Hall. Amendments and corrections: “Testing for a unit root in the presence of moving average errors” [Biometrika 76 (1989), no. 1, 49–56; MR0991422 (90g:62218)]. Biometrika, 77(1): 235, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336072.

See [Hal89a].

Hamilton:1986:CRP

[Ham86] David Hamilton. Confidence regions for parameter subsets in nonlinear regression. Biometrika, 73(1):57–64, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336271.

Harris:1985:AEN

[Har85a] P. Harris. An asymptotic expansion for the null distribution of the efficient score statistic. Biometrika, 72(3):653–659, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336738. See cor-

rection [Har87].

Harris:1985:TVH

[Har85b] P. Harris. Testing for variance homogeneity of correlated variables. Biometrika, 72(1):103–107, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336339.

Harris:1986:NBA

[Har86] P. Harris. A note on Bartlett adjustments to likelihood ratio tests. Biometrika, 73(3):735–737, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336541.

Harris:1987:ACA

[Har87] P. Harris. Amendments and corrections: “An asymptotic expansion for the null distribution of the efficient score statistic” [Biometrika 72 (1985), no. 3, 653–659; MR0817580 (87a:62036)]. Biometrika, 74(3):667, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336715. See [Har85a].

Harris:1989:PFN

[Har89] Ian R. Harris. Predictive fit for natural exponential families. Biometrika, 76(4):675–684, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336627.

Hsieh:1983:STS

[HCT83] F. Y. Hsieh, John Crowley, and Douglass C. Tormey. Some test statistics for use in multistate survival analysis. Biometrika, 70 (1):111–119, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335948.

Harrell:1982:NDF

[HD82] Frank E. Harrell and C. E. Davis. A new distribution-free quantile estimator. Biometrika, 69(3):635–640, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335999.

Henery:1986:IAR

[Hen86] Robert J. Henery. Interpretation of average ranks. Biometrika, 73(1):224–227, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336291.

Harrington:1982:CRT

[HF82] David P. Harrington and Thomas R. Fleming. A class of rank test procedures for censored survival data. Biometrika, 69(3):553–566, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335991.

Hall:1988:SCN

[HH88] Peter Hall and E. J. Hannan. On stochastic complexity and nonparametric density estimation. Biometrika, 75(4):705–714, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336311.

Heckman:1989:ICR

[HH89] James J. Heckman and Bo E. Honoré. The identifiability of the competing risks model. Biometrika, 76(2):325–330, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336666.

Hsuan:1982:CSP

[HHP82] F. Hsuan, F. K. Hwang, and M. Parnes. A class of selection problems for which more sampling is more informative. Biometrika, 69(1):280–283, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335887.

Hill:1981:ARE

[Hil81] Catherine Hill. Asymptotic relative efficiency of survival tests with covariates. Biometrika, 68(3):699–702, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335456.

Hillier:1986:JTZ

[Hil86] Grant H. Hillier. Joint tests for zero restrictions on nonnegative regression coefficients. Biometrika, 73(3):657–669, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336530.

Hinkley:1980:LAP

[Hin80] D. V. Hinkley. Likelihood as approximate pivotal distribution. Biometrika, 67(2):287–292, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335473.

Hinkley:1985:TDL

[Hin85] David Hinkley. Transformation diagnostics for linear models. Biometrika, 72(3):487–496, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336721.

Hinkley:1988:MST

[Hin88] D. V. Hinkley. More on score tests for transformation in regression. Biometrika, 75(2):366–369, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336186.

Hirotsu:1982:UCE

[Hir82] C. Hirotsu. Use of cumulative efficient scores for testing ordered alternatives in discrete models. Biometrika, 69(3):567–577, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335992.

Hirotsu:1983:DPA

[Hir83] C. Hirotsu. Defining the pattern of association in two-way contingency tables. Biometrika, 70(3):579–589, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336493.

Hirotsu:1986:CCS

[Hir86] C. Hirotsu. Cumulative chi-squared statistic as a tool for testing goodness of fit. Biometrika, 73(1):165–173, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336282.

Hall:1981:NIB

[HJ81] W. B. Hall and R. G. Jarrett. Nonresolvable incomplete block designs with few replicates. Biometrika, 68(3):617–627, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335444.

Hall:1982:RSD

[HJ82] David L. Hall and Brian L. Joiner. Representations of the space of distributions useful in robust estimation of location. Biometrika, 69(1):55–59, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335852.

Hannan:1984:MAM

[HK84a] E. J. Hannan and L. Kavalieris. A method for autoregressivemoving average estimation. Biometrika, 71(2):273–280, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336243.

Huang:1984:CSI

[HK84b] J. S. Huang and Samuel Kotz. Correlation structure in iterated Farlie–Gumbel–Morgenstern distributions. Biometrika, 71 (3):633–636, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336577.

Hannan:1986:REL

[HKM86] E. J. Hannan, L. Kavalieris, and M. Mackisack. Recursive estimation of linear systems. Biometrika, 73(1):119–133, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336277.

Halperin:1988:SIA

[HLH88] Max Halperin, K. K. Gordon Lan, and Mohamed I. Hamdy. Some implications of an alternative definition of the multiple comparison problem. Biometrika, 75(4):773–778, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336318.

Hallin:1988:TRU

[HLP88] Marc Hallin, Claude Lefèvre, and Madan L. Puri. On timereversibility and the uniqueness of moving average representations for non-Gaussian stationary time series. Biometrika, 75(1):170–171, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336451.

Huda:1984:MMV

[HM84] S. Huda and Rahul Mukerjee. Minimizing the maximum variance of the difference between two estimated responses. Biometrika, 71(2):381–385, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336256.

Hardle:1985:ANS

[HM85] W. Härdle and J. S. Marron. Asymptotic nonequivalence of some bandwidth selectors in nonparametric regression. Biometrika, 72 (2):481–484, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336104.

Hall:1988:BRI

[HM88] Peter Hall and Michael A. Martin. On bootstrap resampling and iteration. Biometrika, 75(4):661–671, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336307.

Haggan:1981:MNR

[HO81] V. Haggan and T. Ozaki. Modelling nonlinear random vibrations using an amplitude-dependent autoregressive time series model. Biometrika, 68(1):189–196, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335819.

Hochberg:1988:SBP

[Hoc88] Yosef Hochberg. A sharper Bonferroni procedure for multiple tests of significance. Biometrika, 75(4):800–802, December 1988. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336325.

Holgate:1981:SHP

[Hol81] Philip Holgate. Studies in the history of probability and statistics. XXXIX. Buffon’s cycloid. Biometrika, 68(3):712–716, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335459. See

correction [Hol82].

Holgate:1982:ACS

[Hol82] Philip Holgate. Amendments and corrections: “Studies in the history of probability and statistics. XXXIX. Buffon’s cycloid” [Biometrika 68 (1981), no. 3, 712–716; MR 82m:01040]. Biometrika, 69(2):491, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335431. See [Hol81].

Holgate:1986:SHP

[Hol86] P. Holgate. Studies in the history of probability and statistics. XLI. Waring and Sylvester on random algebraic equations. Biometrika, 73(1):228–231, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336292.

Hommel:1988:SRM

[Hom88] G. Hommel. A stagewise rejective multiple test procedure based on a modified Bonferroni test. Biometrika, 75(2):383–386, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336190.

Hommel:1989:CTM

[Hom89] Gerhard Hommel. A comparison of two modified Bonferroni procedures. Biometrika, 76(3):624–625, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336131.

Horn:1982:FDS

[Hor82] Paul S. Horn. The finite distribution of t statistics based on two symmetric order statistics as a sum of orthant probabilities.

Biometrika, 69(3):681–682, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336011.

Hosking:1980:ADS

[Hos80] J. R. M. Hosking. The asymptotic distribution of the sample inverse autocorrelations of an autoregressive-moving average process. Biometrika, 67(1):223–226, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335338. See correction [Hos82].

Hosking:1981:FD

[Hos81] J. R. M. Hosking. Fractional differencing. Biometrika, 68(1):165– 176, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335817.

Hosking:1982:ACA

[Hos82] J. R. M. Hosking. Amendments and corrections: “The asymptotic distribution of the sample inverse autocorrelations of an autoregressive-moving average process” [Biometrika 67 (1980), no. 1, 223–226; MR 81h:62155]. Biometrika, 69(2):491, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335432. See [Hos80].

Hosking:1984:TWS

[Hos84] J. R. M. Hosking. Testing whether the shape parameter is zero in the generalized extreme- value distribution. Biometrika, 71 (2):367–374, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336254.

Hosoya:1988:SOF

[Hos88] Yuzo Hosoya. The second-order Fisher information. Biometrika, 75(2):265–274, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336174.

Hougaard:1984:LTM

[Hou84] Philip Hougaard. Life table methods for heterogeneous populations: Distributions describing the heterogeneity. Biometrika, 71(1):75– 83, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336399.

Hougaard:1986:CMF

[Hou86a] Philip Hougaard. A class of multivariate failure time distributions. Biometrika, 73(3):671–678, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336531. See correction [Hou88a].

Hougaard:1986:SMH

[Hou86b] Philip Hougaard. Survival models for heterogeneous populations derived from stable distributions. Biometrika, 73(2):387–396, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336215. See

correction [Hou88b].

Hougaard:1988:ACC

[Hou88a] Philip Hougaard. Amendments and corrections: “A Class of Multivariate Failure Time Distributions”. Biometrika, 75(2):395, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336196. See

[Hou86a].

Hougaard:1988:ACS

[Hou88b] Philip Hougaard. Amendments and corrections: “Survival models for heterogeneous populations derived from stable distributions” [Biometrika 73 (1986), no. 2, 387–396; MR0855898 (88c:62154)]. Biometrika, 75(2):395, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336195. See [Hou86b].

Hollander:1975:TMR

[HP75] Myles Hollander and Frank Proschan. Tests for the mean residual life. Biometrika, 62(3):585–593, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335514. See correction [HP76,

HP80b].

Hollander:1976:CAT

[HP76] Myles Hollander and Frank Proschan. Corrections and amendments: “Tests for the mean residual life” (Biometrika 62 (1975), 585–593). Biometrika, 63(2):412, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335641. See [HP75].

Harris:1980:LPE

[HP80a] P. Harris and H. W. Peers. The local power of the efficient scores test statistic. Biometrika, 67(3):525–529, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335120.

Hollander:1980:ACT

[HP80b] M. Hollander and F. Proschan. Amendments and corrections: “Tests for the mean residual life” [Biometrika 62 (1975), no. 3, 585–593; MR 52 #15917]. Biometrika, 67(1):259, April 1980. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335351. See [HP75].

Hayakawa:1985:ADL

[HP85] Takesi Hayakawa and Madan L. Puri. Asymptotic distributions of likelihood ratio criteria for testing latent roots and latent vectors of a covariance matrix under an elliptical population. Biometrika, 72(2):331–338, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336085.

Hannan:1982:REM

[HR82] E. J. Hannan and J. Rissanen. Recursive estimation of mixed autoregressive-moving average order. Biometrika, 69(1):81–94, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335856. See

correction [HR83].

Hannan:1983:ACR

[HR83] E. J. Hannan and J. Rissanen. Amendments and corrections: “Recursive estimation of mixed autoregressive-moving average order” [Biometrika 69 (1982), no. 1, 81–94; MR0655673 (84e:62136)]. Biometrika, 70(1):303, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335976. See [HR82].

Hermans:1985:NST

[HR85] M. Hermans and J. P. Rasson. A new Sobolev test for uniformity on the circle. Biometrika, 72(3):698–702, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336748.

Hanumara:1980:NAP

[HS80] R. Choudary Hanumara and William F. Strain. A note on the application of the percentage points of the extreme roots of a Wishart matrix. Biometrika, 67(2):501–502, August 1980. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335500.

Heathcote:1981:MMS

[HS81] C. R. Heathcote and M. J. Silvapulle. Minimum mean squared estimation of location and scale parameters under misspecification of the model. Biometrika, 68(2):501–514, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335599.

Hinkley:1987:CBM

[HS87] David Hinkley and Edna Schechtman. Conditional bootstrap methods in the mean-shift model. Biometrika, 74(1):85–93, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336023.

Hinkley:1989:ISN

[HS89] D. V. Hinkley and S. Shi. Importance sampling and the nested bootstrap. Biometrika, 76(3):435–446, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336110.

Hayre:1981:EOR

[HT81a] Lakhbir S. Hayre and Bruce W. Turnbull. Estimation of the odds ratio in the two-armed bandit problem. Biometrika, 68(3):661–668, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335450.

Hayre:1981:SET

[HT81b] Lakhbir S. Hayre and Bruce W. Turnbull. Sequential estimation in two-armed exponential clinical trials. Biometrika, 68(2):411–416, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335586.

Hurvich:1989:RTS

[HT89] Clifford M. Hurvich and Chih-Ling Tsai. Regression and time series model selection in small samples. Biometrika, 76(2):297–307, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336663.

Huggins:1989:SAC

[Hug89] R. M. Huggins. On the statistical analysis of capture experiments. Biometrika, 76(1):133–140, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336377.

Hurvich:1988:MSE

[Hur88] Clifford M. Hurvich. A mean squared error criterion for time series data Windows. Biometrika, 75(3):485–490, September 1988. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336598.

Hall:1980:CBS

[HW80] W. J. Hall and Jon A. Wellner. Confidence bands for a survival curve from censored data. Biometrika, 67(1):133–143, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335326.

Hall:1983:TNB

[HW83] Peter Hall and A. H. Welsh. A test for normality based on the empirical characteristic function. Biometrika, 70(2):485–489, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335564. See

correction [HW84a].

Hall:1984:ACT

[HW84a] P. Hall and A. H. Welsh. Amendments and corrections: “A test for normality based on the empirical characteristic function” [Biometrika 70 (1983), no. 2, 485–489; MR0712037 (85b:62041)]. Biometrika, 71(3):655, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336585. See [HW83].

Hinkley:1984:IJC

[HW84b] David Hinkley and Bo Cheng Wei. Improvements of jackknife confidence limit methods. Biometrika, 71(2):331–339, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336250.

Hamilton:1987:CFR

[HW87] David Hamilton and Douglas Wiens. Correction factors for F ratios in nonlinear regression. Biometrika, 74(2):423–425, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336157.

Hall:1988:NDU

[HW88] Peter Hall and Matthew P. Wand. On nonparametric discrimination using density differences. Biometrika, 75(3):541–547, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336605.

Hall:1987:KDE

[HWC87] Peter Hall, G. S. Watson, and Javier Cabrera. Kernel density estimation with spherical data. Biometrika, 74(4):751–762, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336469.

Amari:1982:GTA

[iA82] Shun ichi Amari. Geometrical theory of asymptotic ancillarity and conditional inference. Biometrika, 69(1):1–17, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335848. See correction [iA83].

Amari:1983:ACG

[iA83] Shun ichi Amari. Amendments and corrections: “Geometrical theory of asymptotic ancillarity and conditional inference” [Biometrika 69 (1982), no. 1, 1–17; MR0655666 (84g:62005)]. Biometrika, 70(1): 303, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335974.

See [iA82].

Ipinyomi:1985:NGC

[IJ85] R. A. Ipinyomi and J. A. John. Nested generalized cyclic rowcolumn designs. Biometrika, 72(2):403–409, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336092.

Isogawa:1980:LPF

[IO80] Yoshiko Isogawa and Masashi Okamoto. Linear prediction in the factor analysis model. Biometrika, 67(2):482–484, August 1980.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335495.

Ireson:1985:IES

[IR85] M. J. Ireson and P. V. Rao. Interval estimation of slope with rightcensored data. Biometrika, 72(3):601–608, December 1985. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336732.

Inaba:1986:MDN

[IS86] T. Inaba and S. Shirahata. Measures of dependence in normal models and exponential models by information gain. Biometrika, 73(2):345–352, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336210.

Jacroux:1980:DCO

[Jac80] Mike Jacroux. On the determination and construction of E-optimal block designs with unequal numbers of replicates. Biometrika, 67(3):661–667, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335136.

James:1986:EEC

[Jam86] Ian R. James. On estimating equations with censored data. Biometrika, 73(1):35–42, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336269.

James:1987:TLS

[Jam87] Ian R. James. Tests for location with k samples and censored data. Biometrika, 74(3):599–607, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336699.

Jarrett:1984:BEF

[Jar84] R. G. Jarrett. Bounds and expansions for Fisher information when the moments are known. Biometrika, 71(1):101–113, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336402.

John:1986:RCD

[JE86] J. A. John and J. A. Eccleston. Row-column α-designs. Biometrika, 73(2):301–306, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336206.

Jennrich:1983:NBL

[Jen83] Robert I. Jennrich. A note on the behaviour of the log rank permutation test under unequal censoring. Biometrika, 70(1):133–137, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335950.

Jennrich:1984:SET

[Jen84] Robert I. Jennrich. Some exact tests for comparing survival curves in the presence of unequal right censoring. Biometrika, 71(1):57–64, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336397.

Jensen:1986:STS

[Jen86] J. L. Jensen. Similar tests and the standardized log likelihood ratio statistic. Biometrika, 73(3):567–572, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336520.

Jennison:1987:EGS

[Jen87] Christopher Jennison. Efficient group sequential tests with unpredictable group sizes. Biometrika, 74(1):155–165, March 1987. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336030.

Jewell:1985:LSR

[Jew85] Nicholas P. Jewell. Least squares regression with data arising from stratified samples of the dependent variable. Biometrika, 72(1):11– 21, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336330.

Jeyaratnam:1982:SCC

[Jey82] S. Jeyaratnam. A sufficient condition on the covariance matrix for F tests in linear models to be valid. Biometrika, 69(3):679–680, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336010.

James:1987:TCP

[JJS87] Barry James, Kang Ling James, and David Siegmund. Tests for a change-point. Biometrika, 74(1):71–83, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336022.

Joshi:1986:TTO

[JL86] P. C. Joshi and S. Lalitha. Tests for two outliers in a linear model. Biometrika, 73(1):236–239, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336294.

Jupp:1980:GCC

[JM80] P. E. Jupp and K. V. Mardia. A general correlation coefficient for directional data and related regression problems. Biometrika, 67(1):163–173, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2335329. See correction [JM81].

Jupp:1981:ACG

[JM81] P. E. Jupp and K. V. Mardia. Amendments and corrections: “A general correlation coefficient for directional data and related regression problems” [Biometrika 67 (1980), no. 1, 163–173; MR 82b:62068]. Biometrika, 68(3):738, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335467. See [JM80].

Joe:1987:EQM

[Joe87] Harry Joe. Estimation of quantiles of the maximum of N observations. Biometrika, 74(2):347–354, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336149.

Johansen:1980:WJA

[Joh80] Søren Johansen. The Welch–James approximation to the distribution of the residual sum of squares in a weighted linear regression. Biometrika, 67(1):85–92, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2335320. See amendments and correction

[Joh82].

Johansen:1982:ACW

[Joh82] Søren Johansen. Amendments and corrections: “The Welch–James approximation to the distribution of the residual sum of squares in a weighted linear regression” [Biometrika 67 (1980), no. 1, 85– 92; MR 81h:62121]. Biometrika, 69(2):491, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335433. See [Joh80].

Johnson:1985:IML

[Joh85] Wesley Johnson. Influence measures for logistic regression: Another point of view. Biometrika, 72(1):59–65, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336335.

Jones:1983:SAE

[Jon83] D. A. Jones. Statistical analysis of empirical models fitted by optimization. Biometrika, 70(1):67–88, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335944.

Jorgensen:1983:MLE

[Jør83] Bent Jørgensen. Maximum likelihood estimation and large-sample inference for generalized linear and nonlinear regression models. Biometrika, 70(1):19–28, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335940.

Jorgensen:1987:JFP

[Jor87] Murray A. Jorgensen. Jackknifing fixed points of iterations. Biometrika, 74(1):207–211, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336037.

Johannes:1980:AIS

[JR80] James M. Johannes and Robert H. Rasche. Additional information on significance values for Durbin’s c^{+}c^{−} and c statistics. Biometrika, 67(2):511–514, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335504.

Jewell:1981:DOC

[JR81] Nicholas P. Jewell and Gillian M. Raab. Difficulties in obtaining consistent estimators of variance parameters. Biometrika, 68 (1):221–226, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335822.

Jennison:1985:RCI

[JT85] Christopher Jennison and Bruce W. Turnbull. Repeated confidence intervals for the median survival time. Biometrika, 72(3):619–625, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336734.

Jupp:1987:NCC

[Jup87] P. E. Jupp. A nonparametric correlation coefficient and a twosample test for random vectors or directions. Biometrika, 74 (4):887–890, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336485.

Jones:1979:SFL

[JW79] David Jones and John Whitehead. Sequential forms of the log rank and modified Wilcoxon tests for censored data. Biometrika, 66(1):105–113, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2335249. See correction [JW81].

Jones:1981:ACS

[JW81] David Jones and John Whitehead. Amendments and corrections: “Sequential forms of the log rank and modified Wilcoxon tests for censored data” [Biometrika 66 (1979), no. 1, 105–113; MR 80i:62069]. Biometrika, 68(2):576, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335610. See [JW79].

Kohn:1983:FIE

[KA83] Robert Kohn and Craig F. Ansley. Fixed interval estimation in state space models when some of the data are missing or aggregated. Biometrika, 70(3):683–688, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336505.

Kohn:1984:NKF

[KA84] Robert Kohn and Craig F. Ansley. A note on Kalman filtering for the seasonal moving average model. Biometrika, 71(3):648–650, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336582.

Kohn:1985:EEP

[KA85] Robert Kohn and Craig F. Ansley. Efficient estimation and prediction in time series regression models. Biometrika, 72(3):694–697, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336747.

Kohn:1986:FFS

[KA86] Robert Kohn and Craig F. Ansley. Fast filtering for seasonal moving average models. Biometrika, 73(2):522–524, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336233.

Kohn:1987:SEF

[KA87] Robert Kohn and Craig F. Ansley. Signal extraction for finite nonstationary time series. Biometrika, 74(2):411–421, June 1987. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336156.

Kohn:1989:FAS

[KA89] Robert Kohn and Craig F. Ansley. A fast algorithm for signal extraction, influence and cross-validation in state space models. Biometrika, 76(1):65–79, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336370.

Kaiser:1983:AEE

[Kai83] Lee Kaiser. Asymptotic equivalence of an expansion test and an approximate degrees of freedom test. Biometrika, 70(2):505–509, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335569.

Kamps:1989:EBE

[Kam89] U. Kamps. Estimation based on equioverlapping samples. Biometrika, 76(4):799–802, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336641.

Kang:1987:TEF

[Kan87] Heejoon Kang. The tapering estimation of the first-order autoregressive parameters. Biometrika, 74(3):643–645, September 1987.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336706.

Kim:1987:DAG

[KD87] Kyungmann Kim and David L. DeMets. Design and analysis of group sequential tests based on the Type I error spending rate function. Biometrika, 74(1):149–154, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336029.

Keefe:1982:RBT

[Kee82] Thomas J. Keefe. On the relationship between two tests for homogeneity of the marginal distributions in a two-way classification. Biometrika, 69(3):683–684, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336012. See correction [Kee84].

Keefe:1984:ACR

[Kee84] Thomas J. Keefe. Amendments and corrections: “On the Relationship Between two Tests for Homogeneity of the Marginal Distributions in a Two-Way Classification”. Biometrika, 71(1):219, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336420. See

[Kee82].

Keenan:1985:TNT

[Kee85] Daniel MacRae Keenan. A Tukey nonadditivity-type test for time series nonlinearity. Biometrika, 72(1):39–44, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336333.

Kempthorne:1984:AVS

[Kem84] Peter J. Kempthorne. Admissible variable-selection procedures when fitting regression models by least squares for prediction. Biometrika, 71(3):593–597, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336569.

Kent:1982:ACR

[Ken82a] John T. Kent. Amendments and corrections: “Robust properties of likelihood ratio tests”. Biometrika, 69(2):492, August 1982. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335434. See [Ken82b].

Kent:1982:RPL

[Ken82b] John T. Kent. Robust properties of likelihood ratio tests. Biometrika, 69(1):19–27, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335849. See correction [Ken82a].

Kent:1983:IGG

[Ken83] John T. Kent. Information gain and a general measure of correlation. Biometrika, 70(1):163–173, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335954.

Kent:1986:USN

[Ken86] John T. Kent. The underlying structure of nonnested hypothesis tests. Biometrika, 73(2):333–343, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336209.

Knott:1983:VEJ

[KF83] M. Knott and C. C. Frangos. Variance estimation for the jackknife using von Mises expansions. Biometrika, 70(2):501–504, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335568.

Kochar:1987:CKT

[KG87] Subhash C. Kochar and R. P. Gupta. Competitors of the Kendall– Tau test for testing independence against positive quadrant dependence. Biometrika, 74(3):664–666, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336712.

Khan:1981:NPT

[Kha81] Rasul A. Khan. A note on Page’s two-sided cumulative sum procedure. Biometrika, 68(3):717–719, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335460.

Kumon:1984:ESP

[KiA84] Masayuki Kumon and Shun ichi Amari. Estimation of a structural parameter in the presence of a large number of nuisance parameters. Biometrika, 71(3):445–459, December 1984. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336554.

Kiefer:1982:TDM

[Kie82] Nicholas M. Kiefer. Testing for dependence in multivariate probit models. Biometrika, 69(1):161–166, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335864.

Kimber:1985:TEW

[Kim85] A. C. Kimber. Tests for the exponential, Weibull and Gumbel distributions based on the stabilized probability plot. Biometrika, 72(3):661–663, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336739.

Kim:1988:IAE

[Kim88] Kyungmann Kim. Improved approximation for estimation following closed sequential tests. Biometrika, 75(1):121–128, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336442.

Korn:1983:IEM

[KL83] Edward L. Korn and Ping Yu Liu. Interactive effects of mixtures of stimuli in life table analysis. Biometrika, 70(1):103–110, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335947.

Kay:1987:TEV

[KL87] Richard Kay and Sarah Little. Transformations of the explanatory variables in the logistic regression model for binary data. Biometrika, 74(3):495–501, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336688.

Kent:1988:MDC

[KO88] John T. Kent and John O’Quigley. Measures of dependence for censored survival data. Biometrika, 75(3):525–534, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336603.

Kochar:1981:NDF

[Koc81] Subhash C. Kochar. A new distribution-free test for the equality of two failure rates. Biometrika, 68(2):423–426, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335588.

Kokolakis:1981:EPC

[Kok81] G. E. Kokolakis. On the expected probability of correct classification. Biometrika, 68(2):477–483, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335596.

Konishi:1981:NTS

[Kon81] Sadanori Konishi. Normalizing transformations of some statistics in multivariate analysis. Biometrika, 68(3):647–651, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335448.

Kott:1986:SAR

[Kot86] Phillip S. Kott. Some asymptotic results for the systematic and stratified sampling of a finite population. Biometrika, 73(2):485– 491, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/

2336226. See correction [Kot87].

Kott:1987:ACS

[Kot87] Phillip S. Kott. Amendments and corrections: “Some asymptotic results for the systematic and stratified sampling of a finite population” [Biometrika 73 (1986), no. 2, 485–491]. Biometrika, 74(3):667, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336716.

See [Kot86].

Kott:1988:MBF

[Kot88] Phillip S. Kott. Model-based finite population correction for the Horvitz–Thompson estimator. Biometrika, 75(4):797–799, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336324.

Koutrouvelis:1980:GFT

[Kou80] I. A. Koutrouvelis. A goodness-of-fit test of simple hypotheses based on the empirical characteristic function. Biometrika, 67(1):238–240, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335343.

Koutrouvelis:1982:ELS

[Kou82] Ioannis A. Koutrouvelis. Estimation of location and scale in Cauchy distributions using the empirical characteristic function. Biometrika, 69(1):205–213, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335870.

Koutras:1986:GNC

[Kou86] Markos Koutras. On the generalized noncentral chi-squared distribution induced by an elliptical gamma law. Biometrika, 73 (2):528–532, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336235.

Koziol:1980:GFT

[Koz80] James A. Koziol. Goodness-of-fit tests for randomly censored data. Biometrika, 67(3):693–696, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335143.

Koziol:1982:CIP

[Koz82] James A. Koziol. A class of invariant procedures for assessing multivariate normality. Biometrika, 69(2):423–427, August 1982. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335417.

Kalbfleisch:1981:EAH

[KP81] John D. Kalbfleisch and Ross L. Prentice. Estimation of the average hazard ratio. Biometrika, 68(1):105–112, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335811.

Kedem:1986:AVH

[KR86] Benjamin Kedem and George Reed. On the asymptotic variance of higher order crossings with special reference to a fast white noise test. Biometrika, 73(1):143–149, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336279.

Kremers:1987:UEW

[KR87] Walter K. Kremers and Douglas S. Robson. Unbiased estimation when sampling from renewal processes: the single sample and ksample random means cases. Biometrika, 74(2):329–336, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336147.

Kraemer:1981:ICD

[Kra81] Helena Chmura Kraemer. Intergroup concordance: Definition and estimation. Biometrika, 68(3):641–646, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335447.

Krzanowski:1983:DBP

[Krz83] W. J. Krzanowski. Distance between populations using mixed continuous and categorical variables. Biometrika, 70(1):235–243, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335961.

Krzanowski:1989:CRC

[Krz89] W. J. Krzanowski. On confidence regions in canonical variate analysis. Biometrika, 76(1):107–116, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336374.

Kedem:1981:GFT

[KS81] Benjamin Kedem and Eric Slud. On goodness of fit of time series models: An application of higher order crossings. Biometrika, 68(2):551–556, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335604.

Kim:1989:LRT

[KS89] Hyune-Ju Kim and David Siegmund. The likelihood ratio test for a change-point in simple linear regression. Biometrika, 76(3):409–423, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336108.

Kass:1989:AMA

[KTK89] Robert E. Kass, Luke Tierney, and Joseph B. Kadane. Approximate methods for assessing influence and sensitivity in Bayesian analysis. Biometrika, 76(4):663–674, December 1989. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336626.

Kuk:1984:ASR

[Kuk84] Anthony Y. C. Kuk. All subsets regression in a proportional hazards model. Biometrika, 71(3):587–592, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336568.

Kuk:1988:EDF

[Kuk88] Anthony Y. C. Kuk. Estimation of distribution functions and medians under sampling with unequal probabilities. Biometrika, 75(1):97–103, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336439.

Kumazawa:1987:NEL

[Kum87] Yoshiki Kumazawa. A note on an estimator of life expectancy with random censorship. Biometrika, 74(3):655–658, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336709.

Kunert:1985:ORM

[Kun85] Joachim Kunert. Optimal repeated measurements designs for correlated observations and analysis by weighted least squares. Biometrika, 72(2):375–389, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336090.

Kunert:1987:NBB

[Kun87a] Joachim Kunert. Neighbour balanced block designs for correlated errors. Biometrika, 74(4):717–724, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336465.

Kunsch:1987:IAR

[Kün87b] Hans R. Künsch. Intrinsic autoregressions and related models on the two-dimensional lattice. Biometrika, 74(3):517–524, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336690.

Kuo:1988:LBE

[Kuo88] Lynn Kuo. Linear Bayes estimators of the potency curve in bioassay. Biometrika, 75(1):91–96, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336438.

Kudo:1982:TTO

[KY82] A. Kudo and J. S. Yao. Tables for testing ordered alternatives in an analysis of variance without replications. Biometrika, 69(1):237– 238, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335875.

Kamakura:1983:ERP

[KY83] T. Kamakura and T. Yanagimoto. Evaluation of the regression parameter estimators in the proportional hazard model. Biometrika, 70(2):530–533, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335575.

Lagakos:1981:GEE

[Lag81] S. W. Lagakos. The graphical evaluation of explanatory variables in proportional hazard regression models. Biometrika, 68(1):93–98, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335809.

Lagakos:1988:LEM

[Lag88] S. W. Lagakos. The loss in efficiency from misspecifying covariates in proportional hazards regression models. Biometrika, 75(1):156– 160, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336448.

Lam:1986:NPS

[Lam86] K. Lam. A new procedure for selecting good populations. Biometrika, 73(1):201–206, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336286.

Larimore:1983:PIS

[Lar83] Wallace E. Larimore. Predictive inference, sufficiency, entropy and an asymptotic likelihood principle. Biometrika, 70(1):175–181, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335955.

Laslett:1982:SCU

[Las82] G. M. Laslett. The survival curve under monotone density constraints with applications to two-dimensional line segment processes. Biometrika, 69(1):153–160, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335863.

Lauder:1981:LVM

[Lau81] I. J. Lauder. Latent variable models for statistical diagnosis. Biometrika, 68(2):365–372, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335581.

Lauder:1983:DKA

[Lau83] I. J. Lauder. Direct kernel assessment of diagnostic probabilities. Biometrika, 70(1):251–256, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335963.

Lawal:1984:CFT

[Law84] H. Bayo Lawal. Comparisons of the X^{2}, Y^{2}, Freeman–Tukey and Williams’s improved G^{2} test statistics in small samples of one-way multinomials. Biometrika, 71(2):415–418, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336263.

Lawless:1986:NLR

[Law86] J. F. Lawless. A note on lifetime regression models. Biometrika, 73(2):509–512, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336229.

Lawrance:1987:SSR

[Law87] A. J. Lawrance. The score statistic for regression transformation. Biometrika, 74(2):275–279, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336141.

Lagakos:1988:NAT

[LBD88] S. W. Lagakos, L. M. Barraj, and V. De Gruttola. Nonparametric analysis of truncated survival data, with application to AIDS. Biometrika, 75(3):515–523, September 1988. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336602.

Lan:1983:DSB

[LD83] K. K. Gordon Lan and David L. DeMets. Discrete sequential boundaries for clinical trials. Biometrika, 70(3):659–663, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336502.

Le:1984:LMC

[Le84a] Chap T. Le. Logistic models for cross-over designs. Biometrika, 71(1):216–217, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336418.

Le:1984:RRE

[Le84b] Chap T. Le. On relative risk estimation under multiple matching. Biometrika, 71(1):195–196, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336413.

Le:1986:ARR

[Le86] Chap T. Le. Addendum: “On relative risk estimation under multiple matching” [Biometrika 71 (1984), no. 1, 195–196; MR0738342 (85e:62212)]. Biometrika, 73(1):245, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336298.

Lee:1985:TFC

[Lee85] Sik Yum Lee. On testing functional constraints in structural equation models. Biometrika, 72(1):125–131, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336342.

Leigh:1988:CEN

[Lei88] G. M. Leigh. A comparison of estimates of natural mortality from fish tagging experiments. Biometrika, 75(2):347–353, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336183.

Leurgans:1983:DFT

[Leu83a] Sue Leurgans. On distribution-free tests for bivariate observations. Biometrika, 70(3):727–728, December 1983. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336513.

Leurgans:1983:TCC

[Leu83b] Sue Leurgans. Three classes of censored data rank tests: Strengths and weaknesses under censoring. Biometrika, 70(3):651–658, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336501.

Leurgans:1987:LMR

[Leu87] Sue Leurgans. Linear models, random censoring and synthetic data. Biometrika, 74(2):301–309, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336144.

Levin:1984:SIC

[Lev84] Bruce Levin. Simple improvements on Cornfield’s approximation to the mean of a noncentral hypergeometric random variable. Biometrika, 71(3):630–632, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336576.

Lim:1980:STT

[LF80] T. K. Lim and Karen Yuen Fung. Sequential trimmed t tests. Biometrika, 67(1):181–186, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335331.

Lakshminarayanan:1984:EPL

[LG84] Mani Y. Lakshminarayanan and Richard F. Gunst. Estimation of parameters in linear structural relationships: Sensitivity to the choice of the ratio of error variances. Biometrika, 71(3):569–573, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336566.

Li:1989:RMT

[LH89] W. K. Li and Y. V. Hui. Robust multiple time series modelling. Biometrika, 76(2):309–315, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336664.

Li:1985:DRA

[Li85] W. K. Li. Distribution of residual autocorrelations in multivariate autoregressive index models. Biometrika, 72(3):686–688, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336745.

Li:1988:GFT

[Li88] W. K. Li. A goodness-of-fit test in robust time series modelling. Biometrika, 75(2):355–361, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336184.

Liang:1983:IAP

[Lia83] Kung-Yee Liang. On information and ancillarity in the presence of a nuisance parameter. Biometrika, 70(3):607–612, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336496.

Liang:1984:AEC

[Lia84] Kung Yee Liang. The asymptotic efficiency of conditional likelihood methods. Biometrika, 71(2):305–313, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336247.

Liang:1985:ORI

[Lia85] Kung-Yee Liang. Odds ratio inference with dependent data. Biometrika, 72(3):678–682, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336743.

Liang:1987:EFA

[Lia87a] Kung-Yee Liang. Estimating functions and approximate conditional likelihood. Biometrika, 74(4):695–702, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336463.

Liang:1987:LMP

[Lia87b] Kung-Yee Liang. A locally most powerful test for homogeneity with many strata. Biometrika, 74(2):259–264, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336139.

Liggett:1988:EEP

[Lig88] Walter Liggett. Estimation of the error probability density from replicate measurements on several items. Biometrika, 75(3):557– 567, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336607.

Liggett:1989:EAD

[Lig89] Walter Liggett. Estimation of an asymmetrical density from several small samples. Biometrika, 76(1):13–21, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336364.

Lindsay:1982:CSF

[Lin82] Bruce Lindsay. Conditional score functions: Some optimality results. Biometrika, 69(3):503–512, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335985.

Ljung:1982:LFS

[Lju82] Greta M. Ljung. The likelihood function for a stationary Gaussian autoregressive-moving average process with missing observations. Biometrika, 69(1):265–268, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335883.

Ljung:1986:DTU

[Lju86] Greta M. Ljung. Diagnostic testing of univariate time series models. Biometrika, 73(3):725–730, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336539.

Loukas:1986:EPE

[LKP86] S. Loukas, C. D. Kemp, and H. Papageorgiou. Even point estimation for the bivariate Poisson distribution. Biometrika, 73(1):222– 223, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336290.

Lam:1983:MWR

[LL83] F. C. Lam and M. T. Longnecker. A modified Wilcoxon rank sum test for paired data. Biometrika, 70(2):510–513, August 1983. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335570.

Lin:1980:STN

[LM80] Ching Chuong Lin and Govind S. Mudholkar. A simple test for normality against asymmetric alternatives. Biometrika, 67(2):455–461, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335489.

Lawless:1984:IAD

[LM84] J. F. Lawless and D. L. McLeish. The information in aggregate data from Markov chains. Biometrika, 71(3):419–430, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336552.

Laara:1985:ETM

[LM85a] E. Laara and J. N. S. Matthews. The equivalence of two models for ordinal data. Biometrika, 72(1):206–207, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336352.

Lockhart:1985:APT

[LM85b] Richard A. Lockhart and Christopher G. McLaren. Asymptotic points for a test of symmetry about a specified median. Biometrika, 72(1):208–210, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336353.

Lee:1986:OPL

[LM86a] Chin-Hui Lee and R. Douglas Martin. Ordinary and proper location M-estimates for autoregressive-moving average models. Biometrika, 73(3):679–686, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336532.

Li:1986:FTS

[LM86b] W. K. Li and A. I. McLeod. Fractional time series modelling. Biometrika, 73(1):217–221, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336289.

Leybourne:1989:DST

[LM89] S. J. Leybourne and B. P. M. McCabe. On the distribution of some test statistics for coefficient constancy. Biometrika, 76(1):169–177, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336381.

Leite:1988:NEM

[LOdBP88] José Galvão Leite, Jorge Oishi, and Carlos Alberto de Bragança Pereira. A note on the exact maximum likelihood estimation of the size of a finite and closed population. Biometrika, 75(1):178– 180, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336454.

Lombard:1987:RTC

[Lom87] F. Lombard. Rank tests for changepoint problems. Biometrika, 74(3):615–624, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336701.

Longford:1987:FSA

[Lon87] Nicholas T. Longford. A fast scoring algorithm for maximum likelihood estimation in unbalanced mixed models with nested random effects. Biometrika, 74(4):817–827, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336476.

Louis:1981:NAA

[Lou81] Thomas A. Louis. Nonparametric analysis of an accelerated failure time model. Biometrika, 68(2):381–390, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335583.

Lagakos:1981:ECP

[LR81] S. W. Lagakos and N. Reid. Estimating convolutions from partially censored data. Biometrika, 68(1):113–117, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335812.

Liang:1985:THO

[LS85a] Kung-Yee Liang and Steven G. Self. Tests for homogeneity of odds ratio when the data are sparse. Biometrika, 72(2):353–358, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336087.

Little:1985:MLE

[LS85b] Roderick J. A. Little and Mark D. Schluchter. Maximum likelihood estimation for mixed continuous and categorical data with missing values. Biometrika, 72(3):497–512, December 1985. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336722.

Lockhart:1985:TFM

[LS85c] R. A. Lockhart and M. A. Stephens. Tests of fit for the von Mises distribution. Biometrika, 72(3):647–652, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336737.

Li:1979:SSP

[LSG79] Shou-Hua Li, Richard M. Simon, and John J. Gart. Small sample properties of the Mantel–Haenszel test. Biometrika, 66(1):181–183, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335263. See

correction [LSG83].

Li:1983:ACS

[LSG83] Shou-Hua Li, Richard M. Simon, and John J. Gart. Amendments and corrections: ‘Small Sample Properties of the Mantel–Haenszel Test’. Biometrika, 70(1):304, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335977. See [LSG79].

Luukkonen:1988:TLA

[LST88] Ritva Luukkonen, Pentti Saikkonen, and Timo Teräsvirta. Testing linearity against smooth transition autoregressive models. Biometrika, 75(3):491–499, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336599.

Lewis:1984:MPC

[LSW84] T. Lewis, I. W. Saunders, and M. Westcott. The moments of the Pearson chi-squared statistic and the minimum expected value in two-way tables. Biometrika, 71(3):515–522, December 1984. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336560. See correction

[LSW89].

Lewis:1989:ACM

[LSW89] T. Lewis, I. W. Saunders, and M. Westcott. Amendments and corrections: “The moments of the Pearson chi-squared statistic and the minimum expected value in two-way tables” [Biometrika 71 (1984), no. 3, 515–522; MR0775397 (86e:62077)]. Biometrika, 76(2): 407, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336678.

See [LSW84].

Lysne:1987:LSP

[LT87] Dan Lysne and Dag Tjøstheim. Loss of spectral peaks in autoregressive spectral estimation. Biometrika, 74(1):200–206, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336036.

Lawal:1980:ADG

[LU80] H. Bayo Lawal and Graham J. G. Upton. An approximation to the distribution of the X^{2} goodness-of-fit statistic for use with small expectations. Biometrika, 67(2):447–453, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335488.

Lubin:1981:EEU

[Lub81] Jay H. Lubin. An empirical evaluation of the use of conditional and unconditional likelihoods for case-control data. Biometrika, 68(2):567–571, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335607.

Lundy:1985:AAA

[Lun85] M. Lundy. Applications of the annealing algorithm to combinatorial problems in statistics. Biometrika, 72(1):191–198, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336349.

Lustbader:1980:TDC

[Lus80] Edward D. Lustbader. Time-dependent covariates in survival analysis. Biometrika, 67(3):697–698, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335144.

Lutkepohl:1982:DPM

[Lüt82] Helmut Lütkepohl. Discounted polynomials for multiple time series model building. Biometrika, 69(1):107–115, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335858.

Liang:1986:LDA

[LZ86] Kung Yee Liang and Scott L. Zeger. Longitudinal data analysis using generalized linear models. Biometrika, 73(1):13–22, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336267.

Machado:1983:TST

[Mac83] S. G. Machado. Two statistics for testing for multivariate normality. Biometrika, 70(3):713–718, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336510.

Mak:1981:LSR

[Mak81] T. K. Mak. Large-sample results in the estimation of a linear transformation. Biometrika, 68(1):323–325, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335835.

Mak:1985:IER

[Mak85] T. K. Mak. Interpretation and estimation in ranking theory. Biometrika, 72(1):227–229, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336358.

Mallet:1986:MLE

[Mal86] A. Mallet. A maximum likelihood estimation method for random coefficient regression models. Biometrika, 73(3):645–656, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336529.

Maller:1988:EMS

[Mal88] R. A. Maller. On the exponential model for survival. Biometrika, 75(3):582–586, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336610.

Mantel:1985:PMH

[Man85] Nathan Mantel. Propriety of the Mantel–Haenszel variance for the log rank test. Biometrika, 72(2):471–472, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336101.

Marshall:1980:AET

[Mar80] R. J. Marshall. Autocorrelation estimation of time series with randomly missing observations. Biometrika, 67(3):567–570, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335124.

Martin:1981:NAF

[Mar81] R. J. Martin. A note on the adequacy of fit of a spatial model to some agricultural uniformity data. Biometrika, 68(1):336–338, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335840.

Marriott:1982:OMC

[Mar82a] F. H. C. Marriott. Optimization methods of cluster analysis. Biometrika, 69(2):417–421, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335416.

Martin:1982:CRB

[Mar82b] R. Douglas Martin. The Cramér–Rao bound and robust Mestimates for autoregressions. Biometrika, 69(2):437–442, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335419.

Martin:1982:SAE

[Mar82c] R. J. Martin. Some aspects of experimental design and analysis when errors are correlated. Biometrika, 69(3):597–612, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335995.

Markowski:1984:IUN

[Mar84] Edward P. Markowski. Inference using near-neighbour trimmed rank statistics for simple linear regression models. Biometrika, 71(1):51–56, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336396.

Martin:1986:DEU

[Mar86] R. J. Martin. On the design of experiments under spatial correlation. Biometrika, 73(2):247–277, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336203. See correction [Mar88].

Martin:1988:ACD

[Mar88] R. J. Martin. Amendments and corrections: “On the design of experiments under spatial correlation” [Biometrika 73 (1986), no. 2, 247–277; MR0855887 (88a:62193)]. Biometrika, 75(2):396, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336197. See

[Mar86].

Masarotto:1987:RCE

[Mas87] Guido Masarotto. Robust and consistent estimates of autoregressivemoving average parameters. Biometrika, 74(4):791–797, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336473.

Matloff:1981:URF

[Mat81] Norman S. Matloff. Use of regression functions for improved estimation of means. Biometrika, 68(3):685–689, December 1981. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335453.

Matthews:1987:OCD

[Mat87] J. N. S. Matthews. Optimal crossover designs for the comparison of two treatments in the presence of carryover effects and autocorrelated errors. Biometrika, 74(2):311–320, June 1987. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336145. See correction

[Mat88b].

Matthews:1988:LBC

[Mat88a] David E. Matthews. Likelihood-based confidence intervals for functions of many parameters. Biometrika, 75(1):139–144, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336444.

Matthews:1988:ACO

[Mat88b] J. N. S. Matthews. Amendments and corrections: “Optimal crossover designs for the comparison of two treatments in the presence of carryover effects and autocorrelated errors” [Biometrika 74 (1987), no. 2, 311–320; MR0903131 (88g:62141)]. Biometrika, 75(2): 396, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336198.

See [Mat87].

Matthews:1989:EDP

[Mat89] J. N. S. Matthews. Estimating dispersion parameters in the analysis of data from crossover trials. Biometrika, 76(2):239–244, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336656.

Mauro:1983:NCK

[Mau83] David W. Mauro. A note on the consistency of Kaplan–Meier least squares estimators. Biometrika, 70(2):534–535, August 1983. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335576.

Mielke:1988:CMA

[MB88] Paul W. Mielke, Jr. and Kenneth J. Berry. Cumulant methods for analysing independence of r-way contingency tables and goodnessof-fit frequency data. Biometrika, 75(4):790–793, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336322.

McGregor:1989:SCD

[MB89] J. R. McGregor and J. C. Babb. Serially correlated differences in the paired comparison of time series. Biometrika, 76(4):735–739, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336633.

Mielke:1981:CNT

[MBBW81] Paul W. Mielke, Kenneth J. Berry, Peter J. Brockwell, and James S. Williams. A class of nonparametric tests based on multiresponse permutation procedures. Biometrika, 68(3):720–724, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335461.

McAleer:1983:ETM

[McA83] Michael McAleer. Exact tests of a model against nonnested alternatives. Biometrika, 70(1):285–288, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335969.

McCullagh:1982:SAQ

[McC82] Peter McCullagh. Some applications of quasisymmetry. Biometrika, 69(2):303–308, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335403.

McCullagh:1984:LS

[McC84a] Peter McCullagh. Local sufficiency. Biometrika, 71(2):233–244, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336239.

McCullagh:1984:TNC

[McC84b] Peter McCullagh. Tensor notation and cumulants of polynomials. Biometrika, 71(3):461–476, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336555.

McLeod:1984:DOP

[McL84] A. I. McLeod. Duality and other properties of multiplicative seasonal autoregressive- moving average models. Biometrika, 71 (1):207–211, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336416.

Mardia:1989:SAS

[MD89] K. V. Mardia and I. L. Dryden. The statistical analysis of shape data. Biometrika, 76(2):271–281, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336660.

Mardia:1982:WDR

[ME82] K. V. Mardia and R. Edwards. Weighted distributions and rotating caps. Biometrika, 69(2):323–330, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335406.

Mehta:1981:SCT

[Meh81] Cyrus R. Mehta. Sequential comparison of two exponential distributions with censored survival data. Biometrika, 68(3):669–675, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335451.

Matthews:1985:SLC

[MF85] D. E. Matthews and V. T. Farewell. On a singularity in the likelihood for a change-point hazard rate model. Biometrika, 72 (3):703–704, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336749.

Miller:1980:REQ

[MH80] Rupert G. Miller and Jerry W. Halpern. Robust estimators for quantal bioassay. Biometrika, 67(1):103–110, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335322.

Miller:1982:RCD

[MH82] Rupert Miller and Jerry Halpern. Regression with censored data. Biometrika, 69(3):521–531, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335987.

Mukerjee:1985:MST

[MH85] Rahul Mukerjee and S. Huda. Minimax second- and third-order designs to estimate the slope of a response surface. Biometrika, 72(1):173–178, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336347.

Mukerjee:1988:ODE

[MH88] Rahul Mukerjee and S. Huda. Optimal design for the estimation of variance components. Biometrika, 75(1):75–80, March 1988. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336436.

Martinez:1981:TDN

[MI81] Jorge Martinez and Boris Iglewicz. A test for departure from normality based on a biweight estimator of scale. Biometrika, 68(1):331–333, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335838.

Michael:1983:SPP

[Mic83] John R. Michael. The stabilized probability plot. Biometrika, 70(1):11–17, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335939.

Milhoj:1981:TFT

[Mil81] Anders Milhøj. A test of fit in time series models. Biometrika, 68(1):177–187, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335818.

Milhoj:1984:BCF

[Mil84] Anders Milhøj. Bias correction in the frequency domain estimation of time series models. Biometrika, 71(1):91–99, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336401.

Minkin:1983:AQA

[Min83] Salomon Minkin. Assessing the quadratic approximation to the log likelihood function in nonnormal linear models. Biometrika, 70(2):367–372, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335550.

Mazumdar:1983:MLE

[MJ83] M. Mazumdar and T. R. Jefferson. Maximum likelihood estimates for multinomial probabilities via geometric programming. Biometrika, 70(1):257–261, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335964.

Munson:1989:CSE

[MJ89] Peter J. Munson and Robert W. Jernigan. A cubic spline extension of the Durbin–Watson test. Biometrika, 76(1):39–47, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336367.

Mitchell:1985:MDE

[MK85a] Ann F. S. Mitchell and Wojtek J. Krzanowski. The Mahalanobis distance and elliptic distributions. Biometrika, 72(2):464–467, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336099. See

correction [MK89].

Mukerjee:1985:RAR

[MK85b] Rahul Mukerjee and Sanpei Kageyama. On resolvable and affine resolvable variance-balanced designs. Biometrika, 72(1):165–172, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336346.

Mitchell:1989:ACM

[MK89] Ann F. S. Mitchell and Wojtek J. Krzanowski. Amendments and corrections: The Mahalanobis Distance and Elliptic Distributions. Biometrika, 76(2):407, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336677. See [MK85a].

Mak:1988:NME

[ML88a] T. K. Mak and W. K. Li. A new method for estimating subgroup means under misclassification. Biometrika, 75(1):105–111, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336440.

Meester:1988:TNE

[ML88b] S. G. Meester and R. A. Lockhart. Testing for normal errors in designs with many blocks. Biometrika, 75(3):569–575, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336608.

Mardia:1984:MLE

[MM84] K. V. Mardia and R. J. Marshall. Maximum likelihood estimation of models for residual covariance in spatial regression. Biometrika, 71(1):135–146, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336405.

Mudholkar:1989:CTE

[MM89] Govind S. Mudholkar and Michael P. McDermott. A class of tests for equality of ordered means. Biometrika, 76(1):161–168, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336380.

Mehrotra:1982:RBT

[MMM82] Kishan G. Mehrotra, Joel E. Michalek, and Daniel Mihalko. A relationship between two forms of linear rank procedures for censored data. Biometrika, 69(3):674–676, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336008.

Mak:1981:AFD

[MN81] T. K. Mak and K. W. Ng. Analysis of familial data: Linearmodel approach. Biometrika, 68(2):457–461, August 1981. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335593.

Moreau:1986:DSA

[MOL86] T. Moreau, J. O’Quigley, and J. Lellouch. On D. Schoenfeld’s approach for testing the proportional hazards assumption. Biometrika, 73(2):513–515, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336230.

Monahan:1984:NES

[Mon84] John F. Monahan. A note on enforcing stationarity in autoregressive-moving average models. Biometrika, 71(2):403–404, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336259.

Moore:1980:MRT

[Moo80] Bruce R. Moore. A modification of the Rayleigh test for vector data. Biometrika, 67(1):175–180, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335330.

Moore:1986:APM

[Moo86] Dirk F. Moore. Asymptotic properties of moment estimators for overdispersed counts and proportions. Biometrika, 73(3):583–588, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336522.

Moran:1980:CND

[Mor80] P. A. P. Moran. Calculation of the normal distribution function. Biometrika, 67(3):675–676, December 1980. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335138.

Morton:1981:EEE

[Mor81a] R. Morton. Efficiency of estimating equations and the use of pivots. Biometrika, 68(1):227–233, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335823.

Morton:1981:EEU

[Mor81b] R. Morton. Estimating equations for an ultrastructural relationship. Biometrika, 68(3):735–737, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335465.

Morton:1987:AEN

[Mor87a] Richard Morton. Asymmetry of estimators in nonlinear regression. Biometrika, 74(4):679–685, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336461.

Morton:1987:GLM

[Mor87b] Richard Morton. A generalized linear model with nested strata of extra-Poisson variation. Biometrika, 74(2):247–257, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336138.

Merikoski:1983:NEP

[MP83] J. K. Merikoski and T. M. Pukkila. A note on the expectation of products of autocorrelations. Biometrika, 70(2):528–529, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335574. See

correction [MP84].

Merikoski:1984:ACN

[MP84] J. K. Merikoski and T. M. Pukkila. Amendments and corrections: “A note on the expectation of products of autocorrelations” [Biometrika 70 (1983), no. 2, 528–529; MR0712046 (84m:62116)]. Biometrika, 71(3):655, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336586. See [MP83].

Mackisack:1989:AFE

[MP89] M. S. Mackisack and D. S. Poskitt. Autoregressive frequency estimation. Biometrika, 76(3):565–575, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336121.

Mehta:1988:CES

[MPW88] Cyrus R. Mehta, Nitin R. Patel, and L. J. Wei. Constructing exact significance tests with restricted randomization rules. Biometrika, 75(2):295–302, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336177.

Manly:1987:CSC

[MR87] Bryan F. J. Manly and J. C. W. Rayner. The comparison of sample covariance matrices using likelihood ratio tests. Biometrika, 74(4):841–847, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336478.

McLeish:1986:LMD

[MS86] D. L. McLeish and Christopher G. Small. Likelihood methods for the discrimination problem. Biometrika, 73(2):397–403, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336216.

Muller:1988:DDS

[MS88] Hans-Georg Müller and Ulrich Stadtmüller. Detecting dependencies in smooth regression models. Biometrika, 75(4):639–650, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336305.

Mukerjee:1989:OEF

[MS89] Rahul Mukerjee and S. Sengupta. Optimal estimation of finite population total under a general correlated model. Biometrika, 76(4):789–794, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336639.

Muller:1987:BCC

[MSS87] Hans-Georg Müller, U. Stadtmüller, and Thomas Schmitt. Bandwidth choice and confidence intervals for derivatives of noisy data.

Biometrika, 74(4):743–749, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336468.

Mudholkar:1980:NAD

[MT80] Govind S. Mudholkar and Madhusudan C. Trivedi. A normal approximation for the distribution of the likelihood ratio statistic in multivariate analysis of variance. Biometrika, 67(2):485–488, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335496.

Murota:1981:SEC

[MT81] Kazuo Murota and Kei Takeuchi. The Studentized empirical characteristic function and its application to test for the shape of distribution. Biometrika, 68(1):55–65, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335805.

McLeish:1983:EEQ

[MT83] D. McLeish and D. Tosh. The estimation of extreme quantiles in logit bioassay. Biometrika, 70(3):625–632, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336498.

Mukerjee:1988:ORI

[Muk88] Hari Mukerjee. Order restricted inference in a repeated measures model. Biometrika, 75(3):616–617, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336618.

Murray:1983:NMO

[Mur83] Gordon D. Murray. Nonconvergence of the minimax order algorithm. Biometrika, 70(2):490–491, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335565.

Mallows:1982:BEE

[MV82] C. L. Mallows and Y. Vardi. Bounds on the efficiency of estimates based on overlapping data. Biometrika, 69(2):287–296, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335401.

Muirhead:1980:ADC

[MW80] Robb J. Muirhead and Christine M. Waternaux. Asymptotic distributions in canonical correlation analysis and other multivariate procedures for nonnormal populations. Biometrika, 67(1):31–43, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335313.

Mardia:1989:MLS

[MW89] K. V. Mardia and A. J. Watkins. On multimodality of the likelihood in the spatial linear model. Biometrika, 76(2):289–295, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336662.

Nagarsenker:1980:TES

[Nag80] P. B. Nagarsenker. On a test of equality of several exponential survival distributions. Biometrika, 67(2):475–478, August 1980. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335493.

Nagaraja:1982:NLF

[Nag82] H. N. Nagaraja. A note on linear functions of ordered correlated normal random variables. Biometrika, 69(1):284–285, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335888.

Nagarsenker:1984:BTH

[Nag84] P. B. Nagarsenker. On Bartlett’s test for homogeneity of variances. Biometrika, 71(2):405–407, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336260.

Nair:1981:PTG

[Nai81] V. N. Nair. Plots and tests for goodness of fit with randomly censored data. Biometrika, 68(1):99–103, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335810.

Nair:1986:TAO

[Nai86] Vijay N. Nair. On testing against ordered alternatives in analysis of variance models. Biometrika, 73(2):493–499, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336227.

Nayak:1988:EPS

[Nay88] Tapan K. Nayak. Estimating population size by recapture sampling. Biometrika, 75(1):113–120, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336441.

Nazaret:1987:BLL

[Naz87] William A. Nazaret. Bayesian log linear estimates for three-way contingency tables. Biometrika, 74(2):401–410, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336155.

Newbold:1980:ETT

[New80a] Paul Newbold. The equivalence of two tests of time series model adequacy. Biometrika, 67(2):463–465, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335490.

Newton:1980:EEM

[New80b] H. J. Newton. Efficient estimation of multivariate moving average autocovariances. Biometrika, 67(1):227–231, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335339.

Ng:1980:EPD

[Ng80] Vee Ming Ng. On the estimation of parametric density functions. Biometrika, 67(2):505–506, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335502.

Nagelkerke:1980:SCS

[NH80] N. J. D. Nagelkerke and A. A. M. Hart. The sequential comparison of survival curves. Biometrika, 67(1):247–249, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335346.

Newton:1984:SCB

[NP84] H. Joseph Newton and Marcello Pagano. Simultaneous confidence bands for autoregressive spectra. Biometrika, 71(1):197–202, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336414.

Nelder:1987:EQL

[NP87] J. A. Nelder and D. Pregibon. An extended quasi-likelihood function. Biometrika, 74(2):221–232, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336136.

Nguyen:1984:ECP

[NRW84] H. T. Nguyen, G. S. Rogers, and E. A. Walker. Estimation in change-point hazard rate models. Biometrika, 71(2):299–304, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336246.

Nomakuchi:1988:TMI

[NS88] Kentaro Nomakuchi and Ning Zhong Shi. A test for a multiple isotonic regression problem. Biometrika, 75(1):181–184, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336455.

Nurminen:1981:AEG

[Nur81] Markku Nurminen. Asymptotic efficiency of general noniterative estimators of common relative risk. Biometrika, 68(2):525–530, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335601.

Oakes:1986:SIM

[Oak86] David Oakes. Semiparametric inference in a model for association in bivariate survival data. Biometrika, 73(2):353–361, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336211.

OBrien:1980:LTM

[O’B80] Peter C. O’Brien. A likelihood test for multivariate serial correlation. Biometrika, 67(3):531–537, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335121. See correction [O’B82].

OBrien:1982:ACL

[O’B82] Peter C. O’Brien. Amendments and corrections: “A likelihood test for multivariate serial correlation” [Biometrika 67 (1980), no. 3, 531–537; MR 82c:62087]. Biometrika, 69(2):492, August 1982. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335435. See [O’B80].

OHagan:1981:MI

[O’H81] A. O’Hagan. A moment of indecision. Biometrika, 68(1):329–330, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335837.

Oja:1981:TLS

[Oja81] Hannu Oja. Two location and scale-free goodness-of-fit tests. Biometrika, 68(3):637–640, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335446.

Oja:1983:NTN

[Oja83] Hannu Oja. New tests for normality. Biometrika, 70(1):297–299, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335972.

ONeill:1982:THF

[O’N82] M. E. O’Neill. On tests of homogeneity of factorially-structure populations; an additive approach. Biometrika, 69(1):167–174, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335865.

Ochi:1984:LIC

[OP84] Y. Ochi and Ross L. Prentice. Likelihood inference in a correlated probit regression model. Biometrika, 71(3):531–543, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336562.

Oman:1988:SBD

[OS88] Samuel D. Oman and Esther Seiden. Switch-back designs. Biometrika, 75(1):81–89, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336437.

Olkin:1981:MLE

[OV81] Ingram Olkin and Michael Væth. Maximum likelihood estimation in a two-way analysis of variance with correlated errors in one classification. Biometrika, 68(3):653–660, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335449.

Owen:1988:ELR

[Owe88] Art B. Owen. Empirical likelihood ratio confidence intervals for a single functional. Biometrika, 75(2):237–249, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336172.

Ozturk:1986:TEV

[Özt86] Aydın Öztürk. On the W test for the extreme value distribution. Biometrika, 73(3):738–740, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336542.

Palmgren:1981:FIM

[Pal81] Juni Palmgren. The Fisher information matrix for log linear models arguing conditionally on observed explanatory variables. Biometrika, 68(2):563–566, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335606.

Palmgren:1987:PDS

[Pal87] Juni Palmgren. Precision of double sampling estimators for comparing two probabilities. Biometrika, 74(4):687–694, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336462.

Parr:1983:NJB

[Par83] William C. Parr. A note on the jackknife, the bootstrap and the delta method estimators of bias and variance. Biometrika, 70(3):719–722, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336511.

Paterson:1983:CEI

[Pat83a] Lindsay Paterson. Circuits and efficiency in incomplete block designs. Biometrika, 70(1):215–225, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335959.

Paterson:1983:UBM

[Pat83b] Lindsay Paterson. An upper bound for the minimum canonical efficiency factor of incomplete block designs. Biometrika, 70 (2):441–446, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335558.

Patefield:1985:IML

[Pat85a] W. M. Patefield. Information from the maximized likelihood function. Biometrika, 72(3):664–668, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336740.

Patel:1985:AID

[Pat85b] H. I. Patel. Analysis of incomplete data in a two-period crossover design with reference to clinical trials. Biometrika, 72(2):411–418, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336093.

Patel:1986:ARM

[Pat86] H. I. Patel. Analysis of repeated measures designs with changing covariates in clinical trials. Biometrika, 73(3):707–715, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336536.

Petruccelli:1986:PTS

[PD86] Joseph D. Petruccelli and Neville Davies. A portmanteau test for self-exciting threshold autoregressive-type nonlinearity in time series. Biometrika, 73(3):687–694, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336533.

Pearlman:1980:AEL

[Pea80] J. G. Pearlman. An algorithm for the exact likelihood of a high-order autoregressive- moving average process. Biometrika, 67(1):232–233, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335340.

Pericchi:1981:BAT

[Per81] L. R. Pericchi. A Bayesian approach to transformations to normality. Biometrika, 68(1):35–43, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335803.

Pericchi:1984:ASB

[Per84] L. R. Pericchi. An alternative to the standard Bayesian procedure for discrimination between normal linear models. Biometrika, 71(3):575–586, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336567.

Pesaran:1984:APC

[Pes84] M. H. Pesaran. Asymptotic power comparisons of tests of separate parametric families by Bahadur’s approach. Biometrika, 71(2):245–252, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336240.

Petkau:1980:FPT

[Pet80a] A. John Petkau. Frequentist properties of three stopping rules for comparative clinical trials. Biometrika, 67(3):690–692, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335142.

Pettitt:1980:SCS

[Pet80b] A. N. Pettitt. A simple cumulative sum type statistic for the changepoint problem with zero-one observations. Biometrika, 67(1):79–84, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335319.

Pettitt:1981:PPC

[Pet81] A. N. Pettitt. Posterior probabilities for a change-point using ranks. Biometrika, 68(2):443–450, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335591.

Pettitt:1982:PTA

[Pet82] A. N. Pettitt. Parametric tests for agreement amongst groups of judges. Biometrika, 69(2):365–375, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335410.

Pettitt:1983:AMU

[Pet83] A. N. Pettitt. Approximate methods using ranks for regression with censored data. Biometrika, 70(1):121–132, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335949.

Pettitt:1984:TGC

[Pet84] A. N. Pettitt. Tied, grouped continuous and ordered categorical data: A comparison of two models. Biometrika, 71(1):35–42, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336394.

Pettitt:1986:COR

[Pet86] A. N. Pettitt. Censored observations, repeated measures and mixed effects models: An approach using the EM algorithm and normal errors. Biometrika, 73(3):635–643, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336528.

Pierce:1982:TNE

[PG82] Donald A. Pierce and Robert J. Gray. Testing normality of errors in regression models. Biometrika, 69(1):233–236, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335874.

Pettigrew:1986:BHC

[PGT86] Hugh M. Pettigrew, John J. Gart, and Donald G. Thomas. The bias and higher cumulants of the logarithm of a binomial variate. Biometrika, 73(2):425–435, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336219.

Phillips:1982:TCF

[Phi82] P. C. B. Phillips. The true characteristic function of the F distribution. Biometrika, 69(1):261–264, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335882.

Phillips:1987:TUA

[Phi87] P. C. B. Phillips. Towards a unified asymptotic theory for autoregression. Biometrika, 74(3):535–547, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336692.

Pierce:1985:TNA

[Pie85] Donald A. Pierce. Testing normality in autoregressive models. Biometrika, 72(2):293–297, August 1985. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336081.

Papageorgiou:1983:SME

[PKL83] H. Papageorgiou, C. D. Kemp, and S. Loukas. Some methods of estimation for the bivariate Hermite distribution. Biometrika, 70(2):479–484, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335563.

Pukkila:1985:FDE

[PN85] Tarmo Pukkila and Hans Nyquist. On the frequency domain estimation of the innovation variance of a stationary univariate time series. Biometrika, 72(2):317–323, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336083.

Porteous:1985:ILR

[Por85a] B. T. Porteous. Improved likelihood ratio statistics for covariance selection models. Biometrika, 72(1):97–101, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336338.

Porteous:1985:NIL

[Por85b] B. T. Porteous. A note on improved likelihood ratio statistics for generalized log linear models. Biometrika, 72(2):473–475, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336102.

Portier:1986:ETO

[Por86] Christopher J. Portier. Estimating the tumour onset distribution in animal carcinogenesis experiments. Biometrika, 73(2):371–378, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336213.

Porteous:1987:MIH

[Por87] B. T. Porteous. The mutual independence hypothesis for categorical data in complex sampling schemes. Biometrika, 74(4):857–862, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336480.

Poskitt:1987:MHR

[Pos87] D. S. Poskitt. A modified Hannan–Rissanen strategy for mixed autoregressive-moving average order determination. Biometrika, 74(4):781–790, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2336472. See correction [Pos90].

Poskitt:1990:ACM

[Pos90] D. S. Poskitt. Amendments and corrections: “A modified Hannan– Rissanen strategy for mixed autoregressive-moving average order determination” [Biometrika 74 (1987), no. 4, 781–790; MR0919846 (90f:62291)]. Biometrika, 77(3):667, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337010. See [Pos87].

Phillips:1988:TUR

[PP88] Peter C. B. Phillips and Pierre Perron. Testing for a unit root in time series regression. Biometrika, 75(2):335–346, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336182.

Pickands:1987:EAI

[PR87a] J. Pickands III and M. Raghavachari. Exact and asymptotic inference for the size of a population. Biometrika, 74(2):355–363, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336150.

Pigeon:1987:CDC

[PR87b] J. G. Pigeon and D. Raghavarao. Crossover designs for comparing treatments with a control. Biometrika, 74(2):321–328, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336146.

Prentice:1978:LRT

[Pre78] R. L. Prentice. Linear rank tests with right censored data. Biometrika, 65(1):167–179, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335292. See correction [Pre83].

Prentice:1982:CME

[Pre82] R. L. Prentice. Covariate measurement errors and parameter estimation in a failure time regression model. Biometrika, 69 (2):331–342, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2335407. See correction [Pre84b].

Prentice:1983:ACL

[Pre83] R. L. Prentice. Amendments and corrections: “Linear rank tests with right censored data” [Biometrika 65 (1978), no. 1, 167–179; MR0497517 (80a:62060)]. Biometrika, 70(1):304, April 1983. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335978. See [Pre78].

Prentice:1984:DFM

[Pre84a] Michael J. Prentice. A distribution-free method of interval estimation for unsigned directional data. Biometrika, 71(1):147–154, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336406.

Prentice:1984:ACC

[Pre84b] R. L. Prentice. Amendments and corrections: “Covariate measurement errors and parameter estimation in a failure time regression model” [Biometrika 69 (1982), no. 2, 331–342; MR0671971 (83j:62146)]. Biometrika, 71(1):219, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336422. See [Pre82].

Prentice:1986:CCD

[Pre86] R. L. Prentice. A case-cohort design for epidemiologic cohort studies and disease prevention trials. Biometrika, 73(1):1–11, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336266.

Prasad:1980:MHT

[PS80] N. G. Narasimha Prasad and T. Srivenkataramana. A modification to the Horvitz–Thompson estimator under the Midzuno sampling scheme. Biometrika, 67(3):709–711, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335148.

Pettitt:1981:EWS

[PS81] A. N. Pettitt and V. Siskind. Effect of within-sample dependence on the Mann–Whitney–Wilcoxon statistic. Biometrika, 68(2):437–441, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335590.

Prasad:1982:SVU

[PS82] Govind Prasad and Ashok Sahai. Sharper variance upper bound for unbiased estimation in inverse sampling. Biometrika, 69(1):286, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335889. See

correction [PS84].

Prasad:1984:ACS

[PS84] Govind Prasad and Ashok Sahai. Amendments and corrections: “Sharper variance upper bound for unbiased estimation in inverse sampling” [Biometrika 69 (1982), no. 1, 286; MR0655698 (83d:62022)]. Biometrika, 71(1):219, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336421. See [PS82].

Pollak:1985:DPA

[PS85] Moshe Pollak and David Siegmund. A diffusion process and its applications to detecting a change in the drift of Brownian motion. Biometrika, 72(2):267–280, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336079.

Poskitt:1980:TSF

[PT80] D. S. Poskitt and A. R. Tremayne. Testing the specification of a fitted autoregressive-moving average model. Biometrika, 67 (2):359–363, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335478.

Poskitt:1983:POT

[PT83] D. S. Poskitt and A. R. Tremayne. On the posterior odds of time series models. Biometrika, 70(1):157–162, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335953.

Poskitt:1987:DPL

[PT87] D. S. Poskitt and A. R. Tremayne. Determining a portfolio of linear time series models. Biometrika, 74(1):125–137, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336027.

Pang:1989:MLE

[PTF89] Ching-Fai Pang, W. A. Thompson, Jr., and Hamid Fallahi. Maximum likelihood estimation of convex arrival rate. Biometrika, 76(3):626–628, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336132.

Padgett:1980:MLE

[PW80a] W. J. Padgett and L. J. Wei. Maximum likelihood estimation of a distribution function with increasing failure rate based on censored observations. Biometrika, 67(2):470–474, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335492.

Prescott:1980:MLE

[PW80b] P. Prescott and A. T. Walden. Maximum likelihood estimation of the parameters of the generalized extreme-value distribution. Biometrika, 67(3):723–724, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335152.

Padgett:1982:ERS

[PW82] W. J. Padgett and L. J. Wei. Estimation of the ratio of scale parameters in the two-sample problem with arbitrary right censorship. Biometrika, 69(1):252–256, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335880.

Paterson:1986:TEF

[PW86] Lindsay J. Paterson and Peter Wild. Triangles and efficiency factors. Biometrika, 73(2):289–299, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336205.

Prentice:1981:RAM

[PWP81] R. L. Prentice, B. J. Williams, and A. V. Peterson. On the regression analysis of multivariate failure time data. Biometrika, 68(2):373–379, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335582.

Quesenberry:1986:RAB

[QJ86] Charles P. Quesenberry, Jr. and Nicholas P. Jewell. Regression analysis based on stratified samples. Biometrika, 73(3):605–614, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336525.

Raftery:1986:BAP

[RA86] A. E. Raftery and V. E. Akman. Bayesian analysis of a Poisson process with a change-point. Biometrika, 73(1):85–89, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336274.

Raftery:1988:IBP

[Raf88] Adrian E. Raftery. Inference for the binomial N parameter: A hierarchical Bayes approach. Biometrika, 75(2):223–228, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336170.

Rayner:1986:NTS

[RB86] J. C. W. Rayner and D. J. Best. Neyman-type smooth tests for location-scale families. Biometrika, 73(2):437–446, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336220.

Reid:1985:IFP

[RC85] N. Reid and H. Crépeau. Influence functions for proportional hazards regression. Biometrika, 72(1):1–9, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336329.

Regal:1980:TTC

[Reg80] Ronald Regal. The F test with time-censored exponential data. Biometrika, 67(2):479–481, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335494.

Reid:1981:EMS

[Rei81] Nancy Reid. Estimating the median survival time. Biometrika, 68(3):601–608, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335442.

Reinsel:1983:SRM

[Rei83] Gregory Reinsel. Some results on multivariate autoregressive index models. Biometrika, 70(1):145–156, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335952.

Rencher:1988:UCI

[Ren88] Alvin C. Rencher. On the use of correlations to interpret canonical functions. Biometrika, 75(2):363–365, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336185.

Reschenhofer:1989:ATW

[Res89] E. Reschenhofer. Adaptive test for white noise. Biometrika, 76 (3):629–632, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336133.

Reilman:1985:SME

[RGL85] Miriam A. Reilman, Richard F. Gunst, and Mani Y. Lakshminarayanan. Structural model estimation with correlated measurement errors. Biometrika, 72(3):669–672, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336741.

Ritov:1989:ESN

[Rit89] Ya’acov Ritov. Estimating a signal with noisy nuisance parameters. Biometrika, 76(1):31–37, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336366.

Rao:1984:SUV

[RL84] J. N. K. Rao and J. Lanke. Simplified unbiased variance estimation for multistage designs. Biometrika, 71(2):387–395, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336257.

Ryan:1988:UCR

[RO88] Louise M. Ryan and E. John Orav. On the use of covariates for rodent bioassay and screening experiments. Biometrika, 75 (4):631–637, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2336304. See correction [RO89].

Ryan:1989:ACU

[RO89] Louise M. Ryan and E. John Orav. Amendments and corrections: “On the Use of Covariates for Rodent Bioassay and Screening Experiments”. Biometrika, 76(2):407, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336679. See [RO88].

Robinson:1983:CST

[Rob83] Derek Robinson. A comparison of sequential treatment allocation rules. Biometrika, 70(2):492–495, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335566.

Rocke:1983:RSA

[Roc83] David M. Rocke. Robust statistical analysis of interlaboratory studies. Biometrika, 70(2):421–431, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335556.

Rocke:1986:ORS

[Roc86] David M. Rocke. Outlier resistance in small samples. Biometrika, 73(1):175–181, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2336283. See correction [Roc90].

Rocke:1990:ACO

[Roc90] David M. Rocke. Amendments and corrections: “Outlier resistance in small samples” [Biometrika 73 (1986), no. 1, 175–181; MR0836445 (88b:62069)] by D. M. Rocke. Biometrika, 77(1):235– 236, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336073.

See [Roc86].

Rosenbaum:1987:SAC

[Ros87] Paul R. Rosenbaum. Sensitivity analysis for certain permutation inferences in matched observational studies. Biometrika, 74(1):13– 26, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336017.

See correction [Ros88a].

Rosenbaum:1988:ACS

[Ros88a] Paul R. Rosenbaum. Amendments and corrections: “Sensitivity analysis for certain permutation inferences in matched observational studies” [Biometrika 74 (1987), no. 1, 13–26; MR0885915 (88c:62114)]. Biometrika, 75(2):396, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336199. See [Ros87].

Rosenbaum:1988:SAM

[Ros88b] Paul R. Rosenbaum. Sensitivity analysis for matching with multiple controls. Biometrika, 75(3):577–581, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336609.

Roy:1989:ACS

[Roy89] Roch Roy. Asymptotic covariance structure of serial correlations in multivariate time series. Biometrika, 76(4):824–827, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336648.

Royall:1982:BSR

[RP82] Richard M. Royall and Dany Pfeffermann. Balanced samples and robust Bayesian inference in finite population sampling. Biometrika, 69(2):401–409, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335414.

Rosenbaum:1983:CRP

[RR83] Paul R. Rosenbaum and Donald B. Rubin. The central role of the propensity score in observational studies for causal effects. Biometrika, 70(1):41–55, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335942.

Rice:1988:FE

[RR88] John A. Rice and Murray Rosenblatt. On frequency estimation. Biometrika, 75(3):477–484, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336597.

Roberts:1987:LRA

[RRK87] G. Roberts, J. N. K. Rao, and S. Kumar. Logistic regression analysis of sample survey data. Biometrika, 74(1):1–12, March 1987.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336016.

Ray:1980:EFR

[RS80] S. K. Ray and Ashok Sahai. Efficient families of ratio and producttype estimators. Biometrika, 67(1):211–215, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335336.

Rubin:1982:FML

[RS82] Donald B. Rubin and Ted H. Szatrowski. Finding maximum likelihood estimates of patterned covariance matrices by the EM algorithm. Biometrika, 69(3):657–660, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336003.

Rosner:1988:ECI

[RT88] Gary L. Rosner and Anastasios A. Tsiatis. Exact confidence intervals following a group sequential trial: A comparison of methods. Biometrika, 75(4):723–729, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336313.

Rukhin:1988:ELE

[Ruk88] Andrew L. Rukhin. Estimating the loss of estimators of a binomial parameter. Biometrika, 75(1):153–155, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336447.

Russell:1980:BCE

[Rus80] K. G. Russell. Balancing carry-over effects in round robin tournaments. Biometrika, 67(1):127–131, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335325.

Robertson:1982:TOA

[RW82] Tim Robertson and F. T. Wright. Testing for ordered alternatives with increased precision in one of the samples. Biometrika, 69(3):579–586, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335993.

Robertson:1983:ALP

[RW83] Tim Robertson and F. T. Wright. On approximation of the level probabilities and associated distributions in order restricted inference. Biometrika, 70(3):597–606, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336495.

Reinsel:1987:ADP

[RW87] Gregory C. Reinsel and Michael A. Wincek. Asymptotic distribution of parameter estimators for nonconsecutively observed time series. Biometrika, 74(1):115–124, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336026.

Rao:1989:SCP

[RW89] C. Radhakrishna Rao and Yue Hua Wu. A strongly consistent procedure for model selection in a regression problem. Biometrika, 76(2):369–374, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336671.

Ryall:1981:ECL

[Rya81] T. A. Ryall. Extensions of the concept of local ancillarity. Biometrika, 68(3):677–683, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335452.

Raab:1987:ECL

[RZ87] Gillian M. Raab and Yu Jing Zhou. The effect of changes of location on least-squares estimators for samples stratified on the dependent variable. Biometrika, 74(1):216–219, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336039.

Samiuddin:1981:SPU

[SA81] M. Samiuddin and H. Asad. A simple procedure of unequal probability sampling. Biometrika, 68(3):728–731, December 1981. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335463.

Samuels:1981:MDE

[Sam81] Myra L. Samuels. Matching and design efficiency in epidemiological studies. Biometrika, 68(3):577–588, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335440.

Sarndal:1980:IWV

[Sär80] Carl-Erik Särndal. On π-inverse weighting versus best linear unbiased weighting in probability sampling. Biometrika, 67(3):639–650, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335134.

Sasabuchi:1980:TMN

[Sas80] S. Sasabuchi. A test of a multivariate normal mean with composite hypotheses determined by linear inequalities. Biometrika, 67(2):429–439, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335486.

Saw:1983:DUV

[Saw83] John G. Saw. Dependent unit vectors. Biometrika, 70(3):665–671, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336503.

Sandland:1984:SIP

[SC84] R. L. Sandland and R. M. Cormack. Statistical inference for Poisson and multinomial models for capture- recapture experiments. Biometrika, 71(1):27–33, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336393.

Suen:1985:BFD

[SC85] Chung-Yi Suen and I. M. Chakravarti. Balanced factorial designs with two-way elimination of heterogeneity. Biometrika, 72 (2):391–402, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336091.

Stefanski:1987:CSO

[SC87] Leonard A. Stefanski and Raymond J. Carroll. Conditional scores and optimal scores for generalized linear measurement- error models. Biometrika, 74(4):703–716, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336464.

Schoenfeld:1980:CSG

[Sch80] David Schoenfeld. Chi-squared goodness-of-fit tests for the proportional hazards regression model. Biometrika, 67(1):145–153, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335327.

Schervish:1981:AEM

[Sch81a] Mark J. Schervish. Asymptotic expansions for the means and variances of error rates. Biometrika, 68(1):295–299, April 1981. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335830.

Schoenfeld:1981:APN

[Sch81b] David Schoenfeld. The asymptotic properties of nonparametric tests for comparing survival distributions. Biometrika, 68(1):316– 319, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335833.

Schoenfeld:1982:PRP

[Sch82] David Schoenfeld. Partial residuals for the proportional hazards regression model. Biometrika, 69(1):239–241, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335876.

Schneider:1984:SHE

[Sch84a] Helmut Schneider. Simple and highly efficient estimators for censored normal samples. Biometrika, 71(2):412–414, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336262.

Schott:1984:OBD

[Sch84b] James R. Schott. Optimal bounds for the distributions of some test criteria for tests of dimensionality. Biometrika, 71(3):561–567, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336565.

Schluchter:1985:ART

[Sch85] Mark D. Schluchter. An aligned rank test for censored data from randomized block designs. Biometrika, 72(3):609–618, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336733. See cor-

rection [Sch87b].

Schafer:1987:CME

[Sch87a] Daniel W. Schafer. Covariate measurement error in generalized linear models. Biometrika, 74(2):385–391, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336153.

Schluchter:1987:ACA

[Sch87b] M. D. Schluchter. Amendments and corrections: “An aligned rank test for censored data from randomized block designs” [Biometrika 72 (1985), no. 3, 609–618; MR0817575 (87c:62091)]. Biometrika, 74(3):668, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336717. See [Sch85].

Schott:1988:CPC

[Sch88a] James R. Schott. Common principal component subspaces in two groups. Biometrika, 75(2):229–236, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336171. See correction [Sch89].

Schott:1988:TES

[Sch88b] James R. Schott. Testing the equality of the smallest latent roots of a correlation matrix. Biometrika, 75(4):794–796, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336323.

Schott:1989:ACC

[Sch89] James R. Schott. Amendments and corrections: “Common principal component subspaces in two groups” [Biometrika 75 (1988), no. 2, 229–236; MR0946048 (89c:62098)]. Biometrika, 76(2):408, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336680. See

[Sch88a].

Stefanski:1986:OBS

[SCR86] Leonard A. Stefanski, Raymond J. Carroll, and David Ruppert. Optimally bounded score functions for generalized linear models with applications to logistic regression. Biometrika, 73(2):413–424, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336218.

Said:1984:TUR

[SD84] Saïd E. Saïd and David A. Dickey. Testing for unit roots in autoregressive-moving average models of unknown order. Biometrika, 71(3):599–607, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336570.

Santner:1986:NAJ

[SD86] Thomas J. Santner and Diane E. Duffy. A note on A. Albert and J. A. Anderson’s conditions for the existence of maximum likelihood estimates in logistic regression models. Biometrika, 73(3):755–758, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336546.

Spurrier:1986:AOS

[SE86] John D. Spurrier and Don Edwards. An asymptotically optimal subclass of balanced treatment incomplete block designs for comparisons with a control. Biometrika, 73(1):191–199, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336285.

Sen:1982:TLM

[Sen82] Pranab Kumar Sen. On M tests in linear models. Biometrika, 69(1):245–248, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335878.

Sengupta:1989:CUP

[Sen89] S. Sengupta. On Chao’s unequal probability sampling plan. Biometrika, 76(1):192–196, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336386.

Seber:1981:TLP

[SF81] G. A. F. Seber and R. Felton. Tag loss and the Petersen markrecapture experiment. Biometrika, 68(1):211–219, April 1981. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335821.

Struthers:1989:MMT

[SF89] C. A. Struthers and V. T. Farewell. A mixture model for time to AIDS data with left truncation and an uncertain origin. Biometrika, 76(4):814–817, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336645.

Sugiura:1985:PDT

[SG85] Nariaki Sugiura and Akihide Gomi. Pearson diagrams for truncated normal and truncated Weibull distributions. Biometrika, 72(1):219– 222, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336356.

Schrader:1980:RAV

[SH80] Ronald M. Schrader and Thomas P. Hettmansperger. Robust analysis of variance based upon a likelihood ratio criterion. Biometrika, 67(1):93–101, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335321.

Shoemaker:1982:RET

[SH82] Lewis H. Shoemaker and Thomas P. Hettmansperger. Robust estimates and tests for the one- and two-sample scale models. Biometrika, 69(1):47–53, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335851.

Smith:1988:RRL

[SH88] S. P. Smith and K. Hammond. Rank regression with log gamma residuals. Biometrika, 75(4):741–751, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336315.

Shapiro:1985:ADT

[Sha85] Alexander Shapiro. Asymptotic distribution of test statistics in the analysis of moment structures under inequality constraints. Biometrika, 72(1):133–144, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336343.

Shen:1982:MDB

[She82] S. M. Shen. A method for discriminating between models describing compositional data. Biometrika, 69(3):587–595, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335994.

Shibata:1981:OSR

[Shi81a] Ritei Shibata. An optimal selection of regression variables. Biometrika, 68(1):45–54, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335804. See correction [Shi82].

Shirahata:1981:IRT

[Shi81b] S. Shirahata. Intraclass rank tests for independence. Biometrika, 68(2):451–456, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335592.

Shibata:1982:ACO

[Shi82] Ritei Shibata. Amendments and corrections: “An optimal selection of regression variables”. Biometrika, 69(2):492, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335436. See [Shi81a].

Shibata:1984:AES

[Shi84] Ritei Shibata. Approximate efficiency of a selection procedure for the number of regression variables. Biometrika, 71(1):43–49, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336395.

Shukla:1982:THV

[Shu82] G. K. Shukla. Testing the homogeneity of variances in a two-way classification. Biometrika, 69(2):411–416, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335415.

Shuster:1986:CPE

[Shu86] Jonathan J. Shuster. The cross-product estimator in matched casecontrol studies. Biometrika, 73(1):212–216, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336288.

Sichel:1982:AET

[Sic82] H. S. Sichel. Asymptotic efficiencies of three methods of estimation for the inverse Gaussian–Poisson distribution. Biometrika, 69(2):467–472, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335423.

Siegmund:1980:STR

[Sie80] D. Siegmund. Sequential χ^{2} and F tests and the related confidence intervals. Biometrika, 67(2):389–402, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335482.

Siegel:1982:RRU

[Sie82] Andrew F. Siegel. Robust regression using repeated medians. Biometrika, 69(1):242–244, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335877.

Sasabuchi:1983:MVI

[SIK83] S. Sasabuchi, M. Inutsuka, and D. D. S. Kulatunga. A multivariate version of isotonic regression. Biometrika, 70(2):465–472, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335561.

Simes:1986:IBP

[Sim86] R. J. Simes. An improved Bonferroni procedure for multiple tests of significance. Biometrika, 73(3):751–754, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336545.

Sinha:1987:MCR

[Sin87] Kishore Sinha. A method of construction of regular group divisible designs. Biometrika, 74(2):443–444, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336163.

Sirvanci:1982:ELD

[Sir82] Mete Sirvanci. An estimator of location for dependent data. Biometrika, 69(2):473–476, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335424.

Srivastava:1981:CEP

[SJ81] Surendra K. Srivastava and Harbans Singh Jhajj. A class of estimators of the population mean in survey sampling using auxiliary information. Biometrika, 68(1):341–343, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335842.

Sandland:1981:ESM

[SK81] R. L. Sandland and G. P. Kirkwood. Estimation of survival in marked populations with possibly dependent sighting probabilities. Biometrika, 68(2):531–541, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335602.

Sasabuchi:1985:SAN

[SK85] Syoichi Sasabuchi and D. D. Sarath Kulatunga. Some approximations for the null distribution of the Ē^{2} statistic used in order restricted inference. Biometrika, 72(2):476–480, August 1985. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336103.

Struthers:1986:MPH

[SK86] C. A. Struthers and J. D. Kalbfleisch. Misspecified proportional hazard models. Biometrika, 73(2):363–369, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336212.

Srivastava:1988:EIC

[SK88] M. S. Srivastava and K. J. Keen. Estimation of the interclass correlation coefficient. Biometrika, 75(4):731–739, December 1988. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336314.

Skinner:1983:MPS

[Ski83] C. J. Skinner. Multivariate prediction from selected samples. Biometrika, 70(1):289–292, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335970.

Skinner:1988:CMB

[Ski88] C. J. Skinner. On conditioning for model-based inference in survey sampling. Biometrika, 75(2):275–286, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336175.

Skovgaard:1986:SIO

[Sko86] Ib Skovgaard. Successive improvement of the order of ancillarity. Biometrika, 73(2):516–519, August 1986. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336231.

Saunders:1980:MKG

[SL80] Roy Saunders and Purushottam Laud. The multidimensional Kolmogorov goodness-of-fit test. Biometrika, 67(1):237, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335342.

Sirvanci:1982:CMS

[SL82] Mete Sirvanci and Ibrahim Levent. Cramér–von Mises statistic for testing exponentiality with censored samples. Biometrika, 69 (3):641–646, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2336000. See correction [SL84].

Sirvanci:1984:ACC

[SL84] Mete Sirvanci and Ibrahim Levent. Amendments and corrections: “Cramér–von Mises statistic for testing exponentiality with censored samples” [Biometrika 69 (1982), no. 3, 641–646; MR0695210 (84g:62080)]. Biometrika, 71(1):220, April 1984. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336423. See [SL82].

Slud:1982:CEI

[Slu82] Eric V. Slud. Consistency and efficiency of inferences with the partial likelihood. Biometrika, 69(3):547–552, December 1982. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335990.

Simpson:1986:RNT

[SM86] Douglas G. Simpson and Barry H. Margolin. Recursive nonparametric testing for dose-response relationships subject to downturns at high doses. Biometrika, 73(3):589–596, December 1986. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336523.

Small:1989:PMI

[SM89] Christopher G. Small and D. L. McLeish. Projection as a method for increasing sensitivity and eliminating nuisance parameters. Biometrika, 76(4):693–703, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336629.

Smit:1980:DFB

[Smi80a] C. F. Smit. On a distribution-free Behrens–Fisher test by Hettmansperger & Malin. Biometrika, 67(1):241–242, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335344.

Smith:1980:BEU

[Smi80b] J. Q. Smith. Bayes estimates under bounded loss. Biometrika, 67(3):629–638, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335133.

Smith:1985:MLE

[Smi85] Richard L. Smith. Maximum likelihood estimation in a class of nonregular cases. Biometrika, 72(1):67–90, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336336.

Sandvik:1982:NDF

[SO82] Leiv Sandvik and Birgitta Olsson. A nearly distribution-free test for comparing dispersion in paired samples. Biometrika, 69(2):484–485, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335427.

Solomon:1984:EMR

[Sol84] Patricia J. Solomon. Effect of misspecification of regression models in the analysis of survival data. Biometrika, 71(2):291–298, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336245. See

correction [Sol86].

Solomon:1985:TCV

[Sol85] P. J. Solomon. Transformations for components of variance and covariance. Biometrika, 72(2):233–239, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336076.

Solomon:1986:ACE

[Sol86] P. J. Solomon. Amendments and corrections: “Effect of misspecification of regression models in the analysis of survival data” [Biometrika 71 (1984), no. 2, 291–298; MR0767157 (85j:62108)].

Biometrika, 73(1):245, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336299. See [Sol84].

Stewart:1982:ECM

[SP82] William H. Stewart and Donald A. Pierce. Efficiency of Cox’s model in estimating regression parameters with grouped survival data. Biometrika, 69(3):539–545, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335989.

Spiegelhalter:1980:OTN

[Spi80] D. J. Spiegelhalter. An omnibus test for normality for small samples. Biometrika, 67(2):493–496, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335498.

Spiegelhalter:1983:DTD

[Spi83] D. J. Spiegelhalter. Diagnostic tests of distributional shape. Biometrika, 70(2):401–409, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335554.

Sprott:1980:MLS

[Spr80] D. A. Sprott. Maximum likelihood in small samples: Estimation in the presence of nuisance parameters. Biometrika, 67(3):515–523, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335119.

Slud:1983:DCR

[SR83] Eric V. Slud and Lawrence V. Rubinstein. Dependent competing risks and summary survival curves. Biometrika, 70(3):643–649, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336500.

Sreenath:1989:CSB

[Sre89] P. R. Sreenath. Construction of some balanced incomplete block designs with nested rows and columns. Biometrika, 76(2):399–402, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336675.

Srivenkataramana:1980:DRE

[Sri80] T. Srivenkataramana. A dual to ratio estimator in sample surveys. Biometrika, 67(1):199–204, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335334.

Srivastava:1984:EIC

[Sri84] M. S. Srivastava. Estimation of interclass correlations in familial data. Biometrika, 71(1):177–185, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336410.

Singh:1980:SSP

[SS80] Padam Singh and A. K. Srivastava. Sampling schemes providing unbiased regression estimators. Biometrika, 67(1):205–209, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335335.

Srivastava:1981:SPI

[SS81] A. K. Srivastava and D. Singh. A sampling procedure with inclusion probabilities proportional to size. Biometrika, 68(3):732–734, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335464.

Schweder:1982:PVE

[SS82] T. Schweder and E. Spjotvoll. Plots of P -values to evaluate many tests simultaneously. Biometrika, 69(3):493–502, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335984.

Sellke:1983:SAP

[SS83a] T. Sellke and D. Siegmund. Sequential analysis of the proportional hazards model. Biometrika, 70(2):315–326, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335546.

Singh:1983:TAG

[SS83b] B. B. Singh and G. K. Shukla. A test of autoregression in Gaussian spatial processes. Biometrika, 70(2):523–527, August 1983. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335573.

Singh:1984:BIS

[SS84a] Bahadur Singh and J. Sedransk. Bayesian inference and sample design for regression analysis when there is nonresponse. Biometrika, 71(1):161–170, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336408.

Sugden:1984:IID

[SS84b] R. A. Sugden and T. M. F. Smith. Ignorable and informative designs in survey sampling inference. Biometrika, 71(3):495–506, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336558.

Schmitz:1985:RSC

[SS85] N. Schmitz and R. J. R. Swamy. Remarks on: “Sequential comparison of two Markov chains” [Biometrika 70 (1983), no. 1, 293–296; MR0743002 (86b:62135)] by Swamy. Biometrika, 72(2):485–486, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336105.

See [Swa83].

Schucany:1989:JE

[SS89a] William R. Schucany and Simon J. Sheather. Jackknifing Restimators. Biometrika, 76(2):393–398, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336674.

Singh:1989:TPM

[SS89b] Avinash C. Singh and Brajendra C. Sutradhar. Testing proportions for Markov dependent Bernoulli trials. Biometrika, 76(4):809–813, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336644.

Saxena:1986:UPS

[SSS86] R. R. Saxena, Padam Singh, and A. K. Srivastava. An unequal probability sampling scheme. Biometrika, 73(3):761–763, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336548.

Sarndal:1989:WRT

[SSW89] Carl-Erik Särndal, Bengt Swensson, and Jan H. Wretman. The weighted residual technique for estimating the variance of the general regression estimator of the finite population total. Biometrika, 76(3):527–537, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336118.

Schoenfeld:1987:MLR

[ST87] David A. Schoenfeld and Anastasios A. Tsiatis. A modified log rank test for highly stratified data. Biometrika, 74(1):167–175, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336031.

Srivenkataramana:1989:TPS

[ST89] T. Srivenkataramana and D. S. Tracy. Two-phase sampling for selection with probability proportional to size in sample surveys. Biometrika, 76(4):818–821, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336646.

Steele:1980:ECV

[Ste80] J. Michael Steele. Efron’s conjecture on vulnerability to bias in a method for balancing sequential trials. Biometrika, 67(2):503–504, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335501.

Stephens:1982:UMD

[Ste82] Michael A. Stephens. Use of the von Mises distribution to analyse continuous proportions. Biometrika, 69(1):197–203, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335869.

Stefanski:1985:EME

[Ste85a] Leonard A. Stefanski. The effects of measurement error on parameter estimation. Biometrika, 72(3):583–592, December 1985. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336730.

Steinberg:1985:MRR

[Ste85b] David M. Steinberg. Model robust response surface designs: Scaling two-level factorials. Biometrika, 72(3):513–526, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336723.

Stewart:1987:MPE

[Ste87] Warren E. Stewart. Multiresponse parameter estimation with a new and noninformative prior. Biometrika, 74(3):557–562, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336694.

Stigler:1980:SHP

[Sti80] Stephen M. Stigler. Studies in the history of probability and statistics. XXXVIII. R. H. Smith, a Victorian interested in robustness. Biometrika, 67(1):217–221, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335337.

Stigler:1984:SHP

[Sti84] Stephen M. Stigler. Studies in the history of probability and statistics. XL. Boscovich, Simpson and a 1760 manuscript note on fitting a linear relation. Biometrika, 71(3):615–620, December 1984. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336572.

Susarla:1984:BJT

[STV84] V. Susarla, W. Y. Tsai, and J. Van Ryzin. A Buckley–Jamestype estimator for the mean with censored data. Biometrika, 71 (3):624–625, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336574.

Sundberg:1986:TUM

[Sun86] Rolf Sundberg. Tests for underlying Markovian structure from panel data with partially aggregated states. Biometrika, 73(3): 717–721, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336537.

Smith:1980:NBD

[SV80] A. F. M. Smith and I. Verdinelli. A note on Bayes designs for inference using a hierarchical linear model. Biometrika, 67(3):613–619, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335131.

Swanepoel:1986:BAP

[SvW86] J. W. H. Swanepoel and J. W. J. van Wyk. The bootstrap applied to power spectral density function estimation. Biometrika, 73(1):135– 141, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336278.

Smythe:1983:STR

[SW83] R. T. Smythe and L. J. Wei. Significance tests with restricted randomization design. Biometrika, 70(2):496–500, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335567.

Schneider:1986:ELM

[SW86] Helmut Schneider and Lisa Weissfeld. Estimation in linear models with censored data. Biometrika, 73(3):741–745, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336543.

Singh:1987:APS

[SW87] Bahadur Singh and F. T. Wright. Approximations to the power of some order restricted tests with slippage alternatives. Biometrika, 74(4):863–870, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336481.

Scott:1989:HTC

[SW89] A. J. Scott and C. J. Wild. Hypothesis testing in case-control studies. Biometrika, 76(4):806–808, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336643.

Swamy:1983:SCT

[Swa83] R. J. R. Swamy. Sequential comparison of two Markov chains. Biometrika, 70(1):293–296, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335971. See remarks [SS85].

Sweeting:1984:CPD

[Swe84] Trevor J. Sweeting. On the choice of prior distribution for the Box– Cox transformed linear model. Biometrika, 71(1):127–134, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336404.

Sweeting:1987:ABA

[Swe87] Trevor J. Sweeting. Approximate Bayesian analysis of censored survival data. Biometrika, 74(4):809–816, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336475.

Silverman:1987:BSS

[SY87] B. W. Silverman and G. A. Young. The bootstrap: To smooth or not to smooth? Biometrika, 74(3):469–479, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336686.

Tavare:1983:SDO

[TA83] Simon Tavaré and Patricia M. E. Altham. Serial dependence of observations leading to contingency tables, and corrections to chisquared statistics. Biometrika, 70(1):139–144, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335951.

Tableman:1987:SWC

[Tab87] Mara Tableman. A simple way to construct a two-sample sequential confidence interval. Biometrika, 74(3):625–630, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336702.

Tamhane:1982:NUR

[Tam82] Ajit C. Tamhane. A note on the use of residuals for detecting an outlier in linear regression. Biometrika, 69(2):488–489, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335429.

Tam:1984:OES

[Tam84] S. M. Tam. Optimal estimation in survey sampling under a regression superpopulation model. Biometrika, 71(3):645–647, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336581.

Tam:1986:CBM

[Tam86] S. M. Tam. Characterization of best model-based predictors in survey sampling. Biometrika, 73(1):232–235, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336293.

Tam:1987:ORP

[Tam87] S. M. Tam. Optimality of Royall’s predictor under a Gaussian superpopulation model. Biometrika, 74(3):659–660, September 1987.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336710.

Tam:1988:ADU

[Tam88a] S. M. Tam. Asymptotically design-unbiased predictors in survey sampling. Biometrika, 75(1):175–177, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336453.

Tamura:1988:ERE

[Tam88b] H. Tamura. Estimation of rare errors using expert judgement. Biometrika, 75(1):1–9, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336428.

Taniguchi:1982:EIF

[Tan82] Masanobu Taniguchi. On estimation of the integrals of the fourth order cumulant spectral density. Biometrika, 69(1):117–122, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335859.

Tarone:1985:HTB

[Tar85] Robert E. Tarone. On heterogeneity tests based on efficient scores. Biometrika, 72(1):91–95, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336337.

Tawn:1988:BEV

[Taw88] Jonathan A. Tawn. Bivariate extreme value theory: Models and estimation. Biometrika, 75(3):397–415, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336591.

Taylor:1985:PTS

[Tay85] Jeremy M. G. Taylor. Power transformations to symmetry. Biometrika, 72(1):145–152, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336344.

Taylor:1987:AIC

[Tay87] Charles C. Taylor. Akaike’s information criterion and the histogram. Biometrika, 74(3):636–639, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336704.

Taylor:1989:BCS

[Tay89] Charles C. Taylor. Bootstrap choice of the smoothing parameter in kernel density estimation. Biometrika, 76(4):705–712, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336630.

Thomas:1989:AIR

[TC89] William Thomas and R. Dennis Cook. Assessing influence on regression coefficients in generalized linear models. Biometrika, 76 (4):741–749, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336634.

Thisted:1987:DSW

[TE87] Ronald Thisted and Bradley Efron. Did Shakespeare write a newlydiscovered poem? Biometrika, 74(3):445–455, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336684.

Tiao:1980:HPA

[TG80] G. C. Tiao and M. R. Grupe. Hidden periodic autoregressivemoving average models in time series data. Biometrika, 67(2):365– 373, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2335479.

Tang:1989:ALR

[TGG89] Dei-In Tang, Clare Gnecco, and Nancy L. Geller. An approximate likelihood ratio test for a normal mean vector with nonnegative components with application to clinical trials. Biometrika, 76(3):577–583, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336122.

Tarone:1983:AIC

[TGH83] Robert E. Tarone, John J. Gart, and Walter W. Hauck. On the asymptotic inefficiency of certain noniterative estimators of a common relative risk or odds ratio. Biometrika, 70(2):519–522, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335572.

Thompson:1985:OSP

[Tho85] W. A. Thompson, Jr. Optimal significance procedures for simple hypotheses. Biometrika, 72(1):230–232, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336359.

Thorburn:1986:BAD

[Tho86] Daniel Thorburn. A Bayesian approach to density estimation. Biometrika, 73(1):65–75, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336272.

Tibshirani:1988:VSB

[Tib88] Robert Tibshirani. Variance stabilization and the bootstrap. Biometrika, 75(3):433–444, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336593.

Tibshirani:1989:NPO

[Tib89] Robert Tibshirani. Noninformative priors for one parameter of many. Biometrika, 76(3):604–608, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336126.

Tadikamalla:1982:SFC

[TJ82] Pandu R. Tadikamalla and Norman L. Johnson. Systems of frequency curves generated by transformations of logistic variables. Biometrika, 69(2):461–465, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335422.

Titterington:1983:REP

[TJ83] D. M. Titterington and J.-M. Jiang. Recursive estimation procedures for missing-data problems. Biometrika, 70(3):613–624, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336497.

See correction [TJ89].

Titterington:1989:ACR

[TJ89] D. M. Titterington and J.-M. Jiang. Amendments and corrections: “Recursive estimation procedures for missing-data problems” [Biometrika 70 (1983), no. 3, 613–624; MR0725375 (85g:62144)].

Biometrika, 76(2):408, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336681. See [TJ83].

Tsai:1987:NPL

[TJW87] Wei-Yann Tsai, Nicholas P. Jewell, and Mei-Cheng Wang. A note on the product-limit estimator under right censoring and left truncation. Biometrika, 74(4):883–886, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336484.

Tierney:1989:AMD

[TKK89] Luke Tierney, Robert E. Kass, and Joseph B. Kadane. Approximate marginal densities of nonlinear functions. Biometrika, 76 (3):425–433, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2336109. See correction [TKK91].

Tierney:1991:ACA

[TKK91] Luke Tierney, Robert E. Kass, and Joseph B. Kadane. Amendments and corrections: “Approximate marginal densities of nonlinear functions” [Biometrika 76 (1989), no. 3, 425–433; MR1040637 (91c:62015)]. Biometrika, 78(1):233–234, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336918. See [TKK89].

Tuan:1987:VLN

[TMKG87] Pham Dinh Tuan, Joachim Möcks, Walter Köhler, and Theo Gasser. Variable latencies of noisy signals: Estimation and testing in brain potential data. Biometrika, 74(3):525–533, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336691.

Tong:1982:DDP

[Ton82] H. Tong. Discontinuous decision processes and threshold autoregressive time series modelling. Biometrika, 69(1):274–276, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335885. See cor-

rection [Ton83a].

Tong:1983:ACD

[Ton83a] H. Tong. Amendments and corrections: “Discontinuous decision processes and threshold autoregressive time series modelling”

[Biometrika 69 (1982), no. 1, 274–276; MR0655696 (83g:62145)]. Biometrika, 70(1):304, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335979. See [Ton82].

Tong:1983:NDA

[Ton83b] H. Tong. A note on a delayed autoregressive process in continuous time. Biometrika, 70(3):710–712, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336509.

Tong:1988:NLP

[Ton88] H. Tong. A note on local parameter orthogonality and Levinson– Durbin algorithm. Biometrika, 75(4):788–789, December 1988. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336321.

Toyooka:1982:PEL

[Toy82a] Yasuyuki Toyooka. Prediction error in a linear model with estimated parameters. Biometrika, 69(2):453–459, August 1982. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335421.

Toyooka:1982:SOE

[Toy82b] Yasuyuki Toyooka. Second-order expansion of mean squared error matrix of generalized least squares estimator with estimated parameters. Biometrika, 69(1):269–273, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335884.

Tjostheim:1983:BSC

[TP83] Dag Tjøstheim and Jostein Paulsen. Bias of some commonlyused time series estimates. Biometrika, 70(2):389–399, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335553. See cor-

rection [TP84].

Tjostheim:1984:ACB

[TP84] D. Tjøstheim and J. Paulsen. Amendments and corrections: “Bias of some commonly-used time series estimates” [Biometrika 70 (1983), no. 2, 389–399; MR0712026 (84i:62124)]. Biometrika, 71 (3):656, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336587. See [TP83].

Tsiatis:1985:GST

[TRT85] Anastasios A. Tsiatis, Gary L. Rosner, and David L. Tritchler. Group sequential tests with censored survival data adjusting for covariates. Biometrika, 72(2):365–373, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336089.

Tsai:1986:STF

[Tsa86a] Chih-Ling Tsai. Score test for the first-order autoregressive model with heteroscedasticity. Biometrika, 73(2):455–460, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336222.

Tsay:1986:NTT

[Tsa86b] Ruey S. Tsay. Nonlinearity tests for time series. Biometrika, 73(2):461–466, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336223.

Tsai:1988:ESF

[Tsa88] Wei-Yann Tsai. Estimation of the survival function with increasing failure rate based on left truncated and right censored data. Biometrika, 75(2):319–324, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336180.

Thall:1988:TSS

[TSE88] Peter F. Thall, Richard Simon, and Susan S. Ellenberg. Twostage selection and testing designs for comparative clinical trials. Biometrika, 75(2):303–310, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336178.

Tsiatis:1980:NGF

[Tsi80] Anastasios A. Tsiatis. A note on a goodness-of-fit test for the logistic regression model. Biometrika, 67(1):250–251, April 1980. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335347.

Tsiatis:1981:AJD

[Tsi81] Anastasios A. Tsiatis. The asymptotic joint distribution of the efficient scores test for the proportional hazards model calculated over time. Biometrika, 68(1):311–315, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335832.

Tsay:1985:UCA

[TT85] Ruey S. Tsay and George C. Tiao. Use of canonical analysis in time series model identification. Biometrika, 72(2):299–315, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336082.

Tosteson:1988:ARE

[TT88] Tor D. Tosteson and Anastasios A. Tsiatis. The asymptotic relative efficiency of score tests in a generalized linear model with surrogate covariates. Biometrika, 75(3):507–514, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336601.

Turrero:1989:REG

[Tur89] A. Turrero. On the relative efficiency of grouped and censored survival data. Biometrika, 76(1):125–131, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336376.

Tutz:1986:ACS

[Tut86] Gerhard Tutz. An alternative choice of smoothing for kernel-based density estimates in discrete discriminant analysis. Biometrika, 73(2):405–411, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336217.

Tyler:1982:RET

[Tyl82] David E. Tyler. Radial estimates and the test for sphericity. Biometrika, 69(2):429–436, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335418.

Tyler:1983:REP

[Tyl83] David E. Tyler. Robustness and efficiency properties of scatter matrices. Biometrika, 70(2):411–420, August 1983. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335555. See correction [Tyl84].

Tyler:1984:ACR

[Tyl84] David E. Tyler. Amendments and corrections: “Robustness and efficiency properties of scatter matrices” [Biometrika 70 (1983), no. 2, 411–420]. Biometrika, 71(3):656, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336588. See [Tyl83].

Tyler:1987:SAA

[Tyl87] David E. Tyler. Statistical analysis for the angular central Gaussian distribution on the sphere. Biometrika, 74(3):579–589, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336697.

Upton:1986:ACI

[Upt86] Graham J. G. Upton. Approximate confidence intervals for the mean direction of a von Mises distribution. Biometrika, 73(2):525– 527, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336234.

VanDerWatt:1980:NEB

[Van80] P. Van Der Watt. A note on estimation of bounds of random variables. Biometrika, 67(3):712–714, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335149.

Vardi:1989:MCR

[Var89] Y. Vardi. Multiplicative censoring, renewal processes, deconvolution and decreasing density: Nonparametric estimation. Biometrika, 76(4):751–761, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336635.

Vassiliou:1981:LBN

[Vas81] P.-C. G. Vassiliou. On the limiting behaviour of a nonhomogeneous Markov chain model in manpower systems. Biometrika, 68(2):557–561, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335605.

Vecchia:1985:GCM

[Vec85] A. V. Vecchia. A general class of models for stationary twodimensional random processes. Biometrika, 72(2):281–291, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336080.

Verbyla:1986:CGC

[Ver86] A. P. Verbyla. Conditioning in the growth curve model. Biometrika, 73(2):475–483, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336225.

Verbyla:1988:ARM

[Ver88] A. P. Verbyla. Analysis of repeated measures designs with changing covariates. Biometrika, 75(1):172–174, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336452.

Venzon:1988:OIR

[VM88] David J. Venzon and Suresh H. Moolgavkar. Origin-invariant relative risk functions for case-control and survival studies. Biometrika, 75(2):325–333, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336181.

Velu:1986:RRM

[VRW86] Raja P. Velu, Gregory C. Reinsel, and Dean W. Wichern. Reduced rank models for multiple time series. Biometrika, 73(1):105–118, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336276.

Vuchkov:1980:IED

[VS80] I. N. Vuchkov and E. B. Solakov. The influence of experimental design on robustness to nonnormality of the F test in regression analysis. Biometrika, 67(2):489–492, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335497.

Verbyla:1988:EGC

[VV88] A. P. Verbyla and W. N. Venables. An extension of the growth curve model. Biometrika, 75(1):129–138, March 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336443.

Wahrendorf:1980:ICT

[Wah80] Jürgen Wahrendorf. Inference in contingency tables with ordered categories using Plackett’s coefficient of association for bivariate distributions. Biometrika, 67(1):15–21, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335311.

Walden:1988:PLE

[Wal88] A. T. Walden. On phase-lag estimation from non-Gaussian time series. Biometrika, 75(4):785–787, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336320.

Wang:1987:PIB

[Wan87] Yuchung J. Wang. The probability integrals of bivariate normal distributions: A contingency table approach. Biometrika, 74(1):185– 190, March 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336033.

Wei:1981:ELD

[Wei81] L. J. Wei. Estimation of location difference for fragmentary samples. Biometrika, 68(2):471–476, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335595.

Wei:1982:IEL

[Wei82] L. J. Wei. Interval estimation of location difference with incomplete data. Biometrika, 69(1):249–251, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335879. See correction [Wei83].

Wei:1983:ACI

[Wei83] L. J. Wei. Amendments and corrections: “Interval estimation of location difference with incomplete data” [Biometrika 69 (1982), no. 1, 249–251; MR0655691 (83d:62062)]. Biometrika, 70(1):304, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335980. See

[Wei82].

Wei:1988:ETS

[Wei88] L. J. Wei. Exact two-sample permutation tests based on the randomized play-the-winner rule. Biometrika, 75(3):603–606, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336615.

Welch:1983:MSE

[Wel83] William J. Welch. A mean squared error criterion for the design of experiments. Biometrika, 70(1):205–213, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335958.

Welsh:1985:AAL

[Wel85] A. H. Welsh. An angular approach for linear data. Biometrika, 72(2):441–450, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336096.

Welch:1987:RMM

[Wel87] William J. Welch. Rerandomizing the median in matched-pairs designs. Biometrika, 74(3):609–614, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336700.

Wermuth:1989:MES

[Wer89] Nanny Wermuth. Moderating effects of subgroups in linear models. Biometrika, 76(1):81–92, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336371.

West:1987:SMN

[Wes87] Mike West. On scale mixtures of normal distributions. Biometrika, 74(3):646–648, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336707.

Westfall:1988:RPT

[Wes88] Peter Westfall. Robustness and power of tests for a null variance ratio. Biometrika, 75(2):207–214, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336168.

Wolter:1982:EQE

[WF82] Kirk M. Wolter and Wayne A. Fuller. Estimation of the quadratic errors in variables model. Biometrika, 69(1):175–182, April 1982.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335866.

Whittemore:1987:TDC

[WFBH87] Alice S. Whittemore, Nina Friend, Byron W. Brown, Jr., and Elizabeth A. Holly. A test to detect clusters of disease. Biometrika, 74(3):631–635, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2336703. See correction [WFBH88].

Whittemore:1988:ACT

[WFBH88] Alice S. Whittemore, Nina Friend, Byron W. Brown, Jr., and Elizabeth A. Holly. Amendments and corrections: “A test to detect clusters of disease” [Biometrika 74 (1987), no. 3, 631–635; MR0909368 (88i:62184)]. Biometrika, 75(2):396, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336200. See [WFBH87].

Wieand:1989:FNS

[WGJJ89] Sam Wieand, Mitchell H. Gail, Barry R. James, and Kang L. James. A family of nonparametric statistics for comparing diagnostic markers with paired or unpaired data. Biometrika, 76 (3):585–592, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336123.

Wacholder:1989:AVE

[WGPB89] Sholom Wacholder, Mitchell H. Gail, David Pee, and Ron Brookmeyer. Alternative variance and efficiency calculations for the casecohort design. Biometrika, 76(1):117–123, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336375.

Whittemore:1989:TOR

[WH89] Alice S. Whittemore and Jerry Halpern. Testing odds-ratio equality for several diseases. Biometrika, 76(4):795–798, December 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336640.

Wheeler:1980:QEJ

[Whe80] Robert E. Wheeler. Quantile estimators of Johnson curve parameters. Biometrika, 67(3):725–728, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335153.

Whitehead:1986:BML

[Whi86] John Whitehead. On the bias of maximum likelihood estimation following a sequential test. Biometrika, 73(3):573–581, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336521.

Wild:1980:VEM

[Wil80] C. J. Wild. Variance estimation with matched pairs. Biometrika, 67(2):507–510, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335503.

Wild:1983:FTM

[Wil83] C. J. Wild. Failure time models with matched data. Biometrika, 70(3):633–641, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336499.

Williamson:1984:NPH

[Wil84] John A. Williamson. A note on the proof by H. E. Daniels of the asymptotic efficiency of a maximum likelihood estimator. Biometrika, 71(3):651–653, December 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336583.

Williams:1986:NMF

[Wil86] E. R. Williams. A neighbour model for field experiments. Biometrika, 73(2):279–287, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336204.

Winterbottom:1980:AEI

[Win80] Alan Winterbottom. Asymptotic expansions to improve large sample confidence intervals for system reliability. Biometrika, 67 (2):351–357, August 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335477.

Wehrly:1980:BMD

[WJ80] Thomas E. Wehrly and Richard A. Johnson. Bivariate models for dependence of angular observations and a related Markov process. Biometrika, 67(1):255–256, April 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335349.

Wei:1985:CDT

[WJ85] L. J. Wei and Wayne E. Johnson. Combining dependent tests with incomplete repeated measurements. Biometrika, 72(2):359–364, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336088.

Woolson:1980:RTC

[WL80] Robert F. Woolson and Peter A. Lachenbruch. Rank tests for censored matched pairs. Biometrika, 67(3):597–606, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335129.

Woolson:1981:RTC

[WL81] Robert F. Woolson and Peter A. Lachenbruch. Rank tests for censored randomized block designs. Biometrika, 68(2):427–435, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335589. See

correction [WL84].

Wermuth:1983:GRM

[WL83] Nanny Wermuth and Steffen L. Lauritzen. Graphical and recursive models for contingency tables. Biometrika, 70(3):537–552, December 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336490.

Woolson:1984:ACR

[WL84] R. F. Woolson and P. A. Lachenbruch. Amendments and corrections: “Rank tests for censored matched pairs” [Biometrika 67 (1980), no. 3, 597–606; MR0601097 (82e:62073)]. Biometrika, 71(1): 220, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336424.

See [WL81].

Wolak:1988:DTM

[Wol88] Frank A. Wolak. Duality in testing multivariate hypotheses. Biometrika, 75(3):611–615, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336617.

Wong:1989:LES

[Won89] M. Y. Wong. Likelihood estimation of a simple linear regression model when both variables have error. Biometrika, 76(1):141–148, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336378.

Worsley:1982:IBI

[Wor82] K. J. Worsley. An improved Bonferroni inequality and applications. Biometrika, 69(2):297–302, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335402.

Worsley:1983:PLR

[Wor83] K. J. Worsley. The power of likelihood ratio and cumulative sum tests for a change in a binomial probability. Biometrika, 70(2):455–464, August 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335560.

Worsley:1986:CRT

[Wor86] K. J. Worsley. Confidence regions and tests for a change-point in a sequence of exponential family random variables. Biometrika, 73(1):91–104, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336275.

Williams:1980:NAL

[WR80a] E. R. Williams and D. Ratcliff. A note on the analysis of lattice designs with repeats. Biometrika, 67(3):706–708, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335147.

Winsberg:1980:MTA

[WR80b] S. Winsberg and J. O. Ramsay. Monotonic transformations to additivity using splines. Biometrika, 67(3):669–674, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335137.

Walters:1982:CPO

[WR82] D. E. Walters and J. G. Rowell. Comments on a paper by I. Olkin and M. Vaeth on two-way analysis of variance with correlated errors: “Maximum likelihood estimation in a two-way analysis of variance with correlated errors in one classification” [Biometrika 68 (1981), no. 3, 653–660; MR0637784 (83b:62144)]. Biometrika, 69(3):664–666, December 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336005.

Warnes:1987:PLE

[WR87] J. J. Warnes and B. D. Ripley. Problems with likelihood estimation of covariance functions of spatial Gaussian processes. Biometrika, 74(3):640–642, September 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336705.

Wehrly:1981:ICE

[WS81] Thomas E. Wehrly and Eugene P. Shine. Influence curves of estimators for directional data. Biometrika, 68(1):334–335, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335839.

Wei:1989:IER

[WSM89] L. J. Wei, R. T. Smythe, and C. R. Mehta. Interval estimation with restricted randomization rules. Biometrika, 76(2):363–368, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336670.

Wright:1985:ALP

[WT85] F. T. Wright and Tuan Tran. Approximating the level probabilities in order restricted inference: The simple tree ordering. Biometrika, 72(2):429–439, August 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336095.

Wu:1982:EVR

[Wu82] Chien-Fu Wu. Estimation of variance of the ratio estimator. Biometrika, 69(1):183–189, April 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335867.

Wu:1985:AIS

[Wu85] C. F. J. Wu. Asymptotic inference from sequential design in a nonlinear situation. Biometrika, 72(3):553–558, December 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336727.

Wang:1981:CSE

[WV81] Min Chiang Wang and John Van Ryzin. A class of smooth estimators for discrete distributions. Biometrika, 68(1):301–309, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335831.

Wild:1987:CND

[WW87] P. R. Wild and E. R. Williams. The construction of neighbour designs. Biometrika, 74(4):871–876, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336482.

Wynn:1984:ECB

[Wyn84] H. P. Wynn. An exact confidence band for one-dimensional polynomial regression. Biometrika, 71(2):375–379, August 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336255.

Yamamoto:1981:PMA

[Yam81] Taku Yamamoto. Predictions of multivariate autoregressive-moving average models. Biometrika, 68(2):485–492, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335597.

Yang:1978:DSS

[Yan78] Mark C. K. Yang. Determination of sample size for testing the relation between an incident and a set of random variables in a sample survey. Biometrika, 65(3):635–640, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335916. See correction [Yan80].

Yang:1980:ACD

[Yan80] Mark C. K. Yang. Amendments and corrections: “Determination of Sample Size for Testing the Relation Between an Incident and a Set of Random Variables in a Sample Survey”. Biometrika, 67(3):728, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335154.

See [Yan78].

Yanagawa:1984:CCS

[Yan84] Takashi Yanagawa. Case-control studies: Assessing the effect of a confounding factor. Biometrika, 71(1):191–194, April 1984. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336412.

Yeh:1986:CUO

[Yeh86] Ching-Ming Yeh. Conditions for universal optimality of block designs. Biometrika, 73(3):701–706, December 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336535.

Yeh:1988:BAS

[Yeh88] Lam Yeh. Bayesian approach to single variable sampling plans. Biometrika, 75(2):387–391, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336191.

Young:1982:LNN

[You82] Dennis L. Young. The linear nearest neighbour statistic. Biometrika, 69(2):477–480, August 1982. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335425.

Young:1988:NBC

[You88] G. A. Young. A note on bootstrapping the correlation coefficient. Biometrika, 75(2):370–373, June 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336187.

Zamar:1989:REE

[Zam89] Ruben H. Zamar. Robust estimation in the errors-in-variables model. Biometrika, 76(1):149–160, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336379.

Zeger:1988:RMT

[Zeg88] Scott L. Zeger. A regression model for time series of counts. Biometrika, 75(4):621–629, December 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336303.

Zhang:1989:EPG

[ZF89] Nien Fan Zhang and Robert V. Foutz. Estimating partial group delay. Biometrika, 76(1):57–63, March 1989. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336369.

Zimmerman:1989:UPE

[ZH89] Dale L. Zimmerman and David A. Harville. On the unbiasedness of the Papadakis estimator and other nonlinear estimators of treatment contrasts in field-plot experiments. Biometrika, 76(2):253– 259, June 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336658.

Zheng:1988:TSP

[Zhe88] Zu Kang Zheng. A time sequential plan for nonparametric testing of hypotheses with censored data. Biometrika, 75(3):607–610, September 1988. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2336616.

Zidek:1984:MSD

[Zid84] Jim Zidek. Maximal Simpson-disaggregations of 2 × 2 tables. Biometrika, 71(1):187–190, April 1984. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336411.

Zeger:1985:ABL

[ZLS85] Scott L. Zeger, Kung Yee Liang, and Steven G. Self. The analysis of binary longitudinal data with time-independent covariates. Biometrika, 72(1):31–38, April 1985. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336332.